### v2

# references.bib

Abstract
Early detection of fertilizer-induced stress in tomato plants is crucial for timely crop management interventions and yield optimization. Conventional optical methods detect fertilizer stress in young leaves with difficulty. This study proposes a novel, noninvasive technique for quantifying the density of trichomes-elongated hair-like structures found on plant surfaces-on young leaves using a smartphone. This method exhibits superior detection latency, enabling earlier and more accurate identification of fertilizer stress in tomato plants. Our approach combines augmented reality technology and image processing algorithms to analyze smartphone images of a specialized measurement paper. This measurement paper is applied to a tomato leaf to transfer trichomes onto its adhesive surface. The captured images are then processed through a pipeline involving region of interest extraction, perspective transformation, and illumination correction. Trichome detection and spatial distribution analysis of these preprocessed images yield a robust density metric. We validated our method through experiments on hydroponically grown tomatoes under varying fertilizer concentrations. Using leave-one-out cross-validation (LOOCV), our model achieves a mean area under the precision-recall curve of 0.824 and a receiver operating characteristic curve of 0.641 for predicting additional fertilization needs. Based on LOOCV, quantitative analysis revealed a strong relationship between trichome density and explanatory variables, including nitrate ion concentration, explaining 62.48% of the variation ($R^{2}=0.625$). The predicted and actual trichome densities were strongly correlated ($r=0.794$). This straightforward and cost-effective method overcomes the limitations of traditional techniques, demonstrating the potential of using smartphones for practical plant nutrition diagnosis.
Keywords: Tomato, tomato leaf, fertilizer stress, trichome, smartphone, computer vision, machine learning, precision agriculture, digital agriculture
1 Introduction
Tomato (Solanum lycopersicum), a vegetable that is widely cultivated worldwide, is influenced by various environmental factors that affect its production and quality. The nutrient status of plants, one of the most crucial factors influencing tomato yield, is an important consideration in tomato production. Therefore, accurately assessing plant nutrient status is essential for maintaining high tomato yields. Techniques have been developed to analyze the constituents of fertilizers in soil, media, and nutrient solutions, aiming to estimate the amount of nutrients absorbed by plants [sonneveld1990estimating, chen2015crop] . Additionally, researchers have developed invasive methods to analyze plant nutrients [hochmuth1994efficiency, andersen1999relationships] . Moreover, noninvasive techniques such as visual inspection and sensor-based monitoring are available for assessing plant traits.
Morphological changes can serve as valuable indicators for nutritional analysis, particularly in the context of environmental stressors such as drought and nutrient deficiencies. With respect to drought stress, Klepper et al. conducted an observational study and reported that cotton stem diameter, leaf water potential, and relative leaf water content are closely related to the net radiation above the plant canopy [klepper1971stem]. Wakamori et al. optically analyzed the wilting of tomato leaves and changes in stem diameter in an observational study and demonstrated a correlation between these parameters[wakamori2019optical]. In terms of nutrient stress, Padilla et al. provided a comprehensive review of techniques for monitoring the nitrogen status of vegetable crops via optical methods [padilla2018proximal]. Locascio et al. and Coltman demonstrated a causal relationship between the fertilization rate and tomato yield in interventional studies[locascio1997nitrogen, coltman1988yields]. Tei et al. reported a causal relationship between the fertilization rate and tomato leaf area in an interventional study[tei2002critical]. These findings emphasize that environmental stressors alter the vegetative and reproductive growth patterns observed in plants.
The ability to detect and examine these modifications allows for a more sophisticated understanding of the nutrient levels of plants and their responses to varying environmental conditions. Leaf area and chlorophyll content, which are related to leaf color and spectrum, are commonly used as indicators of the adaptive response of plants to fertilizer stress. dela Torre et al. reviewed various optical indices, leaf area, biomass, and other phenotypes for estimating rice yield via remote sensing[delatorre2021remote]. Cartelat et al. demonstrated that optically assessed contents of leaf polyphenols and chlorophyll in wheat can be used as indicators of nitrogen deficiency[cartelat2005optically]. Padilla et al. reported that chlorophyll meters and canopy reflectance sensors serve as sensitive indicators of N status across a wide range of vegetable crops[padilla2020monitoring]. These methods have led to the development of a diagnostic framework based on the relationships among fertilizer application, phenotypic expression, and yield outcomes. Color and spectral data analysis of plant canopies enables predictions of plant health and productivity. These predictions utilize wide-area imagery from satellites, aircraft, and drones combined with data collected from in-field sampling stations. The aforementioned optical observation techniques, which utilize reflected light from the plant body, effectively address the need for noninvasive and digital monitoring of phenotypes [padilla2018proximal].
Furthermore, techniques for detecting fertilizer stress in tomatoes have been extensively researched. Tables 1 and 2 present a comprehensive comparison of various noninvasive methods for early detection of fertilizer stress in tomato plants. We focused on studies that employed interventional experiments with varying fertilizer levels and used noninvasive measurement techniques on tomatoes. Priority was given to methods that utilized digital measurement tools and enabled measurements as close to the growth point as possible for early stress detection. As shown in Table 1, although these methods may appear equally effective at first glance, the complexity of data processing (visualization, modeling, and interpretation) after measurement must be considered, which varies considerably depending on the selected phenotypic indicator and sensor type.
Moreover, in terms of detection latency, which refers to the delay between the onset of fertilizer stress and its detection, evaluating phenotypic indicators based on the metabolic mechanisms of tomato plants is crucial. Tomatoes exhibit apical growth, where cell differentiation occurs at the top of the plant, and differentiated organs move downward as the plant grows. Considering this mechanism, measuring phenotypic indicators near the shoot apical meristem could lead to earlier and more accurate detection of fertilizer stress. However, as shown in Table 2, existing studies primarily utilize "fully expanded leaves" or even the entire plant as a phenotypic indicator for chlorophyll meters, visible-near infrared spectroscopy, fluorescence spectroscopy, stem diameter measurement, and leaf area measurement. A "fully expanded leaf" is generally considered the 3rd to 5th compound leaf from the apex, excluding the apical meristem and leaf primordia, and has reached its final size and shape. However, a universally accepted definition of a fully expanded leaf does not exist, and its specific characteristics are as follows: The precise distance from the apical meristem may vary depending on factors such as plant species, cultivar, growth stage, and environmental conditions. Therefore, the context of each study must be considered when interpreting the term "fully expanded leaf." Consequently, studies in the literature identifying the phenomenon of fertilizer stress are lacking. This scarcity is especially true in the assessment of young tomato leaves, with the definition of "young leaves" itself being subject to variation.
Table 1: Comparison of Noninvasive Methods for Detection of Fertilizer Stress in Tomato Plants
Method Phenotypic Indicator Instrumentation Conditions to Avoid Data Complexity Proposed Method Trichome density Digital camera Direct sunlight 2D image array Leaf Chlorophyll Content [padilla2015threshold] Chlorophyll a and b content Chlorophyll meter Direct sunlight Scalar value Visible-Near Infrared Spectroscopy [padilla2015threshold] Multiple biochemical components VIS-NIR spectrometer Direct sunlight 1D spectral array Fluorescence Spectroscopy [bahadur2015gas] Multiple fluorescent compounds Fluorescence spectrometer Bright light 1D spectral array Leaf Area [xiao2022relationship] Leaf area Ruler Withering Scalar value Stem Diameter [olagunju2024effect] Stem diameter Caliper Withering Scalar value
* This table includes studies that met the following criteria: noninvasive measurement, tested on tomatoes, interventional experiments with varying fertilizer levels. Priority was given to methods with digital measurements and those enabling measurements near the growth point for early stress detection.
Table 2: Measurement Sites and Sensitivity for Early Detection
[…]
a Measurement location not specified in the original study.
b Sensitivity to fertilizer stress was not quantitatively assessed in the original study; however, a correlation was visually observed in the presented figures.
c Specific sensitivity indicator was not provided in the original study.
d mPR and mROC represent the mean area under the precision-recall curve and the mean area under the receiver operating characteristic curve, respectively (see equations 10 and 11). These metrics evaluate the performance of the model in determining the need for additional fertilization.
2 Related Work
Gianquinto et al. conducted a study using a multispectral radiometer positioned directly above the tomato plant canopy to investigate the correlation between light reflectance, chlorophyll concentration, and total Kjeldahl nitrogen content in an intervention study of fertilizer application. By employing this noncontact arrangement, measurements were conducted without physically interacting with the leaves. They discovered a linear relationship between optical reflectance and chlorophyll concentration; however, no such linear relationship was identified between the optical reflectance and total Kjeldahl nitrogen content [gianquinto2011a]. Similarly, Padilla et al. directly observed the most recently fully expanded leaves via an optical leaf-clip chlorophyll meter (Konica Minolta, SPAD-502). A linear relationship was observed between optical observations and the nitrogen nutrition index under conditions of a flattened leaf surface structure [padilla2015threshold]. In both investigations, the absence of any barriers along the path of light between the observing device and the plant leaves is assumed, highlighting a gap between the specifications of diagnostic equipment for farmers and the current state of research and development.
However, in contrast to other crops, tomatoes have numerous trichomes, which are small projections or hair-like structures, on the surface of their leaves (Figure 1). Trichomes are densely concentrated on young, actively growing leaves and sparsely distributed on older leaves [johnson1975plant]. The presence of surface microstructures, such as trichomes, on plant bodies is known to considerably affect their appearance [kinoshita2005structural, middleton2024selfassembled, thomas2010function]. Given their high density on young tomato leaves, it is plausible that trichomes may influence the effectiveness of optical sensing techniques used to assess the nutrient status of tomato plants. Consequently, the presence of trichomes should be considered a potential practical limitation when optical sensing techniques are used on young tomato leaves to evaluate fertilizer stress. However, current methodologies do not adequately address the impact of trichomes on the optical sensing techniques used for the rapid assessment of tomato plant nutrient status.
Figure 1: Many trichomes-small projections or hair-like structures-are present on the surface of tomato leaves.
Counting trichomes, which are abundant on young tomato leaves, is not easily achievable via visual means. Manually counting trichomes via a microscope is one of the few methods available for quantifying their density. Trichomes are abundant on young tomato leaves and are not easily distinguishable by visual means alone. The lack of a specialized device for assessing trichome density may have hindered progress in this area of research.
The chlorophyll content, which is correlated with leaf color, is a widely used indicator of fertilizer stress in plants. However, it is not the only marker. Trichomes have also been identified as potential indicators of fertilizer stress, as they contribute to plant defense against pests by responding to fertilizer stress and altering plant density [barbour1991interaction, hoffland2000nitrogen]. Various types of trichomes can be observed on the surface of tomato leaves, most of which are short, hair-like trichomes and trichomes with large, spherical cells at their tips (known as type VI trichomes, Figure 2) [wilkens1996resource]. The glandular head of type VI trichomes is connected to the stalk cell at the base by fragile cell walls [bergau2015the].
Figure 2: Type VI trichomes on the tomato plant surface.
This structural arrangement allows the glandular part to easily detach from the leaf surface upon physical stimulation. The mucilage within the glandular part is known to inhibit pest activity.
Researchers have observed multiple instances where trichomes respond to fertilizer stress [barbour1991interaction, hoffland2000nitrogen, li2014the]. For example, Barbour et al. revealed an inverse relationship between fertilizer application, leaf nitrogen content, and trichome density [barbour1991interaction]. Similarly, Hoffland et al. reported a positive correlation between the leaf carbon:nitrogen ratio and trichome density [hoffland2000nitrogen]. In contrast, Leite et al. presented data indicating that trichome density increased over time when less fertilizer was used compared with that of conventional farming methods [leite]. Despite these findings, most previous studies have focused on investigating trichome density within a biological context, with limited exploration of techniques to observe trichome density in agricultural production settings without advanced laboratory equipment. This limitation highlights the need for further research to develop practical methods for assessing trichome density in real-world agricultural environments.
To address this gap, the present study focused on tomato leaf trichomes and explored their potential as indicators for assessing the nutrient status of tomato plants. As depicted in Figure 3, our main hypothesis is that leaf trichome density is influenced by fertilizer level. These findings suggest its potential as an additional indicator of plant nutrient status in conjunction with traditional fertilizer stress markers such as the leaf nitrate content. To this end, a tool for assessing trichome density must be developed. The identification of distinct items within a particular setting containing several objects inside the visual field is commonly considered a computer vision problem [zhao2019object, vincent1991watersheds]. Given this context, we hypothesized that the use of cameras could convert the challenge of measuring trichome density to an analogy of an object detection problem.
Figure 3: Hypothesized relationships between fertilizer level, cell differentiation, leaf nitrate content, number of leaf trichomes, and crop yield for tomato plants, serving as the conceptual framework for the study.
Recently, the increasing prevalence of smartphones, with their advanced computational capabilities, high-resolution cameras, and internet access, has spurred the development of proof-of-concept agricultural technologies. These technologies have a wide range of applications, including farm management, crop monitoring, and pest and disease diagnosis, as highlighted in a comprehensive review by Mendes et al. [mendes2020smartphone]. For example, Neumann et al. demonstrated a method for classifying diseases in sugar beet leaves using images captured directly with a smartphone camera [neumann20142014], and Johannes et al. similarly classified wheat diseases via direct leaf images [elhassouny20192019]. Furthermore, Li et al. successfully achieved early detection of late blight in tomato leaves using measurement sheet[li2019noninvasive]. In their study, images of a measurement sheet used to analyze leaf volatiles were captured via a smartphone. Inspired by these advancements, our study aimed to investigate the feasibility of using smartphones as a cost-effective tool for measuring trichome density. These parameters can serve as predictive indicators of tomato plant nutrient status.
3 Materials and Methods
3.1 Development of a Smartphone-Based Trichome Density Measurement Method
3.1.1 Design of a simple kit for collecting trichomes from tomato leaves
In this study, we designed a straightforward kit to measure the density of type VI glandular trichomes on tomato leaves. This kit comprises custom-made measurement paper crafted from inkjet-printable cardstock and cellophane tape (NICHIBAN, CT-18), enabling cost-effective measurements (Figure 4). This paper contains augmented reality (AR) markers, enabling the distance of objects within captured images to be precisely measured. Additionally, there is a 12 mm $\times$ 12 mm opening in the center of the paper. A piece of cellophane tape is affixed to the back of the paper, covering this opening. This tape is used to extract trichomes from the surface of tomato leaves.
Figure 4: A smartphone was used to capture an image of the diagnostic kit.
Optimizing a technique that involves the use of cellophane tape for trichome extraction allows the creation of a highly cost-efficient kit. Although recent advances in sophisticated object detection algorithms have utilized complex neural networks, these neural networks typically demand high-speed parallel processing, considerably increasing computational costs. Our approach, which uses cellophane tape for trichome extraction, circumvents the need for such intensive computational resources and enables the use of less computationally demanding algorithms for trichome detection. Thus, this approach reduces the overall cost of the assessment process, making it more economically viable.
The purpose of this diagnostic kit is to meet the technological progress needs in agricultural production areas that lack access to laboratory-level experimental facilities. Therefore, to minimize the number of components required for measurement, the kit is specifically designed to allow for dual manual operation. The measuring paper is held in one hand, whereas the smartphone is operated in the other hand (Figure 4). The statistical model accounts for the variability in the results caused by varying optical conditions among shots. The kit is designed to prioritize user friendliness as a minimum viable product in lean startup methodology for manufacturing and software development, enabling the measurement of trichome density without requiring technical expertise [mu2013lean, Ries2011-xn]. This technology is well suited for practical application in the field because of its simplicity and cost-effectiveness (Figure 5).
Figure 5: Proposed concept of the tomato cultivation support framework.
3.1.2 Proposed concept of the tomato cultivation support framework
This kit is designed to empower farmers to assess the additional fertilization needs of their tomato plants based on trichome density and receive fertilization recommendations. The proposed framework involves a four-step process. First, trichomes are sampled from tomato leaves via adhesive cellophane tape on measurement paper. Second, a smartphone is used to capture an image of the measurement paper with the transferred trichomes. Third, the captured image is analyzed to determine trichome density, which serves as an indicator of the degree of fertilizer stress. Finally, on the basis of the analyzed trichome density, the system provides a fertilization recommendation to the farmer, guiding their decision on whether additional fertilization is necessary.
3.1.3 Overview of the trichome detection image processing pipeline
This study implemented an image processing pipeline for calculating trichome density. Instead of using deep learning models for direct image processing, tabular data analysis was chosen. Although deep learning often requires substantial computational resources, this architecture leverages the inherent simplicity and reduced dimensionality of tabular data, enabling the use of computationally lighter models. This method effectively minimizes computational costs while bolstering quantitative model interpretation, thereby improving the reliability and interpretability of the results for agricultural applications.
To avoid data leakage, an image processing pipeline was developed using images obtained from preliminary experiments. This approach ensured that the model development and evaluation processes remained independent and unbiased.
The image processing pipeline comprises three main stages. The preprocessing stage involves extracting metadata in the exchangeable image file format (EXIF) embedded in the image files, identifying AR markers within the images, applying homography transformations for image alignment, and adjusting for variations in lighting conditions. This preprocessing stage also includes binarizing images to segment trichomes and applying noise removal filtering algorithms to enhance trichome visibility.
In the feature extraction stage, we focused on identifying trichomes in the images, specifically targeting type VI glandular trichomes, owing to their unique morphology. Furthermore, AR markers were utilized to convert pixel counts to relative distances, enabling accurate distance estimation. The final stage, density calculation, involves calculating the trichome density from images captured under various environmental conditions and camera settings, each associated with distinct EXIF data. Algorithm 1 presents an outlined function of the trichome density computation process using our custom-developed measurement kit. The image processing procedures described above were executed via the software specifications outlined in Table 3.
Algorithm 1 Trichome density analysis process
1:$image$: Image file
2:$density$: Nearest neighbor density
3:$expTime$: Exposure time from EXIF data
4:$iso$: ISO speed from EXIF data
5:$openingPx$: Pixel count of measurement paper opening
6:function ReadExifAndLoad($image$)
7:     Read EXIF data from $image$
8:     $expTime,iso\leftarrow$ Extract from EXIF data
9:     $I(x,y)\leftarrow$ Convert $image$ to the array
10:     return $I(x,y)$, $expTime$, $iso$
11:end function
12:function DetectMarkers($I(x,y)$)
13:     Detecting AR markers in $I(x,y)$
14:     $markers\leftarrow$ Extract coordinates from markers
15:     $openingPx\leftarrow$ Calculate the opening pixel count
16:     return $openingPx$, $markers$
17:end function
18:function ApplyHomography($I(x,y)$, $markers$)
19:     Apply homography transform to $I(x,y)$
20:     return $I_{trans}(x,y)$
21:end function
22:function CorrectIllum($I(x,y)$)
23:     Apply illumination correction to $I(x,y)$
24:     return $I_{corr}(x,y)$
25:end function
26:function RemoveNoise($I(x,y)$)
27:     Apply noise removal to $I(x,y)$
28:     return $I_{clean}(x,y)$
29:end function
30:function SegmentTrichomes($I(x,y)$)
31:     Apply watershed segmentation to $I(x,y)$
32:     return $trichomes$
33:end function
34:function CalcDensity($trichomes$)
35:     $density\leftarrow$ Calculate nearest neighbor density
36:     return $density$
37:end function
38:
39:Main Execution
40:$I(x,y)$, $expTime$, $iso\leftarrow\textsc{ReadExifAndLoad}(image)$
41:$openingPx$, $markers\leftarrow\textsc{DetectMarkers}(I(x,y))$
42:$I_{trans}(x,y)\leftarrow\textsc{ApplyHomography}(I(x,y),markers)$
43:$I_{corr}(x,y)\leftarrow\textsc{CorrectIllum}(I_{trans}(x,y))$
44:$I_{clean}(x,y)\leftarrow\textsc{RemoveNoise}(I_{corr}(x,y))$
45:$trichomes\leftarrow\textsc{SegmentTrichomes}(I_{clean}(x,y))$
46:$density\leftarrow\textsc{CalcDensity}(trichomes)$
Table 3: Versions of Libraries and Operating System Used
Operating System/Library Version
Ubuntu 22.04.2 LTS
opencv-contrib-python 4.6.0.66
imbalanced-learn 0.12.0
lightgbm 3.3.5
shap 0.43.0
ExifRead 3.0.0
3.1.4 Image preprocessing: AR marker detection, illumination correction, and noise reduction
The first step of the developed image processing pipeline is to read EXIF data from the image files. This metadata contains valuable information about the environmental conditions and camera settings at the time of image capture. By analyzing aspects such as exposure time and ISO speed ratings, we can gain a comprehensive understanding of the variations in imaging conditions and their potential impacts on trichome density measurements.
To measure the absolute scale and orientation of the captured images, we incorporated AR markers into our image processing pipeline. These markers enabled us to determine the trichome density of the captured images. For AR marker detection, we utilized the ArUco module from the OpenCV library (version 4.6.0.66) [intelcorporationopencvcontribpython].
The AR marker detection algorithm employs a simple, nonrecursive approach, as the development of AR technology was not the primary focus of this study. The process involves image preprocessing, where the captured image is converted to grayscale and then binarized via Otsu’s method to enhance marker visibility under various lighting conditions. We used a predefined "ArUco dictionary", DICT_4X4_250. This dictionary offers a balance between the number of unique markers and detection robustness. Finally, the OpenCV function cv2.aruco.detectMarkers() is used to identify the ArUco markers in the preprocessed image.
The complete algorithm is presented in Algorithm 2.
Algorithm 2 AR marker detection
1:$I$: Input image
2:$C$: Detected marker corners
3:$ids$: Detected marker IDs
4:function ToGray($img$)
5:Convert $img$ to grayscale return grayscale image
6:end function
7:function Binarize($img$)
8:Apply Otsu’s binarization return binary image
9:end function
10:function GetDict($type$)
11:Create ArUco dictionary return dictionary
12:end function
13:function GetParams
14:Create detector parameters return parameters
15:end function
16:function DetectM($img,dict,params$)
17:Detect ArUco markers return corners and IDs
18:end function
19:
20:Main Execution
21:$G\leftarrow\textsc{ToGray}(I)$
22:$B\leftarrow\textsc{Binarize}(G)$
23:$D\leftarrow\textsc{GetDict}(DICT\_4X4\_250)$
24:$P\leftarrow\textsc{GetParams}$
25:$C,ids\leftarrow\textsc{DetectM}(B,D,P)$
Following AR marker detection, we apply homography transformation to correct perspective distortion in images. This distortion is caused by positional and orientational variations in the measurement paper relative to each camera captured. The homography transformation adjusts the images as though the camera and the measurement paper were parallel. This process involves using the coordinates of the four corners of a region on the measurement paper, obtained from detecting AR markers as reference points to estimate a perspective transformation matrix. By applying this transformation, we ensure that the images are standardized for accurate trichome density analysis, regardless of slight variations in camera positioning during image capture.
After applying the homography transformation, we implemented an illumination correction procedure to address nonuniform lighting conditions across the acquired images. This step is essential for maintaining consistency in trichome detection, regardless of ambient lighting variations during image acquisition.
Our approach uses a flat field correction technique [seibert1998medical], which involves several preprocessing steps. Initially, all the images are resized to a standardized dimension of 1000 pixel $\times$ 1000 pixel to ensure uniform processing. We subsequently apply an averaging filter with a kernel size of 50 pixel $\times$ 50 pixel to generate a smoothed version of the image, representing the overall illumination pattern.
The correction is mathematically described by the following equation:
$C=\frac{R\times m}{F}$ (1)
where $C$ denotes the corrected image, $R$ represents the original (precorrection) image, $m$ is the average luminance of the original image, and $F$ is the smoothed image representing the illumination pattern. The details of this correction are presented in Algorithm 3.
Algorithm 3 Illumination correction
1:$R$: Precorrection image
2:$C$: Corrected image
3:function ApplyBlur($image,kw,kh$)
4:     $w\leftarrow$ width of $image$
5:     $h\leftarrow$ height of $image$
6:     $output\leftarrow$ new image of size $w\times h$
7:     for $y\leftarrow 0$ to $h-1$ do
8:         for $x\leftarrow 0$ to $w-1$ do
9:              $sum\leftarrow 0$
10:              $count\leftarrow 0$
11:              for $j\leftarrow-kh/2$ to $kh/2$ do
12:                  for $i\leftarrow-kw/2$ to $kw/2$ do
13:                       $nx\leftarrow x+i$
14:                       $ny\leftarrow y+j$
15:                       if $0\leq nx<w$ and $0\leq ny<h$ then
16:                           $sum\leftarrow sum+image[ny][nx]$
17:                           $count\leftarrow count+1$
18:                       end if
19:                  end for
20:              end for
21:              $output[y][x]\leftarrow sum/count$
22:         end for
23:     end for
24:     return $output$
25:end function
26:function CalculateAverageLuminance($image$)
27:     $sum\leftarrow 0$
28:     $count\leftarrow 0$
29:     for each pixel $p$ in $image$ do
30:         $sum\leftarrow sum+p$
31:         $count\leftarrow count+1$
32:     end for
33:     return $sum/count$
34:end function
35:
36:Main Execution
37:$F\leftarrow\textsc{ApplyBlur}(R,50,50)$
38:$m\leftarrow\textsc{CalculateAverageLuminance}(R)$
39:$C\leftarrow\frac{R\times m}{F}$
[…]
3.1.5 Feature extraction: trichome identification and density analysis
Following noise removal, a segmentation procedure based on the watershed algorithm was implemented to identify individual trichomes. This method detects boundary surfaces that separate contours in the image. The watershed algorithm was chosen because of its suitability for rule-based segmentation on the basis of the distinctive appearance of trichomes, offering a computationally efficient solution. Although deep learning-based segmentation can often achieve high accuracy, it typically requires substantial computational resources, often necessitating the use of devices with advanced parallel processing capabilities, such as graphics processing units, to be integrated into the computer system. In contrast, the watershed algorithm, which is a traditional image-processing technique, generally operates with lower computational costs and can often perform effectively on standard computer hardware, such as central processing units. This characteristic is particularly important for our application, as it ensures that the tomato nutrient diagnostic kit remains cost-effective for farmers, aligning with our goal of developing accessible and practical tools for agriculture.
[…]
In this paper, we utilize the term "trichomes" to describe the contours of detected objects based on the following criteria:
  • A detachable layer must exist, allowing peeling with cellophane tape;
[…]
  • The lower limit of the shape factor within the survey area is defined as the 1st quartile minus 1.5 times the interquartile range;
  • The upper limit of the shape factor within the survey area is defined as the 3rd quartile plus 1.5 times the interquartile range;
  • An object’s shape should more closely resemble a sphere than a prolate ellipsoid, with the circularity exceeding a defined lower limit;
  • The perimeter of the object falls within a specified range, defined by a lower and upper limit;
  • The area of the object lies within a specified range, defined by a lower and upper limit.
The contours detected through segmentation yield centroid points, denoted as in Equation 2. These points exist within a two-dimensional space and are represented via the pixel coordinates of the image. Figure 6 shows the image processing steps performed up to this point, including preprocessing and object detection.
The watershed segmentation process yields a set of detected contours, from which we extract centroid points. These points are represented in a two-dimensional space corresponding to the pixel coordinates of the image. The set of centroids is denoted as $P$, as defined in Equation 2.
$P={p_{i}\in\mathbb{R}^{2}\mid i=1,2,\ldots,N}$ (2)
where $p_{i}$ represents the i-th centroid point in the two-dimensional real space $\mathbb{R}^{2}$, and $N$ is the total number of detected contours. Each point $p_{i}$ can be expressed as $(x_{i},y_{i})$, where $x_{i}$ and $y_{i}$ are the horizontal and vertical pixel coordinates, respectively.
Algorithm 4 Trichome identification
1:$I$: Preprocessed image
2:$P$: Set of trichome centroid points
3:function IsValidTrichome($contour,C_{th},P_{l},P_{h},A_{l},A_{h}$)
4:     $C\leftarrow\textsc{CalculateCircularity}(contour)$
5:     $P\leftarrow\textsc{CalculatePerimeter}(contour)$
6:     $A\leftarrow\textsc{CalculateArea}(contour)$
7:     if $C<C_{th}$ then
8:         return false
9:     end if
10:     if $P<P_{l}$ or $P>P_{h}$ then
11:         return false
12:     end if
13:     if $A<A_{l}$ or $A>A_{h}$ then
14:         return false
15:     end if
16:     return true
17:end function
18:function CalculateAllMetrics($contours$)
19:     Calculate circularity, perimeter, and area for all contours
20:     return $C_{all},P_{all},A_{all}$
21:end function
22:function CalculateCentroidPoint($contour$)
23:     Calculate the centroid point $(x,y)$ of the $contour$
24:     return $(x,y)$
25:end function
26:
27:Main Execution
28:$contours\leftarrow\textsc{WatershedSegmentation}(I)$
29:$P\leftarrow\emptyset$
30:$C_{all},P_{all},A_{all}\leftarrow\textsc{CalculateAllMetrics}(contours)$
31:$C_{th}\leftarrow 1.5\times Q1(C_{all})$
32:$P_{l},P_{h}\leftarrow 1.5\times Q1(P_{all}),1.5\times Q3(P_{all})$
33:$A_{l},A_{h}\leftarrow 1.5\times Q1(A_{all}),1.5\times Q3(A_{all})$
34:for each $contour$ in $contours$ do
35:     if IsValidTrichome($contour,C_{th},P_{l},P_{h},A_{l},A_{h}$then
36:         $centroidPoint\leftarrow\textsc{CalculateCentroidPoint}(contour)$
37:         $P\leftarrow P\cup\{centroidPoint\}$
38:     end if
39:end for
The detailed process of trichome identification, including the criteria and thresholds used for filtering, is presented in Algorithm 4. This algorithm encapsulates the key steps and decision-making process involved in distinguishing type VI glandular trichomes from other objects in the image. Figure 6 shows the image processing steps performed up to this point, including preprocessing and object detection.
Figure 6: Example of image data preprocessing and object detection.
The density was computed for the centroid points denoted as $P$ in Equation 2. We employed the nearest neighbor distance (NND) method, as described in Equation 3 [clark1954distance], to quantify the spatial distribution of these points. The NND method involves measuring the distance between each point and its nearest neighbor within the set.
$NND=\frac{1}{n}\sum_{i=1}^{n}d(p,q)$ (3)
In this equation, $d(p,q)$ represents the Euclidean distance between point $p$ and its closest neighboring point $q$. The NND is then calculated by averaging these distances over the total number of contours $n$ present in the image. Efficiently computing the nearest neighbors of each point is crucial, especially when many points are dealt with. A brute-force approach, where distances between all point pairs are calculated, would have a computational complexity of $O(nd)$, where $n$ is the number of data points and $d$ is the dimensionality. This "big O" notation, $O(nd)$, describes how the execution time of the algorithm increases as the input size ($n,d$) increases[skiena1998algorithm]. In this case, $O(nd)$ means that the execution time increases linearly with the number of points and the dimensionality. Therefore, the brute-force approach can become computationally expensive as the number of points increases. To address this issue, we utilized a k-dimensional tree algorithm, which organizes the points in a space-partitioning data structure that allows for rapid nearest neighbor queries. This algorithm effectively reduces the computational complexity to $O(d\log n)$ [cunningham2022knearest], meaning that the execution time grows linearly with the dimensionality but only logarithmically with the number of points. This logarithmic growth makes the k-d tree algorithm much more efficient than the brute-force approach, especially for large datasets. This algorithm is particularly suitable for our two-dimensional image analysis ($d=2$) with many trichome centroids. The detailed implementation of the NND calculation is presented in Algorithm 5.
Algorithm 5 Density analysis
1:$P$: Set of trichome centroid points
2:$NND$: nearest neighbor distance
3:function CalcNND($P$)
4:     $T\leftarrow\textsc{BuildKDTree}(P)$
5:     $S\leftarrow 0$
6:     for each $p$ in $P$ do
7:         $q\leftarrow\textsc{FindNN}(T,p)$
8:         $d\leftarrow\textsc{EuclidDist}(p,q)$
9:         $S\leftarrow S+d$
10:     end for
11:     return $S/|P|$
12:end function
13:function BuildKDTree($P$)
14:     Construct a k-d tree from points $P$
15:     return tree
16:end function
17:function FindNN($T,p$)
18:     Find the nearest neighbor of $p$ in tree $T$, excluding $p$
19:     return nearest neighbor
20:end function
21:function EuclidDist($p,q$)
22:     $dx\leftarrow p.x-q.x$
23:     $dy\leftarrow p.y-q.y$
24:     return $\sqrt{dx^{2}+dy^{2}}$
25:end function
26:
27:Main Execution
28:$NND\leftarrow\textsc{CalcNND}(P)$
The NND serves as an inverse indicator of contour density, with higher values denoting sparse arrangements and lower values signifying denser groupings. Glandular type VI trichomes are susceptible to detachment from leaf surfaces because of factors such as insect infestations or physical abrasion. However, NND demonstrates superior robustness to such accidental detachment compared with conventional density indices, which are based on the object count per unit area. This robustness arises because NND represents density as a probability distribution, which enables it to maintain reliable estimates even when localized areas experience trichome detachment, provided that unaffected regions persist. Notably, the NND calculation requires the presence of at least two points. The appendix elucidates this concept through a simulation employing point cloud data generated via a Poisson process in two-dimensional space.
3.2 Evaluation of the New Method for Assessing the Nutrient Status of Tomato Plants
3.2.1 Selection of the cultivation method
A well-controlled cultivation experiment is essential for evaluating the proposed method for assessing the nutrient status of tomato plants with testability. This cultivation experiment aimed to measure phenotypes such as trichome density and leaf nitrate ion concentration and fruit yield under ideal conditions for tomato growth, minimizing the variability inherent in soil-based cultivation. Therefore, we opted for hydroponics rather than soil-based cultivation to circumvent the complexities associated with soil properties such as variations in soil composition and structure. These variations can lead to complexities in cultivation experiments, including inconsistent water retention, fertilizer dynamics, and potential for pest and disease outbreaks. This decision allowed for more precise control over nutrient availability and minimized potential confounding factors in our evaluation of fertilizer stress. This approach also enabled us to investigate the relationships between nutrient levels and plant responses more clearly in an ideal growth environment that approximates the optimal growth conditions found in some commercial cultivation systems.
Among various hydroponic techniques, we selected the deep water culture (DWC) method for its simplicity, minimal equipment requirements and ease of implementation in a randomized block design. In DWC, plant roots are fully submerged in nutrient solution within a production tank, which is aerated to sufficiently oxygenate the roots [hoagland1938the, velazquezgonzalez2022a]. This straightforward approach facilitates the relocation and replication of experimental units regardless of location, contributing to increased falsifiability. Moreover, DWC allows easy control of nutrient solution conditions at the individual plant level, making it suitable for randomized block designs.
3.2.2 Tomato growth conditions in hydroponic culture
[…]
Within these buckets, the nutrient mixture was aerated and mixed by introducing air supplied by a blower (YASUNAGA, AP-40P). The air was divided into eight segments, each channeled through silicone tubing and dispersed into the nutrient solution. The use of air stones ensures proper mixing and oxygenation. The buckets were equipped with lids containing center holes for the tomato stems to prevent algae growth in the nutrient solution.
3.2.3 Fertilizer concentration settings
The nutrient solution was prepared via hydroponic fertilizer (OAT Agrio, Fertilizer Series SA), and maintained at three different total dissolved solids (TDS) concentrations to induce various levels of fertilizer stress:
  • High concentration: 1500 ppm
  • Intermediate concentration: 1000 ppm
  • Low concentration: 500 ppm
This study employed an interventional research design to elucidate the causal relationship between fertilizer stress and tomato phenotypes. We set three fertilizer levels to investigate the presence or absence of intervention effects on tomato phenotypes and their dose-response relationships.
[…]
Based on these findings, we set the low-concentration treatment at 500 ppm TDS. The intermediate- and high-concentration treatments were set at twice this level (1000 ppm) and three times this level (1500 ppm), respectively. Furthermore, to confirm that this range of concentrations did not exert excessively adverse effects on tomato growth, we conducted preliminary cultivation experiments. The results revealed no visually identifiable growth defects or plant death at 500 ppm. Additionally, even at 1500 ppm, no apparent growth inhibition due to increased osmotic pressure was observed. These results confirmed the suitability of the selected fertilizer concentration range for the research objectives.
3.2.4 Experimental design and management
The fertilizer plots were arranged in a randomized block design, as shown in Figure 7. During the preliminary experiment, the tomato plants were initially cultivated in the vacant areas depicted in Figure 7. To evaluate the effects of fertilizer level on outcome variables, such as the concentration of nitrate ions in tomato leaves, the density of trichomes, and yield, this randomized block design was employed to minimize the potential impact of spatial variations in environmental factors [fisher1966design].
Water was added to the buckets as needed when the nutrient solution water level decreased. Fertilizer was applied every two weeks, and nutrient solution concentrations were adjusted to maintain the desired TDS levels, which were measured via a TDS meter (EUTECH, PCSTestr35). The fertilizer used in the experiments closely matched the nutrient uptake ratios of the tomatoes, enabling the use of a closed hydroponic system without frequent nutrient solution replacement [ishihara2007effect]. However, during the cultivation period, the system was not maintained as a completely closed system; nutrient solutions were replaced whenever they became cloudy because algae or root debris was present.
During the cultivation period, two plants were lost before the compound leaves were harvested, leaving 20 plants available for inclusion in the study. Furthermore, an additional two plants were lost by the end of the fruit harvesting period, resulting in 18 plants being available for harvest yield assessment.
Figure 7: Experimental layout with three different fertilizer levels (ppm) in a randomized block design.
3.2.5 Compound leaf sampling and trichome transfer
The experiment involved collecting compound leaves from tomato plants. A total of 40 compound leaf samples were collected during two separate harvesting sessions in March and April 2023. The sampling process followed a systematic approach to ensure consistency and minimize potential biases. The samples were processed sequentially in ascending order of the number of tomato plant samples. These sample numbers were randomly assigned to different fertilizer treatment levels as part of the randomized block design, ensuring that the sampling order was not biased by treatment level.
The compound leaves were sampled 15 cm below the apical meristem of each plant. If no suitable compound leaf was available at this position, the plant was temporarily skipped, allowing for further growth. The sampling process continued iteratively until all the plants had been sampled. For trichome density measurements, we selected all leaflets larger than the opening of the measurement paper from each compound leaf.
The number of measurement papers used corresponded to the number of selected leaflets per compound leaf: This approach allows for an accurate assessment of trichome density distribution across each leaflet. Prior to image capture, glandular trichomes were transferred from the leaflet surface to the measurement paper. Each leaflet was carefully affixed to the adhesive side of the cellophane tape (Figure 8) and gently pressed with a finger to ensure complete contact between the leaf surface and the tape. After the leaflet was removed, the cellophane tape with the transferred trichomes was photographed alongside the measurement paper via a smartphone camera.
Figure 8: Trichome transfer process from the leaflet surface to the measurement paper using cellophane tape.
3.2.6 Image acquisition and dataset preparation
The measurement paper was photographed via the native camera app on a smartphone (OPPO, Reno A). This device has a 16-megapixel camera featuring an F1.7 aperture and a CMOS sensor (Sony, IMX398). The camera supports autofocus functionality, allowing for sharp and clear images even under varying conditions. For our experiments, the camera mode was set to automatically adjust the exposure, ISO sensitivity, and focus, leveraging the device’s built-in optimization algorithms.
To ensure a comprehensive dataset that accounts for various usage scenarios, we manually varied the distance and angle between the camera and the measurement paper. The distance from the camera sensor to the measurement paper ranged from 79 to 218 mm, simulating different user holding positions. The horizontal and vertical offset angles varied from -12.52 to 10.62 degrees and -9.04 to 8.13 degrees, respectively. This range of angles mimics slight misalignments that might occur during practical use of the kit. Furthermore, the measurement paper was positioned a sufficient distance from the background to ensure that the background was blurred and not clearly identifiable in the captured images.
The room used for photography had a north-facing window, which eliminated direct sunlight and provided diffuse natural light. The photographic sessions were conducted at different times of the day, with solar elevation angles ranging from -7.95 to 33.42 degrees. The smartphone’s built-in white LED was activated for auxiliary lighting, and the measurement paper was additionally illuminated by oblique light from the ceiling’s fluorescent light (as shown in Figure 4).
The image composition was carefully photographed to ensure that the measurement paper was centered in the frame.
Immediately after the images were captured, a preliminary analysis was conducted to verify the completeness of the image processing pipeline. All the measurement papers were checked to ensure that they could be included in the dataset. In cases where the analysis program encountered exceptions, rephotography was performed to prevent bias due to missing information from the measurement paper during analysis. Consequently, this study focused exclusively on analyzing trichome density from images where AR markers could be detected.
The accuracy of AR marker detection was assumed to impact the results inconsiderably. This assumption was based on the practical consideration that the detection accuracy of AR markers and other 2D codes depend primarily on the library used and has already reached a practical level for applications. Moreover, in the event of an error, the kit user can rectify the problem by performing rephotography.
After the data collection and preprocessing steps, information from all the measurement papers was included in the final dataset, which comprised 6,855 images suitable for analysis. This comprehensive dataset encompasses a wide range of distances, angles, and lighting conditions and ensures that the developed model is trained and evaluated on a diverse set of scenarios, closely mimicking real-world usage of the kit.
3.2.7 Nitrate concentration measurement in tomato leaves
[…]
Figure 9: Workflow from data acquisition to analysis.
3.2.8 Assessment of tomato fruit harvest under experimental conditions
[…]
3.3 Preliminary Data Analysis and Model Preparation
3.3.1 Configuration of the data analysis environment
The analytic environment was built using a Linux-based virtualization technology to ensure consistency and reproducibility of the analytical results across different computer systems, regardless of the specific hardware specifications. The analysis was conducted using Python [pythonsoftwarefoundationpython] version 3.10. Table 3 provides a comprehensive list of the Python library versions and operating system versions utilized in this study. The opencv-contrib-python, imbalanced-learn [scikitlearncontribimbalancedlearn], lightgbm [kguolinandmicrosoftresearchlightgbm], shap [slundbergandsleeshap], and ExifRead [sianareexifread] libraries were utilized for image processing, dataset preprocessing, machine learning model training, model interpretation, and metadata extraction from image data, respectively.
3.3.2 Statistical analysis of nitrate ion concentrations
We employed a series of nonparametric tests to statistically validate the observations of nitrate ion concentrations across different fertilizer levels. First, we used the Kruskal-Wallis test to examine the overall differences in distribution among the three fertilizer levels. Next, we conducted pairwise comparisons via Mann-Whitney U tests to assess the differences between each pair of fertilizer levels. To account for multiple comparisons, we applied the Bonferroni correction to the p values obtained from the Mann-Whitney U tests.
3.3.3 Multicollinearity and variable interaction assessment
Before building the model, we investigated the potential presence of multicollinearity and interaction effects among the predictor variables. Multicollinearity occurs when predictor variables are strongly intercorrelated, which can result in unstable and unreliable regression coefficient estimates. To detect the presence of multicollinearity, we utilized the variance inflation factor (VIF), which measures the extent to which the variance of a regression coefficient estimate increases owing to the correlation among predictor variables. A VIF value surpassing the cutoff point of 10 is commonly considered a strong indication of problematic multicollinearity [zuur2010a]. To enhance the model interpretability and alleviate the consequences of multicollinearity, we removed variables with VIF values surpassing the threshold of 10 from further analyses. Moreover, we explored the possibility of interaction effects among the predictor variables, which arises when the level of a predictor variable influences the impact of another predictor variable on the response variable. To uncover potential interactions between features, we conducted a stratified analysis involving dividing the data into subgroups on the basis of the levels of a suspected interacting variable and assessing relationships between the predictor and response variables within each subgroup. Specifically, we investigated how image resolution might modulate the association between trichome density and nitrate concentration in tomato leaves.
3.4 Analysis of Tomato Fertilization Requirements Based on Leaf Trichome Density
3.4.1 Problem formulation and objectives
The primary objective of this study was to develop a kit that is suitable for practical use in the field. By leveraging existing knowledge on the correlation between leaf nitrogen-related indicators, we aimed to address a crucial question for farmers, which is affected by the fertilizer stress level and trichome density: Do my tomato plants in the field require additional fertilization?
3.4.2 Model design and variable selection
Consequently, we formulated a binary classification problem to determine the need for fertilization on the basis of leaf nitrate ion concentrations. The threshold for the leaf nitrate ion concentration was set as a parameter. Importantly, there is no universally ideal level of fertilizer stress for tomatoes. This level can vary depending on the environmental conditions and specific goals of the cultivator. Recognizing this variability, our study did not set a specific nitrate ion threshold. Instead, we parameterized the threshold for the leaf nitrate ion concentration in our model. This threshold was used to determine whether additional fertilization was necessary: If the concentration exceeded the threshold, no additional fertilization was needed, whereas concentrations below the threshold indicated a need for fertilization. This binary decision was then used as the target variable in our model.
For the explanatory variables in this model, we used the NND, image resolution and exposure time. We applied standard scaling to these explanatory variables to ensure that all variables were on a comparable scale and to improve the model’s performance. This scaling technique transforms the variables so that their mean is zero and their standard deviation is one.
In our cross-validation process, the data were divided into training and test sets. Importantly, the primary focus of this research was to develop a kit suitable for practical field use, with an emphasis on architectural-level improvements. Specifically, a quantitative method for characterizing tomatoes and refinement of measurement methods were selected. Therefore, we did not employ hyperparameter tuning using a validation dataset. Instead, we used default values for the model’s hyperparameters. This decision may lead to underfitting or overfitting due to insufficient control over model complexity, potentially resulting in reduced generalization performance and missed opportunities for theoretical performance analysis and optimization [mohri2018foundations]. However, we deemed these potential drawbacks acceptable, given that our primary objective was to prioritize the development of a practical and accessible kit for field use, rather than achieving the highest possible model performance.
3.4.3 Cross-validation strategy
To evaluate the model’s generalization performance, we employed leave-one-out cross-validation (LOOCV) for each compound leaf. This method was chosen because of the relatively small-scale dataset, derived from 40 compound leaves harvested from 20 tomato plants from the three intervention groups. These compound leaves yielded 300 leaflets, and we analyzed 6,855 images taken of these leaflets. Given this dataset structure, LOOCV was deemed suitable for minimizing bias from specific samples while maximizing the use of our limited but diverse data points. Furthermore, our feature extraction process enables learning with tabular data, maintaining low computational costs despite the iterative nature of LOOCV.
In each iteration of the LOOCV process, a single unique compound leaf was selected. The data derived from this selected compound leaf served as the test set, whereas the data from the remaining compound leaves constituted the training set.
3.4.4 Evaluation methodology
Because the kit users can easily and cost-effectively perform multiple measurements, we evaluated the model’s predictive performance via a method that averages multiple predictions obtained from multiple images into a single prediction. We randomly extracted 1 to 25 rows of data from the test set and input them into the model. This method simulates scenarios where users capture and utilize 1 to 25 images simultaneously for prediction, which allows us to assess the model’s performance under various measurement intensities. This assessment enables us to understand how the model’s performance improves with an increasing number of measurements. To ensure a comprehensive evaluation, we aggregated the predictions and actual observations from all cross-validation iterations: This combined data is used to calculate various performance metrics.
3.4.5 Parameter range selection
To effectively visualize the analysis results, the range of nitrate ion concentrations and the number of images used for evaluation were narrowed to specific ranges or values. This refinement process was guided by practical considerations and domain knowledge.
With respect to the nitrate ion concentration threshold, setting a single universal value is inappropriate because of the variability in agricultural conditions. We could define a relevant range on the basis of domain knowledge. Two key factors influenced this decision. First, nonlinear models such as LightGBM, which we used for modeling, tend to produce unstable predictions near or outside the boundaries of the training data. Second, agricultural workers and related economic entities primarily sustain their livelihoods through crop production revenues. Consequently, it is crucial to avoid excessive fertilizer stress, which could lead to decreased profitability. Considering these points, we decided to exclude the low nitrate concentration region from the threshold setting. Such a scenario is unlikely to be used in a practical application of the kit.
As our kit allows for multiple measurements to be easily and cost-effectively performed, we set the upper limit for the number of images used in the evaluation to 25. This decision balances the need for comprehensive data with practical considerations of time and resource efficiency in real-world applications.
Based on these considerations, we defined our range of interest for nitrate ion concentrations from the second quartile to the third quartile of the distribution obtained from our experiments. For visualizations requiring specific parameter settings, we used the third quartile value as the nitrate ion concentration threshold and set the number of images for evaluation to 25. We refer to this combination of parameter settings as the "representative parameter setting". This approach allowed us to evaluate and visualize the model’s performance within a practical and statistically meaningful range. To ensure comprehensive reporting, performance evaluations using all the parameters, including those outside the defined range of interest, are presented in the appendix.
3.4.6 Data analysis workflow
Algorithm 6 presents a schematic overview of the data analysis workflow. The main processes of model training and evaluation are described in detail below.
Algorithm 6 Tomato fertilization requirement prediction
1:$X$: Explanatory and moderator variables
2:$Y$: Target variables
3:$t$: Threshold for nitrate ion concentration
4:$n$: Measurement intensity
5:$M$: Metadata containing CompoundLeafIndex etc.
6:$metrics$: Performance metrics
7:$S$: SHAP values
8:function SetLabels($Y,t,M$)
9:     $L\leftarrow\emptyset$
10:     for $i\leftarrow 1$ to $|Y|$ do
11:         $label\leftarrow Y[i]<t?0:1$
12:         $M[i][\text{`ClassLabelIndex'}]\leftarrow label$
13:         $L\leftarrow L\cup\{label\}$
14:     end for
15:     return $M,L$
16:end function
17:function ApplySMOTE($X,Y,M$)
18:     Apply SMOTE to $X$ and $Y$ based on $M$
19:     return balanced $X,Y$
20:end function
21:function TrainModel($X,Y$)
22:     Train model on $X,Y$
23:     $Y{pred}\leftarrow\textsc{MakePredictions}(model,X)$
24:     return $Y,Y{pred},model$
25:end function
26:function Evaluate($model,X,Y,n,M{test}$)
27:     MakePrediction
28:     return $Y,Y_{pred}$
29:end function
30:
31:Main Execution
32:$M,L\leftarrow\textsc{SetLabels}(Y,t,M)$
33:$Y{train\_true},Y{train\_pred},Y{test\_true},Y{test\_pred}\leftarrow\emptyset$
34:$S\leftarrow\emptyset$
35:$I\leftarrow\textsc{ExtractCompoundLeafIndex}(M)$
36:for each unique index $i$ in $I$ do
37:     $X{test}\leftarrow$ Select rows from $X$ where $I$ equals $i$
38:     $Y{test}\leftarrow$ Select rows from $Y$ where $I$ equals $i$
39:     $M{test}\leftarrow$ Select rows from $M$ where $I$ equals $i$
40:     $X{train}\leftarrow$ All rows from $X$ except where $I$ equals $i$
41:     $Y{train}\leftarrow$ All rows from $Y$ except where $I$ equals $i$
42:     $M{train}\leftarrow$ All rows from $M$ except where $I$ equals $i$
43:     $X{train},Y{train}\leftarrow\textsc{ApplySMOTE}(X{train},Y{train},M{train})$
44:     $Y{true},Y{pred},model\leftarrow\textsc{TrainModel}(X{train},Y{train})$
45:     $Y{train\_true}\leftarrow Y{train\_true}\cup Y{true}$
46:     $Y{train\_pred}\leftarrow Y{train\_pred}\cup Y{pred}$
47:     $Y{test},Y{pred}\leftarrow\textsc{Evaluate}(model,X{test},Y{test},n,M{test})$
48:     $Y{test\_true}\leftarrow Y{test\_true}\cup Y{test}$
49:     $Y{test\_pred}\leftarrow Y{test\_pred}\cup Y{pred}$
50:     $S\leftarrow S\cup\textsc{ComputeSHAP}$
51:end for
[…]
3.4.7 Feature preprocessing and label setting
The objective of the modeling was to determine the need for additional fertilization by comparing the nitrate ion concentration, estimated from the trichome density in images, with a predefined threshold. Consequently, leaves were labeled 0 (indicating no need for additional fertilization) or 1 (indicating the need for additional fertilization) based on whether the nitrate ion concentration was below or above the threshold, respectively.
3.4.8 Model training
The dataset was partitioned into training and testing sets before the modeling process. The modeling task involved binary classification, with the data being divided based on each compound leaf. To address the imbalance in both the number of images between compound leaves and leaflets and the class labels, the minority class was oversampled via the synthetic minority oversampling technique (SMOTE)[haixiang2017learning, chawla2002smote].
Modeling was performed via a light gradient boosting machine (LightGBM), a decision tree-based approach that employs gradient boosting techniques to combine multiple decision trees for learning purposes [haixiang2017learning]. In LightGBM, each decision tree uses the residuals from the previous tree to improve the overall predictive accuracy of the model iteratively. The regression predictions $z$ were calculated via Equation 4 [ke2017lightgbm].
$z=\sum_{k=1}^{K}f_{k}(x)$ (4)
In Equation (4), $f_{k}(x)$ is defined as the output generated by the $k$-th decision tree. Here, $k$ represents the total number of decision trees, and $x$ corresponds to the input feature vector. The regression prediction $z$ is then passed through the sigmoid function, as shown in Equation 5, to calculate the probability $\hat{y}$.
$\hat{y}=\frac{1}{1+e^{-z}}$ (5)
The sigmoid function in Equation (5) maps the value of $z$ to a probability between 0 and 1, which is then interpreted as the likelihood of the input belonging to class 1. By applying the sigmoid function to the regression predictions $z$, as shown in Equation (5), we simulate a scenario in which farmers are provided with the predicted probabilities of nutrient deficiency, enabling them to make informed decisions about the application of additional fertilization to their fields. The final class label $\hat{Y}$ was determined based on the computed probability $\hat{y}$, according to the decision rule outlined in Equation 6.
$\hat{Y}=\begin{cases}1&\text{if }\hat{y}\geq 0.5\\ 0&\text{if }\hat{y}<0.5\end{cases}$ (6)
To train the classification model, the default parameters of the LightGBM library were employed, including a learning rate of 0.1 and 100 boosting rounds.
3.4.9 Model evaluation and feature importance
The predicted labels $\hat{Y}$ were compared to the true labels, and performance metrics were computed. The evaluation metric used during training was the area under the precision-recall curve (AUC-PR), as defined by Equation 11. To evaluate the performance of the trained model on the test data, several evaluation metrics were employed, including the precision, recall, $F_{1}$, area under the receiver operating characteristic curve (AUC-ROC), AUC-PR, and mean AUC-PR (mPR). These metrics are described by Equations 7-12.
The precision, as defined by Equation 7, is the ratio of correctly predicted positive observations to the total number of predicted positives [hossin2015a, jeni2013facing].
Precision=True PositiveTrue Positive+False $\text{Precision}=\frac{\text{True Positive}}{\text{True Positive}+\text{False Positive}}$ (7)
Equation 8 defines recall as the proportion of accurately predicted positive observations in relation to the total number of true positives. This metric evaluates the model’s ability to identify all relevant positive instances.
Recall=True PositiveTrue Positive+False $\text{Recall}=\frac{\text{True Positive}}{\text{True Positive}+\text{False Negative}}$ (8)
Equation 9 represents the $F_{1}$ score, which is calculated as the harmonic mean of the precision and recall. This metric serves as a balanced measure of both metrics.
$F_{1}Score=\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (9)
Equation 10 represents the AUC-ROC, which is the area under the curve of the true positive rate versus the false positive rate. $t(f)$ denotes the true positive rate at a given false positive rate $f$. This measure effectively summarizes the model’s ability to discriminate between the positive and negative classes across all possible thresholds, providing a single scalar value that reflects the trade-off between sensitivity (true positive rate) and specificity ($1-false-positiverate$).
[…]
Equation 11 represents the AUC-PR, which defines the area under the precision versus recall curve. $p(r)$ represents the precision at a given recall value $r$. This metric provides a single scalar value that summarizes the trade-off between precision and recall across all possible thresholds. The AUC-PR is particularly useful for datasets with imbalanced class distributions, as it provides a more informative measure than does the AUC-ROC [jeni2013facing].
[…]
$\text{mROC}=\frac{1}{m}\sum_{k=1600}^{1900}\text{AUC-ROC}_{k}$ (12)
[…]
$\text{mPR}=\frac{1}{m}\sum_{k=1600}^{1900}\text{AUC-PR}_{k}$ (13)
3.5 Analysis of Trichome Density in Relation to Fertilizer Stress
3.5.1 Research objectives and scope
The primary objective of this research is to develop practical tools for farmers to address agricultural challenges. Specifically, fertilizer stress in tomato plants has been assessed. This assessment is achieved through reverse inference, utilizing a known natural phenomenon: the change in leaf trichome density in response to variations in fertilizer stress. This approach allows us to infer fertilizer stress levels on the basis of observed trichome density patterns. Although our experimental design is optimized for this applied goal, it also provides an opportunity to contribute, albeit secondarily, to the scientific understanding of the relationship between fertilizer stress and trichome density in tomato plants as a natural phenomenon.
3.5.2 Quantitative analysis approach
Leveraging this opportunity, we employed a machine learning approach with a regression model to investigate the quantitative relationship between fertilizer stress and trichome density in tomato plants. This analysis served as a forward inference step, establishing a predictive model for trichome density based on three features: nitrate ion concentration and two moderator variables of image acquisition: image resolution and exposure time.
3.5.3 Model selection and configuration
We implemented the LightGBM algorithm for our regression model, using default hyperparameters. Specifically, the model was configured with an L2 loss objective function, 100 boosting rounds, and a learning rate of 0.1.
3.5.4 Cross-validation and data preprocessing
To rigorously assess the model’s performance and generalizability, we implemented a LOOCV strategy. This approach involved iteratively using each unique leaflet as a test set, while the remaining data served as the training set.
Prior to model training, we applied standard scaling to the feature variables and the target variable. For each LOOCV fold, we trained a LightGBM model on the scaled training data and used it to predict the NND for the test set. The predictions were then inverse-transformed to the original scale for evaluation.
We assessed the model’s performance via several metrics, including the root mean square error ($RMSE$), Coefficient of determination ($R_{2}$) and the Pearson correlation coefficient ($R$).
4 Results
4.1 Nitrate Ion Concentrations in Compound Leaves as Lagging Indicators of Fertilizer Stress
Figure 10 shows the distribution of nitrate ion concentrations in compound leaves located 15 cm from the apical meristem, measured across experimental plots subjected to three different fertilizer concentrations: high, intermediate, and low. Notably, higher fertilizer application rates corresponded to elevated nitrate ion concentrations within the compound leaves.
Statistical analysis via the Kruskal-Wallis test yielded a p value < 0.05, indicating significant differences among the three groups. Subsequent pairwise comparisons via the Mann-Whitney U test, adjusted for multiple comparisons via the Bonferroni correction, revealed statistically significant differences between the high-low and intermediate-low groups ($\text{adjusted}\,p<0.05$). However, no statistically significant difference was observed between the high-intermediate subgroup ($\text{adjusted}\,p=0.52$). This lack of a significant difference between the high- and intermediate-level fertilizer treatments may reflect the trend of diminishing returns. These findings suggest that the maximum yield is greater in the high-intensity fertilizer treatment than in the low-intensity treatment, beyond which additional fertilizer inputs provide increasingly smaller yield gains.
These results suggest that the compound leaf nitrate ion concentration can serve as a reliable indicator of fertilizer stress in tomatoes, particularly for distinguishing between low and high fertilizer levels. However, it is important to acknowledge the variability observed in the relationship between nitrate ion concentration and fertilizer level. The overlapping distributions and nonsignificant differences between the high- and intermediate-level fertilizer treatments are shown in Figure 10. This variability underscores the importance of directly assessing plant nitrate uptake rather than relying solely on fertilizer application rates to accurately evaluate fertilizer stress. Directly measuring plant nutrient uptake can provide a more comprehensive understanding of nutrient availability, accounting for factors beyond fertilizer application that may influence nutrient absorption.
Figure 10: Distribution of nitrate concentrations in compound leaves at different fertilizer levels.
4.2 Leaf Nitrate Concentration as a Leading Indicator of Fruit Yield
Figure 11 shows a positive correlation between fruit yield and nitrate concentration in compound leaves. This finding indicates that the nitrate ion concentration in compound leaves can serve as an indicator of fruit yield. Generally, crop yield tends to increase with increased fertilizer absorption; however, fertilizer application beyond a certain threshold can reduce yields [andersen1999relationships, locascio1997nitrogen, coltman1988yields]. The findings of this study indicate that fertilizer application in the experimental field fell below this critical threshold. These findings suggest that the tomatoes experienced stress due to insufficient fertilization.
As detailed in the Parameter Range Selection section, various factors were considered when determining an appropriate nitrate concentration threshold for visualization purposes. Based on these considerations, we employed the third quartile value of 1,783 ppm, derived from the measured leaf nitrate concentrations, to illustrate the model’s performance in distinguishing between different yield potentials.
Figure 11: Relationship between the nitrate ion concentration in compound leaves and yield.
[…]
Figure 12: Cumulative fruit yield in the experimental setup.
4.3 Handling of Multicollinearity and Variable Interaction in Data Analysis
The statistical model necessitates the exclusion of ISO speed ratings as a variable because of the manifestation of multicollinearity. A Farrar-Glauber test was conducted to assess multicollinearity, revealing a significant presence ($p<0.001$) [farrar1967multicollinearity]. This result indicates a statistically significant linear dependence among the explanatory variables. Table 4 presents the variance inflation factor (VIF) for each variable. Notably, the ISO speed ratings exhibit an exceptionally high VIF of 47.80, substantially exceeding the conventional threshold of 10 and thus confirming severe multicollinearity.
Table 4: Examination of Multicollinearity through the VIF
Feature VIF
Nearest Neighbor Distance 43.60
Resolution 11.03
Exposure Time 4.98
ISO Speed Ratings 47.80
The ISO speed ratings may inadvertently function as a mediating variable between trichome density and fertilizer stress, attributable to the camera’s automatic exposure (AE) function. The AE system responds to the overall image brightness. This phenomenon is considerably influenced by trichomes appearing as white points due to the reflection of light from the LED of the smartphone. Variations in trichome density, induced by fertilizer stress, alter image brightness, prompting automatic adjustments in ISO sensitivity. Consequently, the ISO speed rating becomes an indirect indicator of trichome density and, by extension, fertilizer stress.
This intricate relationship positions ISO sensitivity as a mediator variable, not through direct measurement of trichomes or fertilizer stress but by reflecting the camera’s response to image characteristics indicative of these factors. The high VIF value for ISO speed ratings can be attributed to this complex interaction, where the camera’s real-time exposure adjustment algorithm effectively renders ISO sensitivity dependent on trichome density, which is, in turn, influenced by fertilizer stress. This unintended mediation effect introduces complications into the dataset: ISO sensitivity becomes correlated with the variable of interest (trichome density) and the target variable (fertilizer stress).
Notably, although multicollinearity may not significantly affect the overall predictive performance of the model as evaluated by cross-validation, it can substantially impact the interpretation of individual variable importance. In the presence of multicollinearity, estimated coefficients for individual predictors may become unstable and unreliable. This instability can lead to biased assessments of variable importance, potentially resulting in overestimation or underestimation of the importance of certain variables in the model.
In the context of this study, where both the generalization performance and model interpretability through cross-validation are paramount, it is imperative to exclude the ISO speed ratings as a variable because of their demonstrated multicollinearity. This exclusion enhances the robustness and interpretability of the statistical model, ensuring more reliable insights into the relationships between trichome density, fertilizer stress, and other relevant variables.
The resolution value exhibits interactions and a shift in the intercept, necessitating its inclusion as a variable within the statistical model. An investigation of the effect of resolution on the relationship between trichome density and nitrate concentration in compound leaves reveals varying coefficients and intercepts in the regression equation. These changes in the coefficients with respect to the resolution are demonstrated in Table 5 and Figure 13. The resolution inherently varies because of factors such as camera performance and the distance between the camera sensor and the measurement paper. Consequently, in the dataset used in this study, the resolution is an indicator of the distance between the camera and the measurement paper. Thus, to avoid omitted variable bias, which could lead to misinterpretation of the effects of one variable as those of another, it is crucial to include resolution as a variable in the model. Furthermore, the varying coefficients demonstrate the advantages of using higher-resolution images for trichome detection in computer vision applications. The results show that increased resolution leads to improved detection accuracy, highlighting the importance of holding the measurement paper close to the camera and using a high-resolution image sensor for such tasks.
Table 5: Interaction when Stratified Based on Resolution
Percentile Resolution (pixel) Equation
Higher than 85th $>$ $1.08\times 10^{6}$ $y=3.06\times 10^{-2}x+170$
15th to 85th $1.08\times 10^{6}-5.56\times 10^{5}$ $y=1.90\times 10^{-2}x+205$
Lower than 15th $<$ $5.56\times 10^{5}$ $y=1.28\times 10^{-2}x+254$
Figure 13: Effect of resolution on the relationship between trichome density and nitrate concentration in leaves.
Table 6: VIF Value after Dimensionality Reduction
Feature VIF
Nearest Neighbor Distance 8.63
Resolution 6.49
Exposure Time 3.10
Table 6 displays the VIF values for all the variables except for the ISO speed ratings. After excluding ISO speed ratings, no variables with a VIF exceeding 10 were identified, indicating that multicollinearity was no longer a concern. To improve the interpretability of the model, ISO speed ratings were excluded as a variable in subsequent analyses because of their strong correlation with other variables and their potential to obscure the individual effects of each predictor. Instead, the modeling process included variables such as NND, resolution, and exposure time, which were deemed to have a more independent impact on the response variable.
4.4 Performance of the Proposed Method
Figure 14 shows the learning curve of the LightGBM model over 100 boosting rounds with default settings. A gradual increase is observed in the AUC-PR values for the training sets after each boosting round. This improvement indicates that the model has been adequately trained.
Figure 14: Learning curve of the LightGBM model.
Table 7 shows the predictive performance of the classification model at the different nitrate concentration thresholds. With respect to the trade-off between precision and recall, the model achieves highly balanced values.
Table 7: Predictive Performance in Terms of the Need for Additional Fertilizer using LOOCV at Various Nitrate Ion (ppm) Thresholds when Using the 25 Image Features
Evaluation Metrics 1600 1750 1900
Precision 0.68 0.76 0.79
Recall 0.66 0.76 0.79
F1 Score 0.67 0.76 0.79
PR AUC 0.76 0.84 0.87
ROC AUC 0.64 0.68 0.58
Table 8: Comparison of Predictive Performance between the Proposed Model and the Baseline Model in a Representative Parameter Setting
Model Precision Recall F1 Score ROC AUC PR AUC
Proposed 0.79 0.78 0.79 0.81 0.87
Baseline 0.62 0.49 0.55 0.50 0.72
For subsequent analyses, we defined a representative scenario using a nitrate concentration threshold of 1,783 ppm and 25 images captured from a single compound leaf. Using this representative scenario, predictions were made, allowing for an accurate assessment of the need for additional fertilization. The predictive performance, as evaluated by the mROC, is 0.64, and, as evaluated by the mPR, it is 0.82 despite variations in the measurement data caused by differences in optical conditions.
Figure 15 shows a confusion matrix depicting true positives, false positives, true negatives, and false negatives. The predictions do not show substantial bias toward false positives or false negatives, indicating a balanced performance in identifying both classes correctly. This lack of bias is also evident from the balanced results presented in Table 7, which includes the F1 score and AUC-PR and AUC-ROC values, as well as the PR curve in Figure 16 and the ROC curve in Figure 17.
Table 8 compares the performance of the proposed model and the baseline model. The baseline model uses a simple majority class predictor. This comparison utilizes the metrics defined in Equations 7-11, using the representative parameter setting. The table demonstrates the improved performance of the proposed model across all evaluated metrics.
Figure 15: Confusion matrix for classification outcomes during LOOCV in a representative parameter setting.
Figure 16: Precision-recall (PR) curve analysis of the model during LOOCV in a representative parameter setting.
Figure 17: Receiver operating characteristic (ROC) curve analysis of the model during LOOCV in a representative parameter setting.
Figure 18 illustrates the relationship between the number of input images and the model’s predictive performance. Using a single image yields a mean AUC-PR (mPR) of 0.80, and increasing the number of input images leads to better predictive performance.
Figure 18: Model prediction performance when various thresholds and numbers of images are set.
Figure 19 illustrates the global feature contributions based on SHAP values, providing insights into the overall model interpretation. NND has a high mean absolute SHAP value, suggesting its significant contribution to the model’s predictions across the dataset.
Figure 19: Feature contribution during LOOCV with a representative model setting.
4.5 Effect of Fertilizer on Trichome Density
To quantitatively evaluate the relationship between trichome density on tomato leaves and nitrogen nutritional status, we conducted modeling via NND. Figure 20 presents the results of an NND prediction model employing nitrate ion concentration, image resolution, and exposure time as explanatory variables. The coefficient of determination ($R^{2}$) for this model is 0.63, indicating that the combination of explanatory variables (nitrate ion concentration) and moderating variables (image resolution, exposure time) can account for approximately 62.48% of the variation in trichome density (represented by NND) of leaflets obtained from fertilizer intervention experiments.
Notably, this study was designed primarily to focus on the development of image-processing technology rather than being a rigorous hypothesis-testing study for interpreting natural phenomena in complex system. Typically, a hypothesis-based approach starts with a specific hypothesis or theory, and experiments are designed to test the validity of this hypothesis in a deductive manner [glass2008brief]. In contrast, while a simple hypothesis existed for conducting experimental work as shown in Figure 3, this study primarily adopted a data-driven approach, where hypotheses and models are generated inductively based on patterns and relationships found within the dataset.
Therefore, it is appropriate to interpret the results of this study as an analysis stemming from a data-driven approach. However, the relationship between trichome density on tomato leaves and nitrogen nutritional status has been quantitatively verified in previous studies through hypothesis-based research [hoffland2000nitrogen]. As this study aims to capture general trends rather than subtle differences, the need for strict hypothesis-based experimentation is relatively low. Consequently, by adopting an inductive research method, the likelihood of fatal bias is considered low, and the data-driven approach of this study supports the possibility that fertilizer stress may affect trichome density, particularly in young leaves.
Nevertheless, the existence of unexplained variation (approximately 37.52%) in the model suggests the need for further investigation of additional factors and model improvements in future research.
Figure 20: Modeling of NND using nitrate concentration, resolution, and exposure time.
5 Discussion
Tomatoes pose a challenge for detecting fertilizer stress through noncontact optical measurements of young leaf chlorophyll. This challenge arises from the nonlinear relationship between optical measurements and leaf nitrogen content, which distinguishes tomatoes from other crops [padilla2018proximal, gianquinto2011a]. To address this problem, our study focused on trichomes, which are densely packed hair-like structures on leaf surfaces that have been less explored in previous research. We investigated the feasibility of detecting fertilizer stress by developing a device that utilizes trichome density as a primary predictor. Our findings suggest the possibility of overcoming the limitations of conventional noncontact optical methods in assessing tomato fertilizer conditions. Our approach opens new avenues for developing diagnostic devices that consider leaf surface structure. However, the diagnostic kits developed in this study have some limitations regarding their performance. We validated the performance of the kit through experiments using a well-designed setup that simulated different levels of fertilizer stress commonly found in agricultural production conditions. In addition, the tomato cultivation method used in the present study involved cost-effective hydroponic cultivation via plastic buckets. The yields achieved were comparable to those obtained in commercial cultivation. This finding suggests that the cultivation method employed in this study closely simulated agricultural production conditions. However, it is uncertain whether all farmers will be able to establish optimal growing conditions for tomatoes to achieve higher yields, considering the financial costs involved in acquiring the necessary equipment. Consequently, environmental stresses beyond fertilizer stress could be encountered, which could decrease the predictive performance of our novel method under these circumstances.
Traditionally, farmers have relied on visual inspection of leaf color, size, texture and overall plant health to assess the nutrient status of plants. However, the accuracy of this method highly depends on the farmers’ experience and motivation, resulting in considerable variations among individuals. Consequently, standardizing the assessment of nutrient status becomes problematic, making it difficult to compare and evaluate the effectiveness of different approaches. The absence of standardization in nutrient status assessment hinders highly advanced evaluation of the usefulness of diagnostic kits for a certain farmer solely on the basis of a comparison of their predictive accuracy. Thus, further demonstrations and validations employing more lagging indicators, such as financial metrics, are necessary to assess the practical applicability of these diagnostic kits in real-world agricultural settings.
Although various nutrient diagnostic kits are commercially available for assessing plant nutrient status, their high cost and potential difficulties in application to delicate plant tissues, such as young tomato leaves, limit their widespread adoption in practical farming settings. To overcome these challenges, we propose a novel, simple, and cost-effective method for assessing nutrient status in tomato plants, with a specific focus on young leaves, which are typically relatively sensitive to nutrient deficiencies and play a crucial role in plant growth and development.
In addition to tomatoes, various plant species, including agriculturally important crops such as watermelon (Citrullus lanatus), egg plants (Solanum melongena), tobacco (Nicotiana tabacum), soybean (Glycine max), and cannabis (Cannabis sativa), are known to have abundant trichomes. The proposed trichome-based method for assessing nutrient status has application potential in other plant species with dense trichomes near their growing points, expanding its scope and potentially improving real-time nutrient management practices in a wide range of crops, thereby enhancing the efficiency and precision of fertilizer application. By optimizing fertilizer application and reducing waste, this method could lead to increased crop yields, lower production costs, and reduced environmental impact, ultimately contributing to more sustainable agricultural practices. Further research is needed to validate the effectiveness of this method in various crop species and investigate the potential integration of this approach with other precision agriculture technologies, such as remote sensing, for more comprehensive and efficient nutrient management strategies.
6 Conclusions
This study focused on developing a novel method for detecting fertilizer stress in tomato plants by measuring the density of trichomes on their leaves. The study also included setting up and evaluating the experimental environment necessary for conducting these assessments. The developed diagnostic kit enables noninvasive and early detection of fertilizer stress in plants, obviating the need for customized sensing device design and permitting the use of a universal diagnostic kit at a lower cost than traditional devices while accurately assessing the nutrient status of tomato plants. The findings of this study present a novel approach to designing diagnostic devices for detecting fertilizer stress in plants by considering the structure of the plant surface.
Appendix
Appendix A Simulation Demonstrating the Robustness of the Nearest Neighbor Distance
This appendix presents a simulation demonstrating the robustness of the nearest neighbor distance (NND) as a density metric compared with the conventional point count per unit area.
The simulation involves generating a set of points within a defined area. Damage is simulated by removing a portion of these points and then comparing the changes in the NND and the point count. To mimic the random spatial distribution of trichomes on a leaf surface, we employed a Poisson point process for point generation. The Poisson process is often used to model random events occurring independently in a given space or time, making it suitable for simulating the distribution of trichomes, which are expected to emerge randomly on the leaf surface.
The detailed steps are outlined in Algorithm 7.
Algorithm 7 Trichome distribution simulation
1:$\lambda$: Expected number of events (trichomes)
2:$A$: Total area of the region
3:$S$: Random seed value
4:$P$: Set of points representing trichome locations
5:$P^{\prime}$: Set of points after simulated damage
6:function SetSeed($S$)
7:Set a random seed to $S$
8:end function
9:function PoissonRandom($\lambda$)
10:Generate random numbers from the Poisson distribution return number of points
11:end function
12:function UniformRandom($a,b$)
13:Generate random numbers from a uniform distribution return random number between $a$ and $b$
14:end function
15:function GeneratePoints($N$)
16:Generate $N$ random points in the area return set of points $P$
17:end function
18:function SimulateDamage($P$)
19:Remove points with $x>500$ return set of remaining points $P^{\prime}$
20:end function
21:
22:Main Execution
23:SetSeed($S$)
24:$N\leftarrow\textsc{PoissonRandom}(\lambda)$
25:$P\leftarrow\textsc{GeneratePoints}(N)$
26:$P^{\prime}\leftarrow\textsc{SimulateDamage}(P)$
Figure 21 shows an example of the point distribution before and after simulated damage. The points were randomly generated via a Poisson process within a 1000 $\times$ 1000 area. The damage was simulated by removing points within the right half of the area, as indicated by the "damage boundary" line. Figure 22 shows the histograms of NNDs before and after simulated damage.
Figure 21: Point distribution before and after simulated damage.
The simulation results demonstrate that NND is more robust to point removal than the conventional density metric based on the point count per unit area. Figure 23 compares the NND and point count before and after simulated damage. Although both metrics decreased after damage, the decrease in NND was smaller than that in the point count.
Figure 22: Histogram of nearest neighbor distances.
To statistically validate this observation, we performed a Wilcoxon signed-rank test to compare the absolute change rates of the NND and the point count after damage, as shown in Figure 24. The test revealed a statistically significant difference between the two metrics ($p<0.001$), supporting the hypothesis that NND is less affected by localized point loss than the point count per unit area is. This finding highlights the advantage of using NND as a density metric in scenarios where localized point loss might occur, such as trichome detachment from leaf surfaces.
Figure 23: Comparison of NND and point count after simulated damage.
Figure 24: Comparison of NND and point count absolute change rates after simulated damage.
Appendix B Comprehensive Model Performance across All Nitrate Ion Concentration Thresholds
Although Table 7 in the main text focused on the nitrate ion concentration range of 1600-1900 ppm, this section provides a more comprehensive evaluation of the model’s performance across the entire range of nitrate ion concentrations observed in our study. Table 9 presents the detailed results for the proposed and baseline models.
For this comprehensive analysis, we divided the full range of observed nitrate ion concentrations into 10 equal intervals: Each interval serves as a threshold for model evaluation. This approach allows for a thorough assessment of model performance across all possible thresholds within our dataset.
Table 9 illustrates the predictive performance of the proposed and baseline models using LOOCV at these various nitrate ion concentration thresholds. Both models utilized features extracted from 25 images per compound leaf. For clarity, the table presents the results for the proposed model as the main values, with the corresponding results for the baseline model shown in parentheses below.
Table 9: Predictive Performance Comparison between Proposed and Baseline Models using LOOCV at Various Nitrate Ion Concentration Thresholds
Threshold Precision Recall F1 Score PR AUC ROC AUC
600 0.15 0.04 0.06 0.13 0.51
(0.08) (0.49) (0.13) (0.30) (0.50)
777 0.12 0.11 0.11 0.18 0.48
(0.15) (0.49) (0.23) (0.36) (0.50)
955 0.39 0.42 0.40 0.49 0.57
(0.30) (0.49) (0.37) (0.47) (0.50)
1133 0.60 0.66 0.63 0.69 0.70
(0.38) (0.49) (0.43) (0.53) (0.50)
1311 0.61 0.66 0.63 0.71 0.67
(0.42) (0.49) (0.46) (0.57) (0.50)
1488 0.61 0.61 0.61 0.70 0.65
(0.45) (0.49) (0.47) (0.59) (0.50)
1666 0.71 0.68 0.69 0.79 0.65
(0.57) (0.49) (0.53) (0.68) (0.50)
1844 0.74 0.72 0.73 0.82 0.63
(0.78) (0.49) (0.60) (0.83) (0.50)
2200 0.95 1.00 0.97 0.97 0.50
(0.95) (0.49) (0.65) (0.96) (0.50)
[…]

### v3

@@
 Abstract
-Early detection of fertilizer-induced stress in tomato plants is crucial for timely crop management interventions and yield optimization.
-Conventional optical methods detect fertilizer stress in young leaves with difficulty.
-This study proposes a novel, noninvasive technique for quantifying the density of trichomes-elongated hair-like structures found on plant surfaces-on young leaves using a smartphone.
-This method exhibits superior detection latency, enabling earlier and more accurate identification of fertilizer stress in tomato plants.
-Our approach combines augmented reality technology and image processing algorithms to analyze smartphone images of a specialized measurement paper.
-This measurement paper is applied to a tomato leaf to transfer trichomes onto its adhesive surface.
-The captured images are then processed through a pipeline involving region of interest extraction, perspective transformation, and illumination correction.
-Trichome detection and spatial distribution analysis of these preprocessed images yield a robust density metric.
-We validated our method through experiments on hydroponically grown tomatoes under varying fertilizer concentrations.
-Using leave-one-out cross-validation (LOOCV), our model achieves a mean area under the precision-recall curve of 0.824 and a receiver operating characteristic curve of 0.641 for predicting additional fertilization needs.
-Based on LOOCV, quantitative analysis revealed a strong relationship between trichome density and explanatory variables, including nitrate ion concentration, explaining 62.48% of the variation ($R^{2}=0.625$).
-The predicted and actual trichome densities were strongly correlated ($r=0.794$).
-This straightforward and cost-effective method overcomes the limitations of traditional techniques, demonstrating the potential of using smartphones for practical plant nutrition diagnosis.
+Early detection of fertilizer-induced stress in tomato plants is crucial for optimizing crop yield through timely management interventions.
+While conventional optical methods struggle to detect fertilizer stress in young leaves, these leaves contain valuable diagnostic information through their microscopic hair-like structures, particularly trichomes, which existing approaches have overlooked.
+This study introduces a smartphone-based noninvasive technique that leverages mobile computing and digital imaging capabilities to quantify trichome density on young leaves with superior detection latency.
+Our method uniquely combines augmented reality technology with image processing algorithms to analyze trichomes transferred onto specialized measurement paper.
+A robust automated pipeline processes these images through region extraction, perspective transformation, and illumination correction to precisely quantify trichome density.
+Validation experiments on hydroponically grown tomatoes under varying fertilizer conditions demonstrated the method’s effectiveness.
+Leave-one-out cross-validation revealed strong predictive performance with the area under the precision-recall curve (PR-AUC: 0.82) and area under the receiver operating characteristic curve (ROC-AUC: 0.64), while the predicted and observed trichome densities exhibited high correlation ($r=0.79$).
+This innovative approach transforms smartphones into precise diagnostic tools for plant nutrition assessment, offering a practical, cost-effective solution for precision agriculture.
 Keywords: Tomato, tomato leaf, fertilizer stress, trichome, smartphone, computer vision, machine learning, precision agriculture, digital agriculture
@@
 Tomato (Solanum lycopersicum), a vegetable that is widely cultivated worldwide, is influenced by various environmental factors that affect its production and quality.
-The nutrient status of plants, one of the most crucial factors influencing tomato yield, is an important consideration in tomato production.
+The nutrient status of plants, which is one of the most crucial factors influencing tomato yield, is an important consideration in tomato production.
 Therefore, accurately assessing plant nutrient status is essential for maintaining high tomato yields.
-Techniques have been developed to analyze the constituents of fertilizers in soil, media, and nutrient solutions, aiming to estimate the amount of nutrients absorbed by plants [sonneveld1990estimating, chen2015crop] .
+The assessment of plant nutrient status has evolved from invasive to noninvasive approaches.
+Initially, techniques were developed to analyze the constituents of fertilizers in soil, media, and nutrient solutions, aiming to estimate the amount of nutrients absorbed by plants [sonneveld1990estimating, chen2015crop].
 Additionally, researchers have developed invasive methods to analyze plant nutrients [hochmuth1994efficiency, andersen1999relationships].
-Moreover, noninvasive techniques such as visual inspection and sensor-based monitoring are available for assessing plant traits.
-Morphological changes can serve as valuable indicators for nutritional analysis, particularly in the context of environmental stressors such as drought and nutrient deficiencies.
-With respect to drought stress, Klepper et al.
+More recently, noninvasive techniques, including visual inspection and sensor-based monitoring, have become available for assessing plant traits, offering the advantage of continuous monitoring without damaging the plant.
+Among these noninvasive techniques, morphological observation has proven particularly valuable for assessing plant responses to environmental stressors, serving as valuable outcome measures of plant stress responses.
+Studies have demonstrated strong relationships between morphological changes and various stress conditions, particularly in the context of environmental stressors such as drought and nutrient deficiencies.
+Under drought stress, Klepper et al.
 conducted an observational study and reported that cotton stem diameter, leaf water potential, and relative leaf water content are closely related to the net radiation above the plant canopy [klepper1971stem].
@@
 optically analyzed the wilting of tomato leaves and changes in stem diameter in an observational study and demonstrated a correlation between these parameters [wakamori2019optical].
-In terms of nutrient stress, Padilla et al.
+In terms of fertilizer stress, Padilla et al.
 provided a comprehensive review of techniques for monitoring the nitrogen status of vegetable crops via optical methods [padilla2018proximal].
@@
 reported a causal relationship between the fertilization rate and tomato leaf area in an interventional study [tei2002critical].
-These findings emphasize that environmental stressors alter the vegetative and reproductive growth patterns observed in plants.
-The ability to detect and examine these modifications allows for a more sophisticated understanding of the nutrient levels of plants and their responses to varying environmental conditions.
-Leaf area and chlorophyll content, which are related to leaf color and spectrum, are commonly used as indicators of the adaptive response of plants to fertilizer stress.
+These findings collectively demonstrate how environmental stressors manifest in observable changes in plant morphology.
+The ability to detect and examine these modifications allows for a more sophisticated understanding of plant responses to varying environmental conditions.
+Building on these morphological indicators, optical measurement techniques have been developed to provide quantitative assessments of plant stress responses.
+These methods focus particularly on leaf characteristics such as area and chlorophyll content, which are commonly used as indicators of the adaptive response of plants to fertilizer stress and can be measured through leaf color and spectral analysis.
 dela Torre et al.
@@
 These methods have led to the development of a diagnostic framework based on the relationships among fertilizer application, phenotypic expression, and yield outcomes.
-Color and spectral data analysis of plant canopies enables predictions of plant health and productivity.
-These predictions utilize wide-area imagery from satellites, aircraft, and drones combined with data collected from in-field sampling stations.
-The aforementioned optical observation techniques, which utilize reflected light from the plant body, effectively address the need for noninvasive and digital monitoring of phenotypes [padilla2018proximal].
-Furthermore, techniques for detecting fertilizer stress in tomatoes have been extensively researched.
-Tables 1 and 2 present a comprehensive comparison of various noninvasive methods for early detection of fertilizer stress in tomato plants.
+Color and spectral data analysis of plant canopies makes it possible to predict plant health and productivity.
+Current optical observation techniques utilize various platforms, from handheld devices to satellites and drones, combined with data collected from in-field sampling stations.
+These techniques, which utilize reflected light from the plant body, enable noninvasive and digital monitoring of plant phenotypes [padilla2018proximal].
+The detection of fertilizer stress in tomatoes has been extensively researched via these approaches.
+Tables 1 and 2 present a comprehensive comparison of various noninvasive methods for the early detection of fertilizer stress in tomato plants.
 We focused on studies that employed interventional experiments with varying fertilizer levels and used noninvasive measurement techniques on tomatoes.
 Priority was given to methods that utilized digital measurement tools and enabled measurements as close to the growth point as possible for early stress detection.
-As shown in Table 1, although these methods may appear equally effective at first glance, the complexity of data processing (visualization, modeling, and interpretation) after measurement must be considered, which varies considerably depending on the selected phenotypic indicator and sensor type.
-Moreover, in terms of detection latency, which refers to the delay between the onset of fertilizer stress and its detection, evaluating phenotypic indicators based on the metabolic mechanisms of tomato plants is crucial.
-Tomatoes exhibit apical growth, where cell differentiation occurs at the top of the plant, and differentiated organs move downward as the plant grows.
-Considering this mechanism, measuring phenotypic indicators near the shoot apical meristem could lead to earlier and more accurate detection of fertilizer stress.
+However, the effectiveness of these methods varies considerably in their practical application.
+The selected phenotypic indicator and sensor type influence multiple aspects of measurement.
+These aspects include the complexity of data processing for visualization, modeling, and interpretation; the necessary prerequisites during measurement; and practical constraints such as measurement time and operational complexity.
+A critical consideration in tomato nutrient assessment is detection latency, the delay between stress onset and detection.
+Evaluating phenotypic indicators based on the metabolic mechanisms of tomato plants is crucial, as tomatoes exhibit apical growth, where cell differentiation occurs at the top of the plant and differentiated organs move downward as the plant grows.
+Considering this mechanism, measuring phenotypic indicators near the shoot apical meristem could theoretically provide earlier and more accurate detection of fertilizer stress.
+The measurement of young leaves should therefore be recognized as a critical nonfunctional requirement for early stress detection.
 However, as shown in Table 2, existing studies primarily utilize "fully expanded leaves" or even the entire plant as a phenotypic indicator for chlorophyll meters, visible-near infrared spectroscopy, fluorescence spectroscopy, stem diameter measurement, and leaf area measurement.
-A "fully expanded leaf" is generally considered the 3rd to 5th compound leaf from the apex, excluding the apical meristem and leaf primordia, and has reached its final size and shape.
-However, a universally accepted definition of a fully expanded leaf does not exist, and its specific characteristics are as follows: The precise distance from the apical meristem may vary depending on factors such as plant species, cultivar, growth stage, and environmental conditions.
-Therefore, the context of each study must be considered when interpreting the term "fully expanded leaf." Consequently, studies in the literature identifying the phenomenon of fertilizer stress are lacking.
-This scarcity is especially true in the assessment of young tomato leaves, with the definition of "young leaves" itself being subject to variation.
+A "fully expanded leaf" is typically considered the 3rd to 5th compound leaf from the apex, excluding the apical meristem and leaf primordia, and has reached its final size and shape.
+Nevertheless, there is no universally accepted definition of a "fully expanded leaf," and its specific characteristics, including the precise distance from the apical meristem, may vary depending on factors such as plant species, cultivar, growth stage, and environmental conditions.
+Therefore, it is essential to consider the context of each study when interpreting the term "fully expanded leaf".
+This approach may delay stress detection, highlighting a significant gap in early stress detection capabilities.
+There is a lack of studies identifying the phenomenon of fertilizer stress, particularly in the assessment of young tomato leaves, with the definition of "young leaves" itself being subject to variation.
+Although young leaf measurement is a critical nonfunctional requirement, current methodologies have not adequately addressed this need.
+To address these limitations and develop more effective methods for early stress detection, a systematic review of existing measurement approaches and their limitations is necessary, particularly with a focus on the challenges specific to tomato plants.
 Table 1: Comparison of Noninvasive Methods for Detection of Fertilizer Stress in Tomato Plants
 Method Phenotypic Indicator Instrumentation Conditions to Avoid Data Complexity Proposed Method Trichome density Digital camera Direct sunlight 2D image array Leaf Chlorophyll Content [padilla2015threshold] Chlorophyll a and b content Chlorophyll meter Direct sunlight Scalar value Visible-Near Infrared Spectroscopy [padilla2015threshold] Multiple biochemical components VIS-NIR spectrometer Direct sunlight 1D spectral array Fluorescence Spectroscopy [bahadur2015gas] Multiple fluorescent compounds Fluorescence spectrometer Bright light 1D spectral array Leaf Area [xiao2022relationship] Leaf area Ruler Withering Scalar value Stem Diameter [olagunju2024effect] Stem diameter Caliper Withering Scalar value
-* This table includes studies that met the following criteria: noninvasive measurement, tested on tomatoes, interventional experiments with varying fertilizer levels.
+* This table includes studies that met the following criteria: noninvasive measurement, tested on tomatoes, and interventional experiments with varying fertilizer levels.
 Priority was given to methods with digital measurements and those enabling measurements near the growth point for early stress detection.
@@
 […]
-a Measurement location not specified in the original study.
+a Measurement location was not specified in the original study.
 b Sensitivity to fertilizer stress was not quantitatively assessed in the original study; however, a correlation was visually observed in the presented figures.
 c Specific sensitivity indicator was not provided in the original study.
-d mPR and mROC represent the mean area under the precision-recall curve and the mean area under the receiver operating characteristic curve, respectively (see equations 10 and 11).
+d mPR and mROC represent the mean area under the precision-recall curve and the mean area under the receiver operating characteristic curve, respectively (see equations 12 and 13).
 These metrics evaluate the performance of the model in determining the need for additional fertilization.
 2 Related Work
-Gianquinto et al.
-conducted a study using a multispectral radiometer positioned directly above the tomato plant canopy to investigate the correlation between light reflectance, chlorophyll concentration, and total Kjeldahl nitrogen content in an intervention study of fertilizer application.
+To address the challenges identified in early stress detection, particularly in young tomato leaves, this section reviews optical measurements that have emerged as promising approaches for detecting fertilizer stress.
+In our literature review, we specifically focused on studies that examined plant responses under varying fertilizer conditions, whether through controlled interventions or through observations across fields with distinct fertilization regimes.
+We distinguished these studies from studies conducted under uniform fertilizer conditions that only identified correlations between chemical analyses and plant phenotypes.
+Previous studies have explored various optical methods for assessing plant nutrient status.
+Of particular interest were studies that combined controlled fertilizer treatments with noninvasive optical assessment methods, as they provide insights into both measurement methodology and its practical limitations.
+For example, Gianquinto et al.
+conducted an intervention study in which a multispectral radiometer positioned directly above the tomato plant canopy was used to investigate the correlations among light reflectance, chlorophyll concentration, and total Kjeldahl nitrogen content.
 By employing this noncontact arrangement, measurements were conducted without physically interacting with the leaves.
-They discovered a linear relationship between optical reflectance and chlorophyll concentration; however, no such linear relationship was identified between the optical reflectance and total Kjeldahl nitrogen content [gianquinto2011a].
+They discovered a linear relationship between the optical reflectance and chlorophyll concentration; however, no such linear relationship between the optical reflectance and total Kjeldahl nitrogen content was identified [gianquinto2011a].
 Similarly, Padilla et al.
 directly observed the most recently fully expanded leaves via an optical leaf-clip chlorophyll meter (Konica Minolta, SPAD-502).
-A linear relationship was observed between optical observations and the nitrogen nutrition index under conditions of a flattened leaf surface structure [padilla2015threshold].
-In both investigations, the absence of any barriers along the path of light between the observing device and the plant leaves is assumed, highlighting a gap between the specifications of diagnostic equipment for farmers and the current state of research and development.
-However, in contrast to other crops, tomatoes have numerous trichomes, which are small projections or hair-like structures, on the surface of their leaves (Figure 1).
-Trichomes are densely concentrated on young, actively growing leaves and sparsely distributed on older leaves [johnson1975plant].
-The presence of surface microstructures, such as trichomes, on plant bodies is known to considerably affect their appearance [kinoshita2005structural, middleton2024selfassembled, thomas2010function].
-Given their high density on young tomato leaves, it is plausible that trichomes may influence the effectiveness of optical sensing techniques used to assess the nutrient status of tomato plants.
-Consequently, the presence of trichomes should be considered a potential practical limitation when optical sensing techniques are used on young tomato leaves to evaluate fertilizer stress.
-However, current methodologies do not adequately address the impact of trichomes on the optical sensing techniques used for the rapid assessment of tomato plant nutrient status.
+A linear relationship was observed between optical observations and the nitrogen nutrition index under conditions where the leaf surface was mechanically flattened by the measurement clips [padilla2015threshold].
+In both investigations, the absence of any barriers along the path of light between the observing device and the plant leaves was assumed, highlighting a gap between the specifications of diagnostic equipment for farmers and the current state of research and development.
+While these optical methods have shown promise in various crops, tomato leaves present unique characteristics that create both challenges and opportunities for nutrient status assessment.
+Unlike many other crops, tomatoes have numerous trichomes, which are small projections or hair-like structures, on the surface of their leaves (Figure 1).
+These trichomes are densely concentrated on young, actively growing leaves and sparsely distributed on older leaves [johnson1975plant].
+The presence of such surface microstructures is known to considerably affect their appearance [kinoshita2005structural, middleton2024selfassembled, thomas2010function].
+Given their high density on young tomato leaves, it is plausible that trichomes significantly influence the effectiveness of optical sensing techniques used to assess plant nutrient status.
+Current methodologies do not adequately address this impact when evaluating fertilizer stress, particularly in young leaves.
+Therefore, the presence of trichomes should be recognized as an overlooked nonfunctional requirement in measurement device design, and it must be incorporated into the design process.
 Figure 1: Many trichomes-small projections or hair-like structures-are present on the surface of tomato leaves.
-Counting trichomes, which are abundant on young tomato leaves, is not easily achievable via visual means.
-Manually counting trichomes via a microscope is one of the few methods available for quantifying their density.
-Trichomes are abundant on young tomato leaves and are not easily distinguishable by visual means alone.
-The lack of a specialized device for assessing trichome density may have hindered progress in this area of research.
-The chlorophyll content, which is correlated with leaf color, is a widely used indicator of fertilizer stress in plants.
-However, it is not the only marker.
-Trichomes have also been identified as potential indicators of fertilizer stress, as they contribute to plant defense against pests by responding to fertilizer stress and altering plant density [barbour1991interaction, hoffland2000nitrogen].
-Various types of trichomes can be observed on the surface of tomato leaves, most of which are short, hair-like trichomes and trichomes with large, spherical cells at their tips (known as type VI trichomes, Figure 2) [wilkens1996resource].
+However, trichomes themselves may offer a novel approach for stress detection.
+In addition to posing a measurement challenge, trichomes have been identified as potential indicators of fertilizer stress, as they contribute to plant defense against pests by responding to fertilizer stress and altering plant density [barbour1991interaction, hoffland2000nitrogen].
+Various types of trichomes can be observed on the surface of tomato leaves, most of which are short, hair-like trichomes and trichomes with large, spherical cells at their tips (known as type VI trichomes; Figure 2) [wilkens1996resource].
 The glandular head of type VI trichomes is connected to the stalk cell at the base by fragile cell walls [bergau2015the].
+This structural arrangement allows the glandular part to easily detach from the leaf surface upon physical stimulation.
+The mucilage within the glandular part is known to inhibit pest activity.
+This detachment mechanism, along with the mucilage that inhibits pest activity, demonstrates how trichomes actively respond to environmental conditions.
+While chlorophyll content, which is correlated with leaf color, remains a widely used indicator of fertilizer stress in plants, the structural and functional characteristics of trichomes suggest their potential as alternative stress indicators.
 Figure 2: Type VI trichomes on the tomato plant surface.
-This structural arrangement allows the glandular part to easily detach from the leaf surface upon physical stimulation. The mucilage within the glandular part is known to inhibit pest activity.
+Multiple studies have demonstrated relationships between trichomes and nutrient status.
 Researchers have observed multiple instances where trichomes respond to fertilizer stress [barbour1991interaction, hoffland2000nitrogen, li2014the].
@@
 Similarly, Hoffland et al.
-reported a positive correlation between the leaf carbon:nitrogen ratio and trichome density [hoffland2000nitrogen].
+reported a positive correlation between the leaf C:N ratio and trichome density [hoffland2000nitrogen].
 In contrast, Leite et al.
-presented data indicating that trichome density increased over time when less fertilizer was used compared with that of conventional farming methods [leite].
-Despite these findings, most previous studies have focused on investigating trichome density within a biological context, with limited exploration of techniques to observe trichome density in agricultural production settings without advanced laboratory equipment.
-This limitation highlights the need for further research to develop practical methods for assessing trichome density in real-world agricultural environments.
-To address this gap, the present study focused on tomato leaf trichomes and explored their potential as indicators for assessing the nutrient status of tomato plants.
-As depicted in Figure 3, our main hypothesis is that leaf trichome density is influenced by fertilizer level.
-These findings suggest its potential as an additional indicator of plant nutrient status in conjunction with traditional fertilizer stress markers such as the leaf nitrate content.
-To this end, a tool for assessing trichome density must be developed.
+reported that trichome density increased over time under reduced fertilizer conditions compared to conventional farming methods [leite].
+However, these studies primarily investigated trichome density within a biological context, with limited exploration of practical measurement techniques for agricultural settings [barbour1991interaction, hoffland2000nitrogen, li2014the].
+Based on these methodological limitations, we identified two potential approaches for overcoming the challenge of early stress detection in young tomato leaves.
+These approaches involve modifying existing optical methods through additional data processing and fundamentally redesigning the measurement strategy to utilize trichomes as indicators.
+To properly account for trichomes in optical measurements, methods for quantifying their density are essential, regardless of which approach is chosen.
+The latter approach is particularly promising because it could satisfy both the prerequisites for optical measurements and the nonfunctional requirement of early stress detection through young leaf measurements with simpler implementation, as it directly utilizes the surface structures that are inherently abundant in young leaves.
+As illustrated in Figure 3, our main hypothesis derived from the literature review is that leaf trichome density is influenced by the fertilizer level and could serve as an additional indicator of plant nutrient status, complementing traditional outcome measures such as leaf nitrate content and fruit yield.
+To address the gap noted above, the present study focused on tomato leaf trichomes and explored their potential as indicators for assessing the nutrient status of tomato plants in real-world agricultural environments.
+Figure 3: Hypothesized relationships between the fertilizer level, cell differentiation, leaf nitrate content, the number of leaf trichomes, and crop yield for tomato plants, serving as the conceptual framework for the study.
+To implement this approach, however, practical methods for quantifying trichome density must be developed.
+Currently, manual counting via microscopy remains one of the few available methods, and the lack of specialized measurement tools may have hindered progress in this area.
 The identification of distinct items within a particular setting containing several objects inside the visual field is commonly considered a computer vision problem [zhao2019object, vincent1991watersheds].
-Given this context, we hypothesized that the use of cameras could convert the challenge of measuring trichome density to an analogy of an object detection problem.
-Figure 3: Hypothesized relationships between fertilizer level, cell differentiation, leaf nitrate content, number of leaf trichomes, and crop yield for tomato plants, serving as the conceptual framework for the study.
-Recently, the increasing prevalence of smartphones, with their advanced computational capabilities, high-resolution cameras, and internet access, has spurred the development of proof-of-concept agricultural technologies.
-These technologies have a wide range of applications, including farm management, crop monitoring, and pest and disease diagnosis, as highlighted in a comprehensive review by Mendes et al.
-[mendes2020smartphone].
+Given this context, we hypothesized that the use of cameras could transform the challenge of measuring trichome density into an analogy of an object detection problem.
+Given these challenges in trichome measurement and the need for practical field-based solutions, we explored the potential of widely available technology.
+As highlighted in a comprehensive review by Mendes et al.
+[mendes2020smartphone], with their portability, advanced computational capabilities, high-resolution cameras, and internet connectivity, smartphones have enabled various agricultural monitoring applications, including farm management, crop monitoring, and pest and disease diagnosis.
+Several successful implementations demonstrate the potential of smartphone-based imaging.
 For example, Neumann et al.
-demonstrated a method for classifying diseases in sugar beet leaves using images captured directly with a smartphone camera [neumann20142014], and Johannes et al.
-similarly classified wheat diseases via direct leaf images [elhassouny20192019].
-Furthermore, Li et al.
-successfully achieved early detection of late blight in tomato leaves using measurement sheet[li2019noninvasive].
-In their study, images of a measurement sheet used to analyze leaf volatiles were captured via a smartphone.
-Inspired by these advancements, our study aimed to investigate the feasibility of using smartphones as a cost-effective tool for measuring trichome density.
-These parameters can serve as predictive indicators of tomato plant nutrient status.
+developed methods for classifying sugar beet leaf diseases [neumann20142014], Elhassouny et al.
+obtained similar results with tomato diseases [elhassouny20192019], and Li et al.
+demonstrated early detection of tomato late blight via smartphone images of measurement sheets [li2019noninvasive].
+Building on these successes in detecting fine-scale plant features, our study investigated the feasibility of using smartphones as a cost-effective tool for measuring trichome density as an indicator of tomato plant nutrient status.
 3 Materials and Methods
@@
 3.1.1 Design of a simple kit for collecting trichomes from tomato leaves
-In this study, we designed a straightforward kit to measure the density of type VI glandular trichomes on tomato leaves.
-This kit comprises custom-made measurement paper crafted from inkjet-printable cardstock and cellophane tape (NICHIBAN, CT-18), enabling cost-effective measurements (Figure 4).
-This paper contains augmented reality (AR) markers, enabling the distance of objects within captured images to be precisely measured.
-Additionally, there is a 12 mm $\times$ 12 mm opening in the center of the paper.
-A piece of cellophane tape is affixed to the back of the paper, covering this opening.
-This tape is used to extract trichomes from the surface of tomato leaves.
+To establish practical nutrient assessment methods for young tomato leaves, we developed a straightforward measurement kit that addresses the need for young leaf measurements, the limitations of existing optical methods in the presence of trichomes, and the need for field-applicable solutions.
+The kit was designed to measure the density of type VI glandular trichomes on tomato leaves as indicators of nutrient status.
+To ensure accessibility and practical application in agricultural settings, we designed the kit using readily available materials, specifically custom-made measurement paper crafted from inkjet-printable cardstock and cellophane tape (NICHIBAN CT-18).
+The selection of cellophane tape was driven by the three-dimensional characteristics of tomato leaves.
+The leaf veins and natural leaf curling create macroscopic undulations with longer periods than those caused by trichome surface structures.
+Given these topographical features, we needed a measurement method compatible with cameras that have a limited depth of field.
+Cellophane tape, which can closely conform to the leaf surface, was selected as both a thin-film substrate and transfer medium, enabling the conversion of three-dimensional trichome distributions into measurable two-dimensional patterns.
+This design choice enables cost-effective measurements while maintaining reliability (Figure 4).
+To achieve consistent measurements across varying field conditions, the measurement paper incorporates augmented reality (AR) markers, which enable precise distance calibration within captured images, regardless of the camera position or environmental conditions.
+Additionally, there is a 12 mm $\times$ 12 mm opening in the center of the paper, which is covered by the cellophane tape for trichome extraction.
 Figure 4: A smartphone was used to capture an image of the diagnostic kit.
-Optimizing a technique that involves the use of cellophane tape for trichome extraction allows the creation of a highly cost-efficient kit.
-Although recent advances in sophisticated object detection algorithms have utilized complex neural networks, these neural networks typically demand high-speed parallel processing, considerably increasing computational costs.
-Our approach, which uses cellophane tape for trichome extraction, circumvents the need for such intensive computational resources and enables the use of less computationally demanding algorithms for trichome detection.
-Thus, this approach reduces the overall cost of the assessment process, making it more economically viable.
-The purpose of this diagnostic kit is to meet the technological progress needs in agricultural production areas that lack access to laboratory-level experimental facilities.
-Therefore, to minimize the number of components required for measurement, the kit is specifically designed to allow for dual manual operation.
-The measuring paper is held in one hand, whereas the smartphone is operated in the other hand (Figure 4).
-The statistical model accounts for the variability in the results caused by varying optical conditions among shots.
-The kit is designed to prioritize user friendliness as a minimum viable product in lean startup methodology for manufacturing and software development, enabling the measurement of trichome density without requiring technical expertise [mu2013lean, Ries2011-xn].
-This technology is well suited for practical application in the field because of its simplicity and cost-effectiveness (Figure 5).
+The development of this measurement approach was guided by the practical constraints of agricultural applications.
+While recent advances in object detection have predominantly relied on complex neural networks requiring high-speed parallel processing, our optimized approach using cellophane tape for trichome extraction circumvents these computational demands.
+This technique, combined with simplified image processing algorithms, creates a practical solution that maintains accuracy while eliminating the need for intensive computational resources, which makes the assessment process more economically viable for agricultural applications.
+This diagnostic kit was developed to meet technological needs in agricultural production areas with limited access to laboratory facilities, fulfilling the requirement for reliable nutrient assessment tools in field conditions.
+To minimize the number of components required for measurement, the design enables dual manual operation, with the measuring paper in one hand and the smartphone in the other (Figure 4), while maintaining measurement reliability through statistical modeling that accounts for varying optical conditions.
+In accordance with lean startup methodology principles [mu2013lean, Ries2011-xn], the kit was developed as a minimum viable product that prioritizes user-friendly operation without requiring technical expertise, enabling rapid assessment of plant nutrient status in the field (Figure 5).
 Figure 5: Proposed concept of the tomato cultivation support framework.
 3.1.2 Proposed concept of the tomato cultivation support framework
-This kit is designed to empower farmers to assess the additional fertilization needs of their tomato plants based on trichome density and receive fertilization recommendations.
-The proposed framework involves a four-step process.
-First, trichomes are sampled from tomato leaves via adhesive cellophane tape on measurement paper.
-Second, a smartphone is used to capture an image of the measurement paper with the transferred trichomes.
-Third, the captured image is analyzed to determine trichome density, which serves as an indicator of the degree of fertilizer stress.
-Finally, on the basis of the analyzed trichome density, the system provides a fertilization recommendation to the farmer, guiding their decision on whether additional fertilization is necessary.
+To translate trichome density measurements into practical agricultural decisions, we developed a systematic support framework that empowers farmers to assess the nutritional needs of their tomato plants.
+This framework streamlines the assessment process through four sequential steps.
+The process begins with sampling trichomes from tomato leaves via adhesive cellophane tape on measurement paper.
+Next, a smartphone is used to capture an image of the transferred trichomes.
+The captured image is then analyzed to determine trichome density as an indicator of fertilizer stress.
+Finally, specific fertilization recommendations are generated based on the analysis results.
+This structured approach enables farmers to make informed decisions about whether additional fertilization is necessary, replacing subjective assessments with objective evaluations.
 3.1.3 Overview of the trichome detection image processing pipeline
-This study implemented an image processing pipeline for calculating trichome density.
-Instead of using deep learning models for direct image processing, tabular data analysis was chosen.
-Although deep learning often requires substantial computational resources, this architecture leverages the inherent simplicity and reduced dimensionality of tabular data, enabling the use of computationally lighter models.
-This method effectively minimizes computational costs while bolstering quantitative model interpretation, thereby improving the reliability and interpretability of the results for agricultural applications.
-To avoid data leakage, an image processing pipeline was developed using images obtained from preliminary experiments.
-This approach ensured that the model development and evaluation processes remained independent and unbiased.
-The image processing pipeline comprises three main stages.
-The preprocessing stage involves extracting metadata in the exchangeable image file format (EXIF) embedded in the image files, identifying AR markers within the images, applying homography transformations for image alignment, and adjusting for variations in lighting conditions.
-This preprocessing stage also includes binarizing images to segment trichomes and applying noise removal filtering algorithms to enhance trichome visibility.
-In the feature extraction stage, we focused on identifying trichomes in the images, specifically targeting type VI glandular trichomes, owing to their unique morphology.
-Furthermore, AR markers were utilized to convert pixel counts to relative distances, enabling accurate distance estimation.
-The final stage, density calculation, involves calculating the trichome density from images captured under various environmental conditions and camera settings, each associated with distinct EXIF data.
-Algorithm 1 presents an outlined function of the trichome density computation process using our custom-developed measurement kit.
-The image processing procedures described above were executed via the software specifications outlined in Table 3.
+The success of this support framework critically depends on accurate and reliable trichome density measurements.
+To realize such measurements while maintaining practical utility in agricultural settings, we developed an image processing pipeline that prioritizes computational efficiency and result interpretability.
+Rather than employing deep learning models for direct image processing, we chose a hybrid approach that combines traditional computer vision techniques with tabular data analysis.
+This architecture leverages the inherent simplicity and reduced dimensionality of tabular data, effectively minimizing computational costs while bolstering quantitative model interpretation.
+This approach enables the use of computationally lighter models while improving both the reliability and interpretability of the results for agricultural applications.
+Our development process prioritized methodological rigor and practical reliability.
+To avoid data leakage, we developed the image processing pipeline using images from preliminary experiments, ensuring that the model development and evaluation processes remained independent and unbiased.
+The pipeline consists of three main stages, beginning with preprocessing, where we extract metadata in the exchangeable image file format (EXIF) embedded in the image files.
+At this stage, we identify AR markers within the images for initial alignment, apply homography transformations, and adjust for variations in lighting conditions.
+Additional preprocessing steps include image binarization for trichome segmentation and noise removal filtering to enhance trichome visibility.
+In the feature extraction stage, we focus on identifying trichomes in the images, specifically targeting type VI glandular trichomes owing to their unique morphology.
+Using the aligned images and calibrated AR markers from the preprocessing stage, we convert the pixel counts to relative distances, enabling accurate distance estimation.
+The final density calculation stage leverages the extracted features to determine trichome density from images captured under various environmental conditions and camera settings, each associated with distinct EXIF data.
+The computational process is detailed in Algorithm 1, with the software specifications provided in Table 3.
 Algorithm 1 Trichome density analysis process
@@
 3.1.4 Image preprocessing: AR marker detection, illumination correction, and noise reduction
-The first step of the developed image processing pipeline is to read EXIF data from the image files.
-This metadata contains valuable information about the environmental conditions and camera settings at the time of image capture.
-By analyzing aspects such as exposure time and ISO speed ratings, we can gain a comprehensive understanding of the variations in imaging conditions and their potential impacts on trichome density measurements.
-To measure the absolute scale and orientation of the captured images, we incorporated AR markers into our image processing pipeline.
-These markers enabled us to determine the trichome density of the captured images.
-For AR marker detection, we utilized the ArUco module from the OpenCV library (version 4.6.0.66) [intelcorporationopencvcontribpython].
-The AR marker detection algorithm employs a simple, nonrecursive approach, as the development of AR technology was not the primary focus of this study.
-The process involves image preprocessing, where the captured image is converted to grayscale and then binarized via Otsu’s method to enhance marker visibility under various lighting conditions.
-We used a predefined "ArUco dictionary", DICT_4X4_250.
-This dictionary offers a balance between the number of unique markers and detection robustness.
-Finally, the OpenCV function cv2.aruco.detectMarkers() is used to identify the ArUco markers in the preprocessed image.
-The complete algorithm is presented in Algorithm 2.
+Image processing begins with reading EXIF metadata from image files.
+This processing provides crucial information about environmental conditions and camera settings during image capture.
+Analysis of specific parameters, such as exposure time and ISO speed ratings, enables us to account for variations in imaging conditions that affect trichome density measurements, ensuring consistency across different environments.
+Our spatial calibration incorporates AR markers into the image processing pipeline.
+These markers serve as reference points for determining the scale and orientation of the captured images, enabling trichome density determination.
+For marker detection, we implemented a nonrecursive approach via the ArUco module from the OpenCV library (version 4.6.0.66) [intelcorporationopencvcontribpython], as the development of AR technology was not the primary focus of this study.
+For robust detection under field conditions, we implemented a systematic image preprocessing approach.
+The algorithm converts the captured image to grayscale, followed by binarization via Otsu’s method to increase marker visibility across lighting conditions.
+We selected the "ArUco dictionary" DICT_4X4_250 due to its balance between marker uniqueness and detection reliability.
+The markers are identified via the OpenCV function cv2.aruco.detectMarkers().
+The complete marker detection algorithm is presented in Algorithm 2.
 Algorithm 2 AR marker detection
@@
 Following AR marker detection, we apply homography transformation to correct perspective distortion in images.
-This distortion is caused by positional and orientational variations in the measurement paper relative to each camera captured.
-The homography transformation adjusts the images as though the camera and the measurement paper were parallel.
-This process involves using the coordinates of the four corners of a region on the measurement paper, obtained from detecting AR markers as reference points to estimate a perspective transformation matrix.
-By applying this transformation, we ensure that the images are standardized for accurate trichome density analysis, regardless of slight variations in camera positioning during image capture.
-After applying the homography transformation, we implemented an illumination correction procedure to address nonuniform lighting conditions across the acquired images.
-This step is essential for maintaining consistency in trichome detection, regardless of ambient lighting variations during image acquisition.
-Our approach uses a flat field correction technique [seibert1998medical], which involves several preprocessing steps.
-Initially, all the images are resized to a standardized dimension of 1000 pixel $\times$ 1000 pixel to ensure uniform processing.
-We subsequently apply an averaging filter with a kernel size of 50 pixel $\times$ 50 pixel to generate a smoothed version of the image, representing the overall illumination pattern.
-The correction is mathematically described by the following equation:
-$C=\frac{R\times m}{F}$ (1)
+This distortion arises from positional and orientational variations in the measurement paper relative to each camera captured.
+The transformation adjusts images as though the camera and measurement paper were parallel, using the coordinates of four corners obtained from AR markers as reference points to estimate a perspective transformation matrix.
+This standardization ensures accurate trichome density analysis, regardless of variations in camera positioning during image capture.
+Image quality optimization begins with illumination correction, which is an essential step to address nonuniform lighting conditions across the acquired images and to maintain consistent trichome detection despite ambient lighting variations.
+For this purpose, we implemented a flat field correction technique [seibert1998medical].
+The process starts by resizing all images to a standardized dimension of 1000 pixels $\times$ 1000 pixels, followed by applying an averaging filter with a kernel size of 50 pixels $\times$ 50 pixels to generate a smoothed version representing the overall illumination pattern.
+The correction is mathematically described in Equation 1.
+$C=\frac{R\cdot m}{F}$ (1)
 where $C$ denotes the corrected image, $R$ represents the original (precorrection) image, $m$ is the average luminance of the original image, and $F$ is the smoothed image representing the illumination pattern.
@@
 38:$m\leftarrow\textsc{CalculateAverageLuminance}(R)$
-39:$C\leftarrow\frac{R\times m}{F}$
+39:$C\leftarrow\frac{R\cdot m}{F}$
 […]
 3.1.5 Feature extraction: Trichome identification and density analysis
-Following noise removal, a segmentation procedure based on the watershed algorithm was implemented to identify individual trichomes.
-This method detects boundary surfaces that separate contours in the image.
-The watershed algorithm was chosen because of its suitability for rule-based segmentation on the basis of the distinctive appearance of trichomes, offering a computationally efficient solution.
-Although deep learning-based segmentation can often achieve high accuracy, it typically requires substantial computational resources, often necessitating the use of devices with advanced parallel processing capabilities, such as graphics processing units, to be integrated into the computer system.
-In contrast, the watershed algorithm, which is a traditional image-processing technique, generally operates with lower computational costs and can often perform effectively on standard computer hardware, such as central processing units.
-This characteristic is particularly important for our application, as it ensures that the tomato nutrient diagnostic kit remains cost-effective for farmers, aligning with our goal of developing accessible and practical tools for agriculture.
+For individual trichome identification, we implemented a segmentation procedure based on the watershed algorithm following the noise removal step.
+This method detects boundary surfaces that separate contours in the image, offering an efficient approach for rule-based segmentation suited to the distinctive appearance of trichomes.
+While deep learning-based segmentation can achieve high accuracy, it typically requires substantial computational resources, including graphics processing units with advanced parallel processing capabilities, to be integrated into the computer system.
+In contrast, the watershed algorithm, which constitutes a traditional image-processing technique, operates with lower computational costs on standard computer hardware such as central processing units, making it particularly suitable for achieving our goal of developing accessible and practical tools for agriculture.
 […]
-In this paper, we utilize the term "trichomes" to describe the contours of detected objects based on the following criteria:
+In this paper, we utilize the term "trichomes" to describe the contours of detected objects on the basis of their specific properties. These criteria for identification include the following:
   • A detachable layer must exist, allowing peeling with cellophane tape;
@@
   • An object’s shape should more closely resemble a sphere than a prolate ellipsoid, with the circularity exceeding a defined lower limit;
-  • The perimeter of the object falls within a specified range, defined by a lower and upper limit;
-  • The area of the object lies within a specified range, defined by a lower and upper limit.
-The contours detected through segmentation yield centroid points, denoted as in Equation 2.
-These points exist within a two-dimensional space and are represented via the pixel coordinates of the image.
-Figure 6 shows the image processing steps performed up to this point, including preprocessing and object detection.
-The watershed segmentation process yields a set of detected contours, from which we extract centroid points.
-These points are represented in a two-dimensional space corresponding to the pixel coordinates of the image.
-The set of centroids is denoted as $P$, as defined in Equation 2.
-$P={p_{i}\in\mathbb{R}^{2}\mid i=1,2,\ldots,N}$ (2)
-where $p_{i}$ represents the i-th centroid point in the two-dimensional real space $\mathbb{R}^{2}$, and $N$ is the total number of detected contours.
-Each point $p_{i}$ can be expressed as $(x_{i},y_{i})$, where $x_{i}$ and $y_{i}$ are the horizontal and vertical pixel coordinates, respectively.
+  • The perimeter of the object falls within a specified range, defined by a lower limit and an upper limit;
+  • The area of the object lies within a specified range, defined by a lower limit and an upper limit.
+The contours detected through segmentation yield centroid points in two-dimensional space, represented via the pixel coordinates of the image.
+Figure 6 shows the image processing steps performed thus far, including preprocessing and object detection, leading to these contour detections.
+The set of centroids, denoted as $P$, is expressed in Equation 2.
+$P=\{\bm{p}_{i}\in\mathbb{Z}^{2}\mid i=1,2,\ldots,N\}$ (2)
+where $\bm{p}_{i}$ represents the $i$-th centroid point in the two-dimensional integer space $\mathbb{Z}^{2}$ and $N$ is the total number of detected contours.
+Each point $\bm{p}_{i}$ is represented by its pixel coordinates $(x_{i},y_{i})$, where $x_{i}$ and $y_{i}$ correspond to the horizontal and vertical positions in the image, respectively.
 Algorithm 4 Trichome identification
@@
 18:function CalculateAllMetrics($contours$)
-19:     Calculate circularity, perimeter, and area for all contours
+19:     Calculate the circularity, perimeter, and area for all contours
 20:     return $C_{all},P_{all},A_{all}$
@@
 The detailed process of trichome identification, including the criteria and thresholds used for filtering, is presented in Algorithm 4.
-This algorithm encapsulates the key steps and decision-making process involved in distinguishing type VI glandular trichomes from other objects in the image.
-Figure 6 shows the image processing steps performed up to this point, including preprocessing and object detection.
+This algorithm encapsulates the key steps and decision-making process for distinguishing type VI glandular trichomes from other objects.
 Figure 6: Example of image data preprocessing and object detection.
-The density was computed for the centroid points denoted as $P$ in Equation 2.
-We employed the nearest neighbor distance (NND) method, as described in Equation 3 [clark1954distance], to quantify the spatial distribution of these points.
-The NND method involves measuring the distance between each point and its nearest neighbor within the set.
-$NND=\frac{1}{n}\sum_{i=1}^{n}d(p,q)$ (3)
-In this equation, $d(p,q)$ represents the Euclidean distance between point $p$ and its closest neighboring point $q$.
-The NND is then calculated by averaging these distances over the total number of contours $n$ present in the image.
-Efficiently computing the nearest neighbors of each point is crucial, especially when many points are dealt with.
-A brute-force approach, where distances between all point pairs are calculated, would have a computational complexity of $O(nd)$, where $n$ is the number of data points and $d$ is the dimensionality.
-This "big O" notation, $O(nd)$, describes how the execution time of the algorithm increases as the input size ($n,d$) increases[skiena1998algorithm].
-In this case, $O(nd)$ means that the execution time increases linearly with the number of points and the dimensionality.
-Therefore, the brute-force approach can become computationally expensive as the number of points increases.
-To address this issue, we utilized a k-dimensional tree algorithm, which organizes the points in a space-partitioning data structure that allows for rapid nearest neighbor queries.
-This algorithm effectively reduces the computational complexity to $O(d\log n)$ [cunningham2022knearest], meaning that the execution time grows linearly with the dimensionality but only logarithmically with the number of points.
-This logarithmic growth makes the k-d tree algorithm much more efficient than the brute-force approach, especially for large datasets.
-This algorithm is particularly suitable for our two-dimensional image analysis ($d=2$) with many trichome centroids.
+To quantify the spatial distribution of the identified trichomes, we computed the density of the centroid points (denoted as $P$ in Equation 2) via the nearest neighbor distance (NND) method [clark1954distance].
+This approach measures the distance between each point and its nearest neighbor within the set, as described in Equation 3.
+In this equation, $\mathrm{d}(\bm{p}_{i},\bm{q}_{i})$ represents the Euclidean distance between point $\bm{p}_{i}$ and its closest neighboring point $\bm{q}_{i}$, with the NND calculated by averaging these distances over the total number of contours $N$ present in the image.
+$\mathrm{NND}=\frac{1}{N}\sum_{i=1}^{N}\mathrm{d}(\bm{p}_{i},\bm{q}_{i})$ (3)
+The efficient computation of nearest neighbors becomes crucial when dealing with numerous points.
+A brute-force approach, which calculates the distances between all point pairs, has a computational complexity of $O(dn^{2})$, where $n$ is the number of data points and $d$ is the dimensionality.
+As described by this "big O" notation [skiena1998algorithm], the execution time of the algorithm increases as the input size ($n,d$) increases, making the brute-force approach computationally expensive as the number of points increases.
+To address this computational challenge, we adopted a k-dimensional (k-d) tree algorithm, which organizes points in a space-partitioning data structure for rapid nearest neighbor queries.
+This approach effectively reduces the computational complexity to $O(n\cdot d\log n)$ [cunningham2022knearest], where the execution time increases linearly with dimensionality and increases more moderately with the number of points compared to the brute-force approach.
+This logarithmic growth characteristic makes the k-d tree algorithm much more efficient than the brute-force approach, especially for large datasets.
+The algorithm is particularly suitable for our two-dimensional image analysis ($d=2$) with numerous trichome centroids.
 The detailed implementation of the NND calculation is presented in Algorithm 5.
@@
 28:$NND\leftarrow\textsc{CalcNND}(P)$
-The NND serves as an inverse indicator of contour density, with higher values denoting sparse arrangements and lower values signifying denser groupings.
-Glandular type VI trichomes are susceptible to detachment from leaf surfaces because of factors such as insect infestations or physical abrasion.
-However, NND demonstrates superior robustness to such accidental detachment compared with conventional density indices, which are based on the object count per unit area.
-This robustness arises because NND represents density as a probability distribution, which enables it to maintain reliable estimates even when localized areas experience trichome detachment, provided that unaffected regions persist.
-Notably, the NND calculation requires the presence of at least two points.
-The appendix elucidates this concept through a simulation employing point cloud data generated via a Poisson process in two-dimensional space.
+The NND serves as an inverse indicator of spatial density, with higher values denoting sparse arrangements and lower values representing denser groupings.
+While conventional density measurements simply count objects per unit area, this measure proves particularly valuable for analyzing type VI glandular trichomes, which are susceptible to detachment from leaf surfaces due to factors such as insect infestations or physical abrasion.
+Compared with conventional counting methods, the NND method demonstrates superior robustness to accidental detachment.
+The reason for this robustness is that the NND method represents density as a probability distribution, enabling reliable estimates even when localized areas experience trichome detachment, provided that unaffected regions persist.
+As a prerequisite for NND calculations, at least two points must be present in the measurement area.
+The appendix further explains this concept through a simulation employing point cloud data generated via a Poisson process in two-dimensional space.
 3.2 Evaluation of the New Method for Assessing the Nutrient Status of Tomato Plants
 3.2.1 Selection of the cultivation method
-A well-controlled cultivation experiment is essential for evaluating the proposed method for assessing the nutrient status of tomato plants with testability.
-This cultivation experiment aimed to measure phenotypes such as trichome density and leaf nitrate ion concentration and fruit yield under ideal conditions for tomato growth, minimizing the variability inherent in soil-based cultivation.
-Therefore, we opted for hydroponics rather than soil-based cultivation to circumvent the complexities associated with soil properties such as variations in soil composition and structure.
-These variations can lead to complexities in cultivation experiments, including inconsistent water retention, fertilizer dynamics, and potential for pest and disease outbreaks.
-This decision allowed for more precise control over nutrient availability and minimized potential confounding factors in our evaluation of fertilizer stress.
-This approach also enabled us to investigate the relationships between nutrient levels and plant responses more clearly in an ideal growth environment that approximates the optimal growth conditions found in some commercial cultivation systems.
-Among various hydroponic techniques, we selected the deep water culture (DWC) method for its simplicity, minimal equipment requirements and ease of implementation in a randomized block design.
-In DWC, plant roots are fully submerged in nutrient solution within a production tank, which is aerated to sufficiently oxygenate the roots [hoagland1938the, velazquezgonzalez2022a].
-This straightforward approach facilitates the relocation and replication of experimental units regardless of location, contributing to increased falsifiability.
-Moreover, DWC allows easy control of nutrient solution conditions at the individual plant level, making it suitable for randomized block designs.
+A well-controlled cultivation experiment is essential for evaluating our proposed method for assessing tomato plant nutrient status while maintaining testability.
+The aim of the experiment was to measure key phenotypes, i.e., trichome density, leaf nitrate ion concentration, and fruit yield, under ideal growth conditions.
+We chose hydroponics over soil-based cultivation to avoid the complexities inherent in soil properties, such as variations in composition and structure that can lead to inconsistent water retention, unpredictable fertilizer dynamics, and increased risk of pest and disease outbreaks.
+This approach enabled precise control over nutrient availability, and it minimized confounding factors in our evaluation of fertilizer stress while providing an ideal growth environment that approximates the optimal conditions found in commercial cultivation systems.
+From the available hydroponic techniques, we selected the deep water culture (DWC) method, where plant roots are fully submerged in nutrient solution within a production tank, which is aerated to sufficiently oxygenate the roots [hoagland1938the, velazquezgonzalez2022a].
+The precise control of nutrient conditions at the individual plant level makes this method particularly suitable for implementation in a randomized block design.
+Additionally, its simplicity of setup and minimal equipment requirements facilitate both the relocation and replication of experimental units across locations, enhancing the experiment’s falsifiability.
 3.2.2 Tomato growth conditions in hydroponic culture
 […]
-Within these buckets, the nutrient mixture was aerated and mixed by introducing air supplied by a blower (YASUNAGA, AP-40P).
-The air was divided into eight segments, each channeled through silicone tubing and dispersed into the nutrient solution.
-The use of air stones ensures proper mixing and oxygenation.
-The buckets were equipped with lids containing center holes for the tomato stems to prevent algae growth in the nutrient solution.
+The nutrient solution in each bucket was continuously aerated and mixed via a blower system (YASUNAGA, AP-40P), which distributed air through eight segments of silicone tubing and dispersed it into the nutrient solution via air stones for proper oxygenation.
+To prevent algae growth in the nutrient solution, the buckets were fitted with lids featuring center holes for tomato stem placement.
 3.2.3 Fertilizer concentration settings
-The nutrient solution was prepared via hydroponic fertilizer (OAT Agrio, Fertilizer Series SA), and maintained at three different total dissolved solids (TDS) concentrations to induce various levels of fertilizer stress:
-  • High concentration: 1500 ppm
-  • Intermediate concentration: 1000 ppm
-  • Low concentration: 500 ppm
-This study employed an interventional research design to elucidate the causal relationship between fertilizer stress and tomato phenotypes.
-We set three fertilizer levels to investigate the presence or absence of intervention effects on tomato phenotypes and their dose-response relationships.
+To investigate the relationship between fertilizer stress and tomato phenotypes, we prepared nutrient solutions using hydroponic fertilizer (OAT Agrio, Fertilizer Series SA) at three concentrations of total dissolved solids (TDS) to induce various levels of fertilizer stress: 1500 ppm (high), 1000 ppm (medium), and 500 ppm (low).
+This interventional research design enabled us to examine both the presence of intervention effects on tomato phenotypes and their dose-response relationships.
 […]
-Based on these findings, we set the low-concentration treatment at 500 ppm TDS.
-The intermediate- and high-concentration treatments were set at twice this level (1000 ppm) and three times this level (1500 ppm), respectively.
-Furthermore, to confirm that this range of concentrations did not exert excessively adverse effects on tomato growth, we conducted preliminary cultivation experiments.
-The results revealed no visually identifiable growth defects or plant death at 500 ppm.
-Additionally, even at 1500 ppm, no apparent growth inhibition due to increased osmotic pressure was observed.
-These results confirmed the suitability of the selected fertilizer concentration range for the research objectives.
+We established the medium and high concentrations at two and three times this baseline.
+Under our greenhouse conditions, preliminary cultivation experiments confirmed the suitability of this range, showing no visually identifiable growth defects or plant death at 500 ppm and no apparent growth inhibition due to increased osmotic pressure even at 1500 ppm.
+These results validated our concentration range selection while supporting our primary goal of detecting fertilizer stress through plant phenotypes to provide actionable fertilization recommendations.
 3.2.4 Experimental design and management
-The fertilizer plots were arranged in a randomized block design, as shown in Figure 7.
-During the preliminary experiment, the tomato plants were initially cultivated in the vacant areas depicted in Figure 7.
-To evaluate the effects of fertilizer level on outcome variables, such as the concentration of nitrate ions in tomato leaves, the density of trichomes, and yield, this randomized block design was employed to minimize the potential impact of spatial variations in environmental factors [fisher1966design].
-Water was added to the buckets as needed when the nutrient solution water level decreased.
-Fertilizer was applied every two weeks, and nutrient solution concentrations were adjusted to maintain the desired TDS levels, which were measured via a TDS meter (EUTECH, PCSTestr35).
-The fertilizer used in the experiments closely matched the nutrient uptake ratios of the tomatoes, enabling the use of a closed hydroponic system without frequent nutrient solution replacement [ishihara2007effect].
-However, during the cultivation period, the system was not maintained as a completely closed system; nutrient solutions were replaced whenever they became cloudy because algae or root debris was present.
-During the cultivation period, two plants were lost before the compound leaves were harvested, leaving 20 plants available for inclusion in the study.
-Furthermore, an additional two plants were lost by the end of the fruit harvesting period, resulting in 18 plants being available for harvest yield assessment.
+To evaluate the effects of fertilizer levels on key outcome variables, i.e., nitrate ion concentration in tomato leaves, trichome density, and yield, we implemented a randomized block design that minimized the potential impact of spatial variations in environmental factors [fisher1966design].
+The fertilizer treatment plots were arranged as shown in Figure 7, with the preliminary experiment utilizing the vacant areas depicted in the figure.
+Nutrient solution management followed a systematic protocol throughout the cultivation period.
+Water was added to the buckets when nutrient solution levels decreased, while fertilizer was applied biweekly, with concentrations adjusted to the desired TDS levels via a TDS meter (EUTECH, PCSTestr35).
+Although the fertilizer composition closely matched the nutrient uptake ratios of tomatoes, enabling operation as a closed hydroponic system without frequent solution replacement [ishihara2007effect], practical considerations necessitated occasional solution replacement when the solution became cloudy due to algae or root debris.
+Plant survival was monitored throughout the experiment, with some attrition occurring at different stages.
+Two plants were lost before the compound leaf harvesting stage, reducing the study population to 20 plants.
+By the end of the fruit harvesting period, the loss of two additional plants further reduced the population to 18 plants for yield assessment.
 Figure 7: Experimental layout with three different fertilizer levels (ppm) in a randomized block design.
 3.2.5 Compound leaf sampling and trichome transfer
-The experiment involved collecting compound leaves from tomato plants.
-A total of 40 compound leaf samples were collected during two separate harvesting sessions in March and April 2023.
-The sampling process followed a systematic approach to ensure consistency and minimize potential biases.
-The samples were processed sequentially in ascending order of the number of tomato plant samples.
-These sample numbers were randomly assigned to different fertilizer treatment levels as part of the randomized block design, ensuring that the sampling order was not biased by treatment level.
-The compound leaves were sampled 15 cm below the apical meristem of each plant.
-If no suitable compound leaf was available at this position, the plant was temporarily skipped, allowing for further growth.
-The sampling process continued iteratively until all the plants had been sampled.
-For trichome density measurements, we selected all leaflets larger than the opening of the measurement paper from each compound leaf.
-The number of measurement papers used corresponded to the number of selected leaflets per compound leaf: This approach allows for an accurate assessment of trichome density distribution across each leaflet.
-Prior to image capture, glandular trichomes were transferred from the leaflet surface to the measurement paper.
-Each leaflet was carefully affixed to the adhesive side of the cellophane tape (Figure 8) and gently pressed with a finger to ensure complete contact between the leaf surface and the tape.
-After the leaflet was removed, the cellophane tape with the transferred trichomes was photographed alongside the measurement paper via a smartphone camera.
+Our experimental protocol included systematic collection of compound leaves from tomato plants, with a total of 40 samples collected across two harvesting sessions in March and April 2023.
+To ensure consistency and minimize potential biases, we implemented a structured sampling process where samples were processed sequentially according to ascending plant identification numbers.
+These IDs had been randomly assigned to different fertilizer treatment levels during the randomized block design setup, preventing any treatment-related bias in the sampling order.
+The sampling location was standardized at 15 cm below the apical meristem of each plant.
+When a suitable compound leaf was unavailable at this position, we temporarily skipped the plant to allow for additional growth, continuing the sampling process iteratively until all the plants were sampled.
+For each compound leaf, we selected all leaflets larger than the measurement paper opening for trichome density measurements.
+The number of measurement papers matched the number of selected leaflets per compound leaf, allowing for accurate assessment of trichome density distribution across each leaflet.
+The glandular trichome transfer process followed a simple yet precise protocol that balanced measurement reliability with practical field applicability.
+Each leaflet was carefully positioned on the adhesive side of the cellophane tape (Figure 8) and gently pressed with a finger to achieve complete contact between the leaf surface and the tape.
+Following leaflet removal, we photographed the cellophane tape containing the transferred trichomes alongside the measurement paper via a smartphone camera.
 Figure 8: Trichome transfer process from the leaflet surface to the measurement paper using cellophane tape.
 3.2.6 Image acquisition and dataset preparation
-The measurement paper was photographed via the native camera app on a smartphone (OPPO, Reno A).
-This device has a 16-megapixel camera featuring an F1.7 aperture and a CMOS sensor (Sony, IMX398).
-The camera supports autofocus functionality, allowing for sharp and clear images even under varying conditions.
-For our experiments, the camera mode was set to automatically adjust the exposure, ISO sensitivity, and focus, leveraging the device’s built-in optimization algorithms.
-To ensure a comprehensive dataset that accounts for various usage scenarios, we manually varied the distance and angle between the camera and the measurement paper.
-The distance from the camera sensor to the measurement paper ranged from 79 to 218 mm, simulating different user holding positions.
-The horizontal and vertical offset angles varied from -12.52 to 10.62 degrees and -9.04 to 8.13 degrees, respectively.
-This range of angles mimics slight misalignments that might occur during practical use of the kit.
-Furthermore, the measurement paper was positioned a sufficient distance from the background to ensure that the background was blurred and not clearly identifiable in the captured images.
-The room used for photography had a north-facing window, which eliminated direct sunlight and provided diffuse natural light.
-The photographic sessions were conducted at different times of the day, with solar elevation angles ranging from -7.95 to 33.42 degrees.
-The smartphone’s built-in white LED was activated for auxiliary lighting, and the measurement paper was additionally illuminated by oblique light from the ceiling’s fluorescent light (as shown in Figure 4).
-The image composition was carefully photographed to ensure that the measurement paper was centered in the frame.
-Immediately after the images were captured, a preliminary analysis was conducted to verify the completeness of the image processing pipeline.
-All the measurement papers were checked to ensure that they could be included in the dataset.
-In cases where the analysis program encountered exceptions, rephotography was performed to prevent bias due to missing information from the measurement paper during analysis.
-Consequently, this study focused exclusively on analyzing trichome density from images where AR markers could be detected.
-The accuracy of AR marker detection was assumed to impact the results inconsiderably.
-This assumption was based on the practical consideration that the detection accuracy of AR markers and other 2D codes depend primarily on the library used and has already reached a practical level for applications.
-Moreover, in the event of an error, the kit user can rectify the problem by performing rephotography.
-After the data collection and preprocessing steps, information from all the measurement papers was included in the final dataset, which comprised 6,855 images suitable for analysis.
-This comprehensive dataset encompasses a wide range of distances, angles, and lighting conditions and ensures that the developed model is trained and evaluated on a diverse set of scenarios, closely mimicking real-world usage of the kit.
+Image acquisition was performed via a smartphone (OPPO, Reno A) equipped with a 16-megapixel camera featuring an F1.7 aperture and a CMOS sensor (Sony, IMX398).
+We utilized the device’s native camera application with its autofocus capability and automatic adjustment of exposure, ISO sensitivity, and focus parameters, leveraging built-in optimization algorithms to ensure clear images across varying conditions.
+To ensure a comprehensive dataset that accounts for various usage scenarios, we manually varied the image capture parameters to validate system robustness.
+The distance between the camera sensor and measurement paper ranged from 79 to 218 mm, with horizontal offset angles ranging from -12.52 to 10.62 degrees and vertical offset angles ranging from -9.04 to 8.13 degrees, simulating natural variations in user handling.
+The measurement paper was positioned with sufficient background separation to ensure that the background was not clearly identifiable in the captured images.
+Photography was conducted in a room with a north-facing window, which eliminated direct sunlight and provided diffuse natural light, supplemented by the smartphone’s built-in white LED and oblique ceiling fluorescent lighting (Figure 4).
+Photographic sessions were scheduled across different times of day, with solar elevation angles ranging from -7.95 to 33.42 degrees.
+Quality control measures were implemented throughout the image acquisition process.
+Each image was composed with the measurement paper centered in the frame, and preliminary analysis was conducted immediately after capture to verify pipeline compatibility.
+All the measurement papers were carefully checked for dataset inclusion, with immediate rephotography performed when the analysis program encountered exceptions, mostly due to failures in two-dimensional code detection.
+The detection accuracy of these codes primarily depends on the code type and detection libraries used.
+As two-dimensional code technology has already achieved practical reliability levels, users in agricultural field conditions can simply retake photographs when the measurement software indicates detection exceptions.
+Given this independence between code detection and our primary objectives of trichome density measurement and fertilization recommendations, we excluded code detection accuracy from our verification targets.
+After the data collection and preprocessing steps, information from all the measurement papers was compiled into the final dataset, which comprised 6,855 images that successfully passed through the image processing pipeline.
+This comprehensive dataset encompasses a wide range of distances, angles, and lighting conditions, ensuring that the developed model is trained on and evaluated based on diverse scenarios that closely mimic real-world usage of the kit.
 3.2.7 Nitrate concentration measurement in tomato leaves
+Following trichome density measurements, we developed a systematic process for extracting and analyzing leaf nitrate concentrations.
+The tomato leaves were first frozen via a refrigerant (HFO-1234ze) to disrupt cell walls through freezing, facilitating subsequent juice extraction.
+The leaves were placed in plastic bags with their zippers partially closed, leaving a small opening for refrigerant introduction.
+During the freezing process, the evaporating refrigerant generated dry gas that displaced the moist air from the bags, thus preventing frost formation.
+[…]
 […]
@@
 3.3.1 Configuration of the data analysis environment
-The analytic environment was built using a Linux-based virtualization technology to ensure consistency and reproducibility of the analytical results across different computer systems, regardless of the specific hardware specifications.
-The analysis was conducted using Python [pythonsoftwarefoundationpython] version 3.10.
-Table 3 provides a comprehensive list of the Python library versions and operating system versions utilized in this study.
-The opencv-contrib-python, imbalanced-learn [scikitlearncontribimbalancedlearn], lightgbm [kguolinandmicrosoftresearchlightgbm], shap [slundbergandsleeshap], and ExifRead [sianareexifread] libraries were utilized for image processing, dataset preprocessing, machine learning model training, model interpretation, and metadata extraction from image data, respectively.
+The analytic environment was built via container virtualization technology to ensure the consistency and reproducibility of the analytical results across different computer systems, regardless of the specific hardware specifications.
+The analysis was conducted via Python [pythonsoftwarefoundationpython] version 3.10; Table 3 provides a comprehensive list of the Python library versions and operating system versions utilized in this study.
+We used opencv-contrib-python for image processing, imbalanced-learn [scikitlearncontribimbalancedlearn] for dataset preprocessing, lightgbm [kguolinandmicrosoftresearchlightgbm] for machine learning model training, shap [slundbergandsleeshap] for model interpretation, and ExifRead [sianareexifread] for metadata extraction from image data.
 3.3.2 Statistical analysis of nitrate ion concentrations
-We employed a series of nonparametric tests to statistically validate the observations of nitrate ion concentrations across different fertilizer levels.
-First, we used the Kruskal-Wallis test to examine the overall differences in distribution among the three fertilizer levels.
-Next, we conducted pairwise comparisons via Mann-Whitney U tests to assess the differences between each pair of fertilizer levels.
-To account for multiple comparisons, we applied the Bonferroni correction to the p values obtained from the Mann-Whitney U tests.
+To statistically validate the observations of nitrate ion concentrations across different fertilizer levels, we employed a series of nonparametric tests.
+First, the Kruskal-Wallis test was used to examine the overall distribution differences among the three fertilizer levels.
+Next, pairwise Mann-Whitney U tests were conducted to assess the differences between each pair of fertilizer levels, with Bonferroni correction applied to the resulting p values to account for multiple comparisons.
 3.3.3 Multicollinearity and variable interaction assessment
-Before building the model, we investigated the potential presence of multicollinearity and interaction effects among the predictor variables.
-Multicollinearity occurs when predictor variables are strongly intercorrelated, which can result in unstable and unreliable regression coefficient estimates.
-To detect the presence of multicollinearity, we utilized the variance inflation factor (VIF), which measures the extent to which the variance of a regression coefficient estimate increases owing to the correlation among predictor variables.
-A VIF value surpassing the cutoff point of 10 is commonly considered a strong indication of problematic multicollinearity [zuur2010a].
-To enhance the model interpretability and alleviate the consequences of multicollinearity, we removed variables with VIF values surpassing the threshold of 10 from further analyses.
-Moreover, we explored the possibility of interaction effects among the predictor variables, which arises when the level of a predictor variable influences the impact of another predictor variable on the response variable.
-To uncover potential interactions between features, we conducted a stratified analysis involving dividing the data into subgroups on the basis of the levels of a suspected interacting variable and assessing relationships between the predictor and response variables within each subgroup.
-Specifically, we investigated how image resolution might modulate the association between trichome density and nitrate concentration in tomato leaves.
+Before model construction, we conducted a comprehensive analysis of the relationships among the predictor variables, specifically examining multicollinearity and interaction effects.
+Multicollinearity, which occurs when predictor variables are strongly intercorrelated, can lead to unstable and unreliable regression coefficient estimates.
+We assessed multicollinearity via the variance inflation factor (VIF), which measures the extent to which the variance of a regression coefficient estimate increases owing to the correlation among predictor variables.
+Following common practice, variables with VIF values exceeding 10 were considered to indicate problematic multicollinearity [zuur2010a], and they were removed from subsequent analyses to enhance model interpretability.
+We also examined the interaction effects among the predictor variables, which arise when the level of a predictor variable influences the impact of another predictor variable on the response variable.
+Our investigation focused particularly on how image resolution might modulate the relationship between trichome density and leaf nitrate concentration.
+This analysis was conducted through stratification, dividing the data into subgroups on the basis of levels of suspected interacting variables and examining the predictor-response relationships within each subgroup.
 3.4 Analysis of Tomato Fertilization Requirements based on Leaf Trichome Density
 3.4.1 Problem formulation and objectives
-The primary objective of this study was to develop a kit that is suitable for practical use in the field.
-By leveraging existing knowledge on the correlation between leaf nitrogen-related indicators, we aimed to address a crucial question for farmers, which is affected by the fertilizer stress level and trichome density: Do my tomato plants in the field require additional fertilization?
+The primary objective of this study was to develop a kit suitable for practical use in the field.
+By leveraging existing knowledge about the correlation between trichome density and leaf nitrogen-related indicators, we aimed to address a crucial question that farmers face, which is affected by the fertilizer stress level and trichome density: Do my tomato plants in the field require additional fertilization?
 3.4.2 Model design and variable selection
-Consequently, we formulated a binary classification problem to determine the need for fertilization on the basis of leaf nitrate ion concentrations.
-The threshold for the leaf nitrate ion concentration was set as a parameter.
-Importantly, there is no universally ideal level of fertilizer stress for tomatoes.
-This level can vary depending on the environmental conditions and specific goals of the cultivator.
-Recognizing this variability, our study did not set a specific nitrate ion threshold.
-Instead, we parameterized the threshold for the leaf nitrate ion concentration in our model.
-This threshold was used to determine whether additional fertilization was necessary: If the concentration exceeded the threshold, no additional fertilization was needed, whereas concentrations below the threshold indicated a need for fertilization.
-This binary decision was then used as the target variable in our model.
-For the explanatory variables in this model, we used the NND, image resolution and exposure time.
-We applied standard scaling to these explanatory variables to ensure that all variables were on a comparable scale and to improve the model’s performance.
-This scaling technique transforms the variables so that their mean is zero and their standard deviation is one.
-In our cross-validation process, the data were divided into training and test sets.
-Importantly, the primary focus of this research was to develop a kit suitable for practical field use, with an emphasis on architectural-level improvements.
-Specifically, a quantitative method for characterizing tomatoes and refinement of measurement methods were selected.
-Therefore, we did not employ hyperparameter tuning using a validation dataset.
-Instead, we used default values for the model’s hyperparameters.
-This decision may lead to underfitting or overfitting due to insufficient control over model complexity, potentially resulting in reduced generalization performance and missed opportunities for theoretical performance analysis and optimization [mohri2018foundations].
-However, we deemed these potential drawbacks acceptable, given that our primary objective was to prioritize the development of a practical and accessible kit for field use, rather than achieving the highest possible model performance.
+We formulated a binary classification problem to determine the need for fertilization on the basis of leaf nitrate ion concentrations.
+Given that there is no universally ideal level of fertilizer stress for tomatoes, as optimal levels vary with environmental conditions and cultivator goals, we designed our model with a parameterized nitrate ion concentration threshold.
+Using this threshold, our model makes binary decisions.
+Concentrations above the threshold indicate that no additional fertilization is needed, whereas concentrations below the threshold suggest that fertilization is needed.
+The model used three explanatory variables: NND, image resolution, and exposure time.
+These variables underwent standard scaling to ensure comparable scales and to improve model performance.
+This scaling technique transformed each variable to have a mean of zero and a standard deviation of one.
+For model evaluation, we employed a cross-validation process that divided the data into training and test sets.
+The primary focus of this research was to develop a kit suitable for practical field use, emphasizing improvements at the architecture level, that is, the system’s fundamental structure and design.
+We aimed to develop objective methods for tomato characterization rather than optimizing model performance.
+Therefore, we used the model’s default hyperparameter values instead of implementing extensive tuning with a validation dataset.
+While this decision might lead to underfitting or overfitting due to insufficient control over model complexity, potentially reducing generalization performance and limiting the opportunities for theoretical performance analysis [mohri2018foundations], we considered these trade-offs to be acceptable, as our priority was creating an accessible and practical tool rather than maximizing theoretical model performance.
 3.4.3 Cross-validation strategy
-To evaluate the model’s generalization performance, we employed leave-one-out cross-validation (LOOCV) for each compound leaf.
-This method was chosen because of the relatively small-scale dataset, derived from 40 compound leaves harvested from 20 tomato plants from the three intervention groups.
-These compound leaves yielded 300 leaflets, and we analyzed 6,855 images taken of these leaflets.
-Given this dataset structure, LOOCV was deemed suitable for minimizing bias from specific samples while maximizing the use of our limited but diverse data points.
-Furthermore, our feature extraction process enables learning with tabular data, maintaining low computational costs despite the iterative nature of LOOCV.
-In each iteration of the LOOCV process, a single unique compound leaf was selected.
-The data derived from this selected compound leaf served as the test set, whereas the data from the remaining compound leaves constituted the training set.
+Our model validation approach was determined by the hierarchical structure of our dataset.
+We analyzed 40 compound leaves harvested from 20 tomato plants across three intervention groups, which yielded 300 leaflets and, ultimately, 6,855 analyzed images.
+Given this relatively small-scale but diverse dataset, we employed leave-one-out cross-validation (LOOCV) at the compound leaf level to evaluate the model’s generalization performance.
+This approach was selected to minimize bias from specific samples while maximizing the use of our limited data points.
+LOOCV implementation was supported by our feature extraction process, which enables learning with tabular data while maintaining low computational costs despite the iterative nature of LOOCV.
+In each iteration, we selected one compound leaf as the test set and used the data from all remaining compound leaves for training.
 3.4.4 Evaluation methodology
-Because the kit users can easily and cost-effectively perform multiple measurements, we evaluated the model’s predictive performance via a method that averages multiple predictions obtained from multiple images into a single prediction.
-We randomly extracted 1 to 25 rows of data from the test set and input them into the model.
-This method simulates scenarios where users capture and utilize 1 to 25 images simultaneously for prediction, which allows us to assess the model’s performance under various measurement intensities.
-This assessment enables us to understand how the model’s performance improves with an increasing number of measurements.
-To ensure a comprehensive evaluation, we aggregated the predictions and actual observations from all cross-validation iterations: This combined data is used to calculate various performance metrics.
+Our evaluation strategy was designed to reflect the practical usage of our kit, where users can easily perform multiple measurements cost-effectively.
+We developed a method to assess model performance by averaging the predictions from multiple images, simulating realistic usage scenarios where farmers might take between 1 and 25 images for a single assessment.
+This approach was implemented by randomly extracting 1 to 25 rows of data from the test set and inputting them into the model for prediction.
+By examining the model’s performance across different numbers of measurements, we could evaluate how prediction accuracy improves with increased measurement intensity.
+For a comprehensive performance assessment, we aggregated predictions and actual observations across all cross-validation iterations to calculate various performance metrics.
+This aggregated evaluation approach provides insights into the model’s reliability under different measurement scenarios while maintaining practical relevance to field applications.
 3.4.5 Parameter range selection
-To effectively visualize the analysis results, the range of nitrate ion concentrations and the number of images used for evaluation were narrowed to specific ranges or values.
-This refinement process was guided by practical considerations and domain knowledge.
-With respect to the nitrate ion concentration threshold, setting a single universal value is inappropriate because of the variability in agricultural conditions.
-We could define a relevant range on the basis of domain knowledge.
-Two key factors influenced this decision.
-First, nonlinear models such as LightGBM, which we used for modeling, tend to produce unstable predictions near or outside the boundaries of the training data.
-Second, agricultural workers and related economic entities primarily sustain their livelihoods through crop production revenues.
-Consequently, it is crucial to avoid excessive fertilizer stress, which could lead to decreased profitability.
-Considering these points, we decided to exclude the low nitrate concentration region from the threshold setting.
-Such a scenario is unlikely to be used in a practical application of the kit.
-As our kit allows for multiple measurements to be easily and cost-effectively performed, we set the upper limit for the number of images used in the evaluation to 25.
-This decision balances the need for comprehensive data with practical considerations of time and resource efficiency in real-world applications.
-Based on these considerations, we defined our range of interest for nitrate ion concentrations from the second quartile to the third quartile of the distribution obtained from our experiments.
-For visualizations requiring specific parameter settings, we used the third quartile value as the nitrate ion concentration threshold and set the number of images for evaluation to 25.
-We refer to this combination of parameter settings as the "representative parameter setting".
-This approach allowed us to evaluate and visualize the model’s performance within a practical and statistically meaningful range.
-To ensure comprehensive reporting, performance evaluations using all the parameters, including those outside the defined range of interest, are presented in the appendix.
-3.4.6 Data analysis workflow
-Algorithm 6 presents a schematic overview of the data analysis workflow. The main processes of model training and evaluation are described in detail below.
+Our predictive model fundamentally depends on two key parameters that govern its operation.
+These parameters are the nitrate ion concentration threshold and the number of images used for evaluation.
+To effectively visualize and analyze the results, we carefully defined specific ranges for these parameters, considering both practical utility and model reliability.
+This selection was particularly crucial, as it directly impacts our model’s ability to provide actionable recommendations under the inherent variability of agricultural conditions.
+The determination of suitable nitrate ion concentration thresholds was influenced by two critical factors from our domain expertise.
+First, nonlinear models such as light gradient boosting machine (LightGBM) typically produce unstable predictions near or outside their training data boundaries.
+Second, agricultural workers and related actors sustain their livelihoods through crop production revenues, where excessive fertilizer stress directly leads to decreased crop productivity and, consequently, reduced profitability.
+Based on these considerations, we excluded the low nitrate concentration region from the threshold settings, as such conditions would be impractical and economically unsustainable in real-world agricultural applications.
+For the number of evaluation images, we balanced measurement comprehensiveness with practical field constraints.
+While our kit enables multiple measurements to be performed easily and cost-effectively, we established an upper limit of 25 images per evaluation.
+This limit provides sufficient data for reliable analysis while remaining feasible within the time and resource efficiency requirements.
+We defined our primary range of interest for nitrate ion concentrations from the second to third quartiles of our experimental distribution.
+For specific visualizations, we adopted a "representative parameter setting" using the third quartile value as the nitrate ion concentration threshold and 25 images for evaluation, providing a practical and statistically robust foundation that reflects real-world agricultural conditions.
+Comprehensive performance evaluations across all parameter ranges, including those outside our defined range of interest, are presented in the appendix, along with the detailed data analysis workflow (Algorithm 7).
+3.4.6 Feature preprocessing and label setting
+Our modeling objective was to determine the need for additional fertilization by comparing the nitrate ion concentration, estimated from trichome density measurements in images, with a predefined threshold.
+We assigned binary labels to leaves based on their nitrate ion concentration: 0 for concentrations below the threshold, indicating no need for additional fertilization, and 1 for concentrations above the threshold, indicating that fertilization was needed.
+3.4.7 Model training
+Prior to modeling, we partitioned our dataset into training and testing sets based on compound leaves for binary classification.
+To address the imbalances in both image numbers across compound leaves and leaflets and in class labels, we applied the synthetic minority oversampling technique (SMOTE) [haixiang2017learning, chawla2002smote] to oversample the minority class.
+For classification, we employed LightGBM, a decision tree-based approach that combines multiple trees through gradient boosting techniques [ke2017lightgbm].
+In this approach, each decision tree iteratively improves the model’s predictive accuracy by learning from the residuals of previous trees.
+The regression predictions $z$ were calculated via Equation 4, where $\mathrm{f}_{k}(\bm{x})$ represents the output of the $k$-th decision tree, with $K$ denoting the total number of trees and $\bm{x}$ representing the input feature vector.
+$z=\sum_{k=1}^{K}\mathrm{f}_{k}(\bm{x})$ (4)
+These predictions were transformed into probabilities through the sigmoid function shown in Equation 5, which maps values to probabilities between 0 and 1, representing the likelihood of the input belonging to class 1.
+This transformation provides farmers with interpretable probabilities of nutrient deficiency, facilitating informed decision making about additional fertilization.
+The final classification ($\hat{Y}$) was determined from the computed probability ($\hat{y}$) following the decision rule in Equation 6.
+The model was trained via LightGBM’s default parameters, with a learning rate of 0.1 and 100 boosting rounds.
+$\hat{y}=\frac{1}{1+\mathrm{e}^{-z}}$ (5)
+$\hat{Y}=\begin{cases}1&\text{if }\hat{y}\geq 0.5\\ 0&\text{if }\hat{y}<0.5\end{cases}$ (6)
+3.4.8 Model evaluation and feature contribution
+We evaluated our model by comparing predicted labels $\hat{Y}$ with the true labels via multiple performance metrics.
+The area under the precision-recall curve (AUC-PR) served as the primary evaluation metric during training, as defined by Equation 11.
+For comprehensive model assessment based on test data, we employed precision, recall, the $\mathrm{F}_{1}\text{ Score}$, the area under the receiver operating characteristic curve (AUC-ROC), the area under the precision recall curve (AUC-PR), mean AUC-ROC (mROC), and mean AUC-PR (mPR), as described in Equations 7-13.
+[…]
+Precision=True PositiveTrue Positive+False $\text{Precision}=\frac{\text{True Positive}}{\text{True Positive}+\text{False Positive}}$ (7)
+Recall=True PositiveTrue Positive+False $\text{Recall}=\frac{\text{True Positive}}{\text{True Positive}+\text{False Negative}}$ (8)
+$\mathrm{F}_{1}\text{ Score}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (9)
+[…]
+[…]
+[…]
+$\text{mROC}=\frac{1}{M}\sum_{c=1600}^{1900}\text{AUC-ROC}_{c}$ (12)
+$\text{mPR}=\frac{1}{M}\sum_{c=1600}^{1900}\text{AUC-PR}_{c}$ (13)
+3.5 Analysis of Trichome Density in Relation to Fertilizer Stress
+3.5.1 Research objectives and scope
+This research primarily aims to develop practical agricultural tools for farmers to address cultivation challenges by investigating the relationship between fertilizer stress and trichome density in tomato plants.
+Our approach leverages reverse inference, utilizing a known natural phenomenon where leaf trichome density changes in response to levels of fertilizer stress.
+While our experimental design primarily focuses on creating practical tools for farmers to assess fertilizer stress through observed trichome density patterns, it also provides an opportunity to contribute, as a secondary outcome, to the broader scientific understanding of this biological response in tomato plants.
+3.5.2 Analysis methodology
+To quantitatively investigate the relationship between fertilizer stress and trichome density, we employed a machine learning regression approach as a forward inference step, addressing the interests of researchers studying the natural phenomenon itself, rather than its reverse inference application for farmers.
+The analysis was designed to investigate how much of the variation in leaflet trichome density under fertilizer stress conditions can be explained by fertilizer stress indicators.
+Our model’s prediction was built upon three key features.
+We used nitrate ion concentration as the primary predictor, while image resolution and exposure time served as moderator variables for image acquisition parameters.
+We implemented the LightGBM algorithm with default hyperparameters and configured it with an L2 loss objective function, 100 boosting rounds, and a learning rate of 0.1.
+The model’s performance and generalizability were rigorously evaluated via LOOCV, where each unique leaflet iteratively served as the test set.
+We applied standard scaling to both the feature variables and the target variable before training.
+Each LOOCV fold involved training a LightGBM model on the scaled training data to predict the NND, followed by inverse transformation to the original scale for evaluation.
+Model performance was assessed via multiple metrics, including the root mean square error ($\mathrm{RMSE}$), coefficient of determination ($R_{2}$), and Pearson correlation coefficient ($r$).
+4 Results and Discussion
+4.1 Leaf Nitrate Concentration as a Lagging Indicator of Fertilizer Stress
+The distribution of nitrate ion concentrations in compound leaves, measured 15 cm from the apical meristem across three fertilizer treatment levels (high, medium, and low), is presented in Figure 10.
+Notably, higher fertilizer application rates corresponded to elevated nitrate ion concentrations.
+Statistical analysis via the Kruskal-Wallis test revealed significant differences among the three groups ($p<0.05$).
+Pairwise comparisons via Mann-Whitney U tests with Bonferroni correction revealed significant differences between the high-low and medium-low groups ($\text{adjusted}\,p<0.05$), whereas the high-medium comparison revealed no significant difference ($\text{adjusted}\,p=0.52$).
+The absence of a significant difference between the high- and medium-intensity fertilizer treatments suggests a diminishing returns effect.
+These findings indicate that while the maximum yield is greater in the high-intensity fertilizer treatment than in the low-intensity treatment, additional fertilizer inputs beyond the medium level yield increasingly smaller gains.
+While these results indicate that compound leaf nitrate ion concentration can serve as a reliable indicator of fertilizer stress in tomatoes, particularly for distinguishing between low and high fertilizer levels, the relationship shows notable variability.
+As illustrated in Figure 10, the overlapping distributions between the high and medium treatments highlight this variability.
+These findings emphasize the importance of directly measuring plant nitrate uptake rather than relying solely on fertilizer application rates to accurately evaluate fertilizer stress.
+Direct measurement of nutrient uptake provides more comprehensive insight into nutrient availability by accounting for various factors beyond fertilizer application that may influence nutrient absorption.
+Figure 10: Distribution of nitrate concentrations in compound leaves at different fertilizer levels.
+4.2 Leaf Nitrate Concentration as a Leading Indicator of Fruit Yield
+Regarding the relationship between fruit yield and compound leaf nitrate concentration, the results demonstrated a positive correlation, as shown in Figure 11, indicating that leaf nitrate ion concentration can serve as a yield indicator.
+While crop yield generally increases with fertilizer absorption, excessive application beyond a critical threshold can reduce yields [andersen1999relationships, locascio1997nitrogen, coltman1988yields].
+Our experimental results suggest that fertilizer application remained below this critical threshold, indicating that the tomatoes experienced stress from insufficient fertilization.
+Figure 11: Relationship between the nitrate ion concentration in compound leaves and yield.
+[…]
+Figure 12: Cumulative fruit yield in the experimental setup.
+4.3 Handling of Multicollinearity and Variable Interaction in Data Analysis
+Our initial statistical analysis revealed significant multicollinearity among the explanatory variables (Farrar-Glauber test, $p<0.001$) [farrar1967multicollinearity], indicating a statistically significant linear dependence among them.
+An examination of VIFs showed that ISO speed ratings presented an exceptionally high VIF of 47.80, far exceeding the conventional threshold of 10 (Table 4).
+This multicollinearity arose from the camera’s automatic exposure (AE) function responding to image brightness, which was influenced by light reflection from trichomes that appeared as white points.
+As trichome density varied with fertilizer stress, the resulting changes in image brightness triggered automatic adjustments in ISO sensitivity.
+This mechanism positioned ISO sensitivity as a mediator variable, not through direct measurement but rather by reflecting the camera’s response to image characteristics, which created an unintended correlation with both the variable of interest (trichome density) and the target variable (fertilizer stress).
+Table 4: Examination of Multicollinearity through the VIF
+Feature VIF
+Nearest Neighbor Distance 43.60
+Resolution 11.03
+Exposure Time 4.98
+ISO Speed Ratings 47.80
+Although multicollinearity may not significantly impact the model’s predictive performance during cross-validation, it can substantially affect the interpretation of individual variable importance, leading to unstable and unreliable coefficient estimates.
+Given our study’s focus on both generalization performance and model interpretability, we excluded ISO speed ratings from subsequent analyses.
+After this exclusion, no remaining variables had VIF values exceeding 10 (Table 5), resolving the multicollinearity concerns.
+Table 5: VIF Value after Dimensionality Reduction
+Feature VIF
+Nearest Neighbor Distance 8.63
+Resolution 6.49
+Exposure Time 3.10
+The resolution value, however, demonstrated interactions and intercept shifts in the relationship between trichome density and nitrate concentration, with varying coefficients and intercepts in the regression equation, as shown in Table 6 and Figure 13.
+Resolution variations can arise from multiple factors, including camera performance and the distance between the camera sensor and measurement paper.
+In this study, as all images were captured via the same camera device, resolution primarily served as an indicator of the camera-to-paper distance.
+To avoid omitted variable bias, which could lead to a misinterpretation of the effects of one variable as those of another, we included resolution as a key variable in the model.
+As shown in Table 6, higher resolution values were associated with larger coefficients in our regression analysis, suggesting that optimal trichome detection in our computer vision application requires close camera proximity and high-resolution sensors.
+Based on these analyses, our final model included the NND, resolution, and exposure time as predictor variables, each of which had independent effects on the response variable.
+Table 6: Interaction when Stratified Based on Resolution
+Percentile Resolution (pixel) Equation
+Higher than 85th $>$ $1.08\times 10^{6}$ $y=3.06\times 10^{-2}x+170$
+15th to 85th $1.08\times 10^{6}-5.56\times 10^{5}$ $y=1.90\times 10^{-2}x+205$
+Lower than 15th $<$ $5.56\times 10^{5}$ $y=1.28\times 10^{-2}x+254$
+Figure 13: Effect of resolution on the relationship between trichome density and nitrate concentration in leaves.
+4.4 Performance of the Proposed Method
+Our LightGBM model demonstrated consistent improvement in performance over 100 boosting rounds via default parameters, as evidenced by the gradual increase in AUC-PR values for training sets shown in Figure 14, indicating that the model was adequately trained.
+For comprehensive evaluation, we established a representative scenario using a nitrate concentration threshold of 1,783 ppm with 25 images captured from a single compound leaf.
+Under this scenario, the model obtained notable performance metrics with an mROC of 0.64 and an mPR of 0.82, despite variations in measurement data due to different optical conditions, enabling accurate assessment of additional fertilization needs.
+Table 7 presents the model’s predictive performance across different nitrate concentration thresholds, showing well-balanced precision and recall values.
+Figure 14: Learning curve of the LightGBM model.
+Table 7: Predictive Performance in Terms of the Need for Additional Fertilizer using LOOCV at Various Nitrate Ion (ppm) Thresholds when Using the 25 Image Features
+Evaluation Metrics 1600 1750 1900
+Precision 0.68 0.76 0.79
+Recall 0.66 0.76 0.79
+F1 Score 0.67 0.76 0.79
+PR AUC 0.76 0.84 0.87
+ROC AUC 0.64 0.68 0.58
+The model’s balanced classification performance is evidenced by the confusion matrix in Figure 15, which shows no substantial bias toward either false positives or false negatives.
+This balanced performance is further supported by the F1 scores, AUC-PR values, and AUC-ROC values presented in Table 7 and visualized through the PR and ROC curves in Figures 16 and 17, respectively.
+Compared with a baseline majority class predictor model, our proposed model demonstrated superior performance across all evaluation metrics defined in Equations 7-11 (Table 8).
+Figure 15: Confusion matrix for classification outcomes during LOOCV in a representative parameter setting.
+Figure 16: Precision-recall (PR) curve analysis of the model during LOOCV in a representative parameter setting.
+Figure 17: Receiver operating characteristic (ROC) curve analysis of the model during LOOCV in a representative parameter setting.
+Table 8: Comparison of Predictive Performance between the Proposed Model and the Baseline Model in a Representative Parameter Setting
+Model Precision Recall F1 Score ROC AUC PR AUC
+Proposed 0.79 0.78 0.79 0.81 0.87
+Baseline 0.62 0.49 0.55 0.50 0.72
+Analysis of the model’s practical applicability revealed that even with a single input image, the model obtains a mean AUC-PR (mPR) of 0.80, with performance improving as the number of input images increases (Figure 18).
+The feature contribution analysis using SHAP values (Figure 19) identified the NND as the most significant contributor to the model’s predictions, showing the highest mean absolute SHAP value across the dataset.
+Figure 18: Model prediction performance when various thresholds and numbers of images are set.
+Figure 19: Feature contribution during LOOCV with a representative model setting.
+The performance metrics demonstrate the practical viability of a smartphone-based trichome density measurement system.
+With just a smartphone camera, measurement sheets and analysis software, our method automates traditional microscope-based counting for the rapid assessment of tomato plant nutrient status.
+4.5 Effect of Fertilizer on Trichome Density
+Our quantitative evaluation of the relationship between tomato leaf trichome density and nitrogen nutritional status employed a prediction model using the NND as a measure of trichome density.
+As shown in Figure 20, the model incorporated nitrate ion concentration as an explanatory variable, while image resolution and exposure time served as moderating variables.
+This combination of variables explained 62.48% of the variation in trichome density ( $R^{2}$ = 0.63) observed in leaflets from our fertilizer intervention experiments.
+The remaining unexplained variation (37.52%) suggests opportunities for future research to investigate additional influencing factors.
+Importantly, while our study provides evidence supporting the relationship between fertilizer stress and trichome density, our approach was primarily data driven rather than being focused on rigorous hypothesis testing for interpreting natural phenomena in complex systems.
+Unlike traditional hypothesis-based research, which typically begins with specific hypotheses and theories for deductive testing [glass2008brief], our study generated hypotheses and models inductively from the patterns and relationships within the dataset, guided by basic hypotheses, as illustrated in Figure 3.
+While potentially introducing inherent biases when extended to the interpretation of underlying biological phenomena, this methodology provides adequate evidence for technological validation in the context of kit development.
+These potential methodological biases are not considered fatal in our context because previous hypothesis-based research has already established the quantitative relationship between tomato leaf trichome density and nitrogen nutritional status [hoffland2000nitrogen].
+Additionally, our actual task was to capture established general trends for practical applications rather than identifying subtle biological differences in complex plant systems where hypothesis-based research would be prioritized.
+Our findings support the influence of fertilizer stress on trichome density in young leaves, although further investigations could enhance model performance and our understanding of this relationship.
+Figure 20: Modeling of the NND via nitrate concentration, resolution, and exposure time.
+4.6 Method Performance and Limitations
+Tomatoes pose a challenge in detecting fertilizer stress through noncontact optical measurements of young leaf chlorophyll.
+This challenge is due to the nonlinear relationship between optical measurements and leaf nitrogen content, which distinguishes tomatoes from other crops [padilla2018proximal, gianquinto2011a].
+To address this problem, our study focused on trichomes, which are densely packed hair-like structures on leaf surfaces that have been less explored in previous research.
+We investigated the feasibility of detecting fertilizer stress by developing a device that utilizes trichome density as a primary predictor.
+Our findings suggest the possibility of overcoming the limitations of conventional noncontact optical methods in assessing tomato fertilizer conditions, although our diagnostic kits have certain performance limitations.
+To validate this approach, we conducted experiments using a well-designed setup that simulated different levels of fertilizer stress common in agricultural production.
+Our cost-effective hydroponic cultivation system with plastic buckets achieved yields comparable to those of commercial cultivation, suggesting that agricultural conditions were effectively simulated.
+However, we acknowledge that not all farmers may be able to establish optimal growing conditions due to equipment costs.
+As a result, environmental stresses beyond fertilizer deficiency could decrease our method’s predictive performance.
+4.7 Practical Implications
+Traditional nutrient status assessment through visual inspection of leaf color, size, texture, and overall plant health heavily relies on individual farmers’ experience and motivation, leading to considerable variations and standardization challenges.
+Even when specific machines or systems are utilized for nutrient status assessment, the combinations of phenotypic indicators and sensor types create endless possibilities for evaluating utility across multiple dimensions, including the complexity of data processing, the prerequisites for measurement, operational complexity, time requirements, and detection latency from stress onset.
+This variability creates situations where certain measurement methods work well under specific conditions, making it difficult to evaluate the effectiveness of diagnostic kits through predictive accuracy comparisons alone.
+More comprehensive validation using lagging indicators, particularly financial metrics, is necessary to assess practical applicability in real-world settings.
+While various commercial nutrient diagnostic kits exist, their high costs and application challenges with delicate tissues such as young tomato leaves, which are typically sensitive to nutrient deficiencies and crucial for plant growth and development, represent one aspect of the diverse factors affecting their adoption in practical farming settings.
+In response, we propose a novel, simple, and cost-effective method specifically designed for assessing nutrient status in these delicate tissues.
+4.8 Future Research
+The potential applications of our trichome-based method extend beyond tomatoes to other agriculturally important species with abundant trichomes, including watermelon (Citrullus lanatus), eggplant (Solanum melongena), tobacco (Nicotiana tabacum), soybean (Glycine max), cannabis (Cannabis sativa), and potato (Solanum tuberosum).
+This broader applicability could improve real-time nutrient management across various crops, potentially optimizing fertilizer application, reducing waste, and lowering environmental impact.
+These improvements would ultimately contribute to more sustainable agricultural practices.
+Further research is needed to validate the effectiveness of this method across different species and to investigate its potential integration with other precision agriculture technologies for comprehensive nutrient management strategies.
+5 Conclusions
+This study focused on developing a novel method for detecting fertilizer stress in tomato plants by measuring the density of trichomes on the leaves of these plants, including the setup and evaluation of necessary experimental conditions.
+Our developed diagnostic kit enables noninvasive and early detection of fertilizer stress in plants without requiring customized sensing devices.
+This universal approach provides accurate nutrient status assessment at a lower cost than traditional devices.
+The findings present a novel approach for designing diagnostic devices for plant fertilizer stress detection by utilizing the leaf surface structure, offering potential advancements in agricultural practice.
+Appendix
+Appendix A Simulation Demonstrating the Robustness of the Nearest Neighbor Distance
+This appendix presents a simulation demonstrating the robustness of the nearest neighbor distance (NND) as a density metric compared with the conventional point count per unit area.
+The simulation involves generating a set of points within a defined area and simulating damage by removing a portion of these points to compare the changes in both the NND and point count measurements.
+To mimic the random spatial distribution of trichomes on a leaf surface, we employed a Poisson point process for point generation, as this process effectively models random events occurring independently in a given space or time.
+This approach is particularly suitable for simulating trichome distribution, as trichomes are expected to emerge randomly on the leaf surface.
+The detailed steps of this simulation are outlined in Algorithm 6.
+Algorithm 6 Trichome distribution simulation
+1:$\lambda$: Expected number of events (trichomes)
+2:$A$: Total area of the region
+3:$S$: Random seed value
+4:$P$: Set of points representing trichome locations
+5:$P^{\prime}$: Set of points after simulated damage
+6:function SetSeed($S$)
+7:Set a random seed to $S$
+8:end function
+9:function PoissonRandom($\lambda$)
+10:Generate random numbers from the Poisson distribution return number of points
+11:end function
+12:function UniformRandom($a,b$)
+13:Generate random numbers from a uniform distribution return random number between $a$ and $b$
+14:end function
+15:function GeneratePoints($N$)
+16:Generate $N$ random points in the area return set of points $P$
+17:end function
+18:function SimulateDamage($P$)
+19:Remove points with $x>500$ return set of remaining points $P^{\prime}$
+20:end function
+21:
+22:Main Execution
+23:SetSeed($S$)
+24:$N\leftarrow\textsc{PoissonRandom}(\lambda)$
+25:$P\leftarrow\textsc{GeneratePoints}(N)$
+26:$P^{\prime}\leftarrow\textsc{SimulateDamage}(P)$
+Figure 21 shows an example of the point distributions before and after simulated damage.
+For this demonstration, we randomly generated points via the Poisson process within a 1000 $\times$ 1000 area.
+Damage was simulated by removing points within the right half of the area, marked by the "damage boundary" line.
+Figure 22 presents the histograms of NNDs before and after this simulated damage.
+Figure 21: Point distributions before and after simulated damage.
+Figure 22: Histogram of nearest neighbor distances.
+The simulation results demonstrate that the NND is more robust to point removal than the conventional density metric based on the point count per unit area.
+Figure 23 compares the NND and point count before and after simulated damage.
+Although both metrics changed after damage, the change in the NND was smaller than that in the point count.
+To validate this observation statistically, we performed a Wilcoxon signed-rank test to compare the absolute change rates of the NND and the point count after damage, as shown in Figure 24.
+The test revealed a statistically significant difference between the two metrics ($p<0.001$), supporting the hypothesis that the NND is less affected by localized point loss than the point count per unit area.
+This finding demonstrates the advantage of using the NND as a density metric for scenarios where localized point loss might occur, such as trichome detachment from leaf surfaces.
+Figure 23: Comparison of NND and point count after simulated damage.
+Figure 24: Comparison of NND and point count absolute change rates after simulated damage.
+Appendix B Data analysis workflow
+Algorithm 7 presents a comprehensive implementation of our tomato fertilization requirement prediction system.
+The algorithm consists of four main functional components: label setting, class balance adjustment, model training, and evaluation.
+The label setting function (SetLabels) processes the raw nitrate concentration data into binary classification labels based on the predefined threshold.
+The SMOTE application function (ApplySMOTE) addresses class imbalance issues in the training data.
+The model training function (TrainModel) handles the core prediction task, while the evaluation function (Evaluate) assesses model performance.
+The main execution flow implements a leave-one-out cross-validation strategy, where data are iteratively partitioned by the compound leaf index.
+For each iteration, the algorithm extracts test data for the current compound leaf while using the remaining data for training.
+After applying SMOTE to balance the training data, the model is trained and evaluated, with the results accumulated for both the training and test sets.
+The algorithm tracks the true and predicted values throughout the process and computes SHAP values.
+Finally, it calculates comprehensive performance metrics across all iterations.
 Algorithm 7 Tomato fertilization requirement prediction
@@
 […]
-3.4.7 Feature preprocessing and label setting
-The objective of the modeling was to determine the need for additional fertilization by comparing the nitrate ion concentration, estimated from the trichome density in images, with a predefined threshold.
-Consequently, leaves were labeled 0 (indicating no need for additional fertilization) or 1 (indicating the need for additional fertilization) based on whether the nitrate ion concentration was below or above the threshold, respectively.
-3.4.8 Model training
-The dataset was partitioned into training and testing sets before the modeling process.
-The modeling task involved binary classification, with the data being divided based on each compound leaf.
-To address the imbalance in both the number of images between compound leaves and leaflets and the class labels, the minority class was oversampled via the synthetic minority oversampling technique (SMOTE)[haixiang2017learning, chawla2002smote].
-Modeling was performed via a light gradient boosting machine (LightGBM), a decision tree-based approach that employs gradient boosting techniques to combine multiple decision trees for learning purposes [haixiang2017learning].
-In LightGBM, each decision tree uses the residuals from the previous tree to improve the overall predictive accuracy of the model iteratively.
-The regression predictions $z$ were calculated via Equation 4 [ke2017lightgbm].
-$z=\sum_{k=1}^{K}f_{k}(x)$ (4)
-In Equation (4), $f_{k}(x)$ is defined as the output generated by the $k$-th decision tree.
-Here, $k$ represents the total number of decision trees, and $x$ corresponds to the input feature vector.
-The regression prediction $z$ is then passed through the sigmoid function, as shown in Equation 5, to calculate the probability $\hat{y}$.
-$\hat{y}=\frac{1}{1+e^{-z}}$ (5)
-The sigmoid function in Equation (5) maps the value of $z$ to a probability between 0 and 1, which is then interpreted as the likelihood of the input belonging to class 1.
-By applying the sigmoid function to the regression predictions $z$, as shown in Equation (5), we simulate a scenario in which farmers are provided with the predicted probabilities of nutrient deficiency, enabling them to make informed decisions about the application of additional fertilization to their fields.
-The final class label $\hat{Y}$ was determined based on the computed probability $\hat{y}$, according to the decision rule outlined in Equation 6.
-$\hat{Y}=\begin{cases}1&\text{if }\hat{y}\geq 0.5\\ 0&\text{if }\hat{y}<0.5\end{cases}$ (6)
-To train the classification model, the default parameters of the LightGBM library were employed, including a learning rate of 0.1 and 100 boosting rounds.
-3.4.9 Model evaluation and feature importance
-The predicted labels $\hat{Y}$ were compared to the true labels, and performance metrics were computed.
-The evaluation metric used during training was the area under the precision-recall curve (AUC-PR), as defined by Equation 11.
-To evaluate the performance of the trained model on the test data, several evaluation metrics were employed, including the precision, recall, $F_{1}$, area under the receiver operating characteristic curve (AUC-ROC), AUC-PR, and mean AUC-PR (mPR).
-These metrics are described by Equations 7-12.
-The precision, as defined by Equation 7, is the ratio of correctly predicted positive observations to the total number of predicted positives [hossin2015a, jeni2013facing].
-Precision=True PositiveTrue Positive+False $\text{Precision}=\frac{\text{True Positive}}{\text{True Positive}+\text{False Positive}}$ (7)
-Equation 8 defines recall as the proportion of accurately predicted positive observations in relation to the total number of true positives.
-This metric evaluates the model’s ability to identify all relevant positive instances.
-Recall=True PositiveTrue Positive+False $\text{Recall}=\frac{\text{True Positive}}{\text{True Positive}+\text{False Negative}}$ (8)
-Equation 9 represents the $F_{1}$ score, which is calculated as the harmonic mean of the precision and recall. This metric serves as a balanced measure of both metrics.
-$F_{1}Score=\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (9)
-Equation 10 represents the AUC-ROC, which is the area under the curve of the true positive rate versus the false positive rate.
-$t(f)$ denotes the true positive rate at a given false positive rate $f$.
-This measure effectively summarizes the model’s ability to discriminate between the positive and negative classes across all possible thresholds, providing a single scalar value that reflects the trade-off between sensitivity (true positive rate) and specificity ($1-false-positiverate$).
-[…]
-Equation 11 represents the AUC-PR, which defines the area under the precision versus recall curve.
-$p(r)$ represents the precision at a given recall value $r$.
-This metric provides a single scalar value that summarizes the trade-off between precision and recall across all possible thresholds.
-The AUC-PR is particularly useful for datasets with imbalanced class distributions, as it provides a more informative measure than does the AUC-ROC [jeni2013facing].
-[…]
-[…]
-$\text{mROC}=\frac{1}{m}\sum_{k=1600}^{1900}\text{AUC-ROC}_{k}$ (12)
-[…]
-$\text{mPR}=\frac{1}{m}\sum_{k=1600}^{1900}\text{AUC-PR}_{k}$ (13)
-3.5 Analysis of Trichome Density in Relation to Fertilizer Stress
-3.5.1 Research objectives and scope
-The primary objective of this research is to develop practical tools for farmers to address agricultural challenges.
-Specifically, fertilizer stress in tomato plants has been assessed.
-This assessment is achieved through reverse inference, utilizing a known natural phenomenon: the change in leaf trichome density in response to variations in fertilizer stress.
-This approach allows us to infer fertilizer stress levels on the basis of observed trichome density patterns.
-Although our experimental design is optimized for this applied goal, it also provides an opportunity to contribute, albeit secondarily, to the scientific understanding of the relationship between fertilizer stress and trichome density in tomato plants as a natural phenomenon.
-3.5.2 Quantitative analysis approach
-Leveraging this opportunity, we employed a machine learning approach with a regression model to investigate the quantitative relationship between fertilizer stress and trichome density in tomato plants.
-This analysis served as a forward inference step, establishing a predictive model for trichome density based on three features: nitrate ion concentration and two moderator variables of image acquisition: image resolution and exposure time.
-3.5.3 Model selection and configuration
-We implemented the LightGBM algorithm for our regression model, using default hyperparameters.
-Specifically, the model was configured with an L2 loss objective function, 100 boosting rounds, and a learning rate of 0.1.
-3.5.4 Cross-validation and data preprocessing
-To rigorously assess the model’s performance and generalizability, we implemented a LOOCV strategy.
-This approach involved iteratively using each unique leaflet as a test set, while the remaining data served as the training set.
-Prior to model training, we applied standard scaling to the feature variables and the target variable.
-For each LOOCV fold, we trained a LightGBM model on the scaled training data and used it to predict the NND for the test set.
-The predictions were then inverse-transformed to the original scale for evaluation.
-We assessed the model’s performance via several metrics, including the root mean square error ($RMSE$), Coefficient of determination ($R_{2}$) and the Pearson correlation coefficient ($R$).
-4 Results
-4.1 Nitrate Ion Concentrations in Compound Leaves as Lagging Indicators of Fertilizer Stress
-Figure 10 shows the distribution of nitrate ion concentrations in compound leaves located 15 cm from the apical meristem, measured across experimental plots subjected to three different fertilizer concentrations: high, intermediate, and low.
-Notably, higher fertilizer application rates corresponded to elevated nitrate ion concentrations within the compound leaves.
-Statistical analysis via the Kruskal-Wallis test yielded a p value < 0.05, indicating significant differences among the three groups.
-Subsequent pairwise comparisons via the Mann-Whitney U test, adjusted for multiple comparisons via the Bonferroni correction, revealed statistically significant differences between the high-low and intermediate-low groups ($\text{adjusted}\,p<0.05$).
-However, no statistically significant difference was observed between the high-intermediate subgroup ($\text{adjusted}\,p=0.52$).
-This lack of a significant difference between the high- and intermediate-level fertilizer treatments may reflect the trend of diminishing returns.
-These findings suggest that the maximum yield is greater in the high-intensity fertilizer treatment than in the low-intensity treatment, beyond which additional fertilizer inputs provide increasingly smaller yield gains.
-These results suggest that the compound leaf nitrate ion concentration can serve as a reliable indicator of fertilizer stress in tomatoes, particularly for distinguishing between low and high fertilizer levels.
-However, it is important to acknowledge the variability observed in the relationship between nitrate ion concentration and fertilizer level.
-The overlapping distributions and nonsignificant differences between the high- and intermediate-level fertilizer treatments are shown in Figure 10.
-This variability underscores the importance of directly assessing plant nitrate uptake rather than relying solely on fertilizer application rates to accurately evaluate fertilizer stress.
-Directly measuring plant nutrient uptake can provide a more comprehensive understanding of nutrient availability, accounting for factors beyond fertilizer application that may influence nutrient absorption.
-Figure 10: Distribution of nitrate concentrations in compound leaves at different fertilizer levels.
-4.2 Leaf Nitrate Concentration as a Leading Indicator of Fruit Yield
-Figure 11 shows a positive correlation between fruit yield and nitrate concentration in compound leaves.
-This finding indicates that the nitrate ion concentration in compound leaves can serve as an indicator of fruit yield.
-Generally, crop yield tends to increase with increased fertilizer absorption; however, fertilizer application beyond a certain threshold can reduce yields [andersen1999relationships, locascio1997nitrogen, coltman1988yields].
-The findings of this study indicate that fertilizer application in the experimental field fell below this critical threshold.
-These findings suggest that the tomatoes experienced stress due to insufficient fertilization.
-As detailed in the Parameter Range Selection section, various factors were considered when determining an appropriate nitrate concentration threshold for visualization purposes.
-Based on these considerations, we employed the third quartile value of 1,783 ppm, derived from the measured leaf nitrate concentrations, to illustrate the model’s performance in distinguishing between different yield potentials.
-Figure 11: Relationship between the nitrate ion concentration in compound leaves and yield.
-[…]
-Figure 12: Cumulative fruit yield in the experimental setup.
-4.3 Handling of Multicollinearity and Variable Interaction in Data Analysis
-The statistical model necessitates the exclusion of ISO speed ratings as a variable because of the manifestation of multicollinearity.
-A Farrar-Glauber test was conducted to assess multicollinearity, revealing a significant presence ($p<0.001$) [farrar1967multicollinearity].
-This result indicates a statistically significant linear dependence among the explanatory variables.
-Table 4 presents the variance inflation factor (VIF) for each variable.
-Notably, the ISO speed ratings exhibit an exceptionally high VIF of 47.80, substantially exceeding the conventional threshold of 10 and thus confirming severe multicollinearity.
-Table 4: Examination of Multicollinearity through the VIF
-Feature VIF
-Nearest Neighbor Distance 43.60
-Resolution 11.03
-Exposure Time 4.98
-ISO Speed Ratings 47.80
-The ISO speed ratings may inadvertently function as a mediating variable between trichome density and fertilizer stress, attributable to the camera’s automatic exposure (AE) function.
-The AE system responds to the overall image brightness.
-This phenomenon is considerably influenced by trichomes appearing as white points due to the reflection of light from the LED of the smartphone.
-Variations in trichome density, induced by fertilizer stress, alter image brightness, prompting automatic adjustments in ISO sensitivity.
-Consequently, the ISO speed rating becomes an indirect indicator of trichome density and, by extension, fertilizer stress.
-This intricate relationship positions ISO sensitivity as a mediator variable, not through direct measurement of trichomes or fertilizer stress but by reflecting the camera’s response to image characteristics indicative of these factors.
-The high VIF value for ISO speed ratings can be attributed to this complex interaction, where the camera’s real-time exposure adjustment algorithm effectively renders ISO sensitivity dependent on trichome density, which is, in turn, influenced by fertilizer stress.
-This unintended mediation effect introduces complications into the dataset: ISO sensitivity becomes correlated with the variable of interest (trichome density) and the target variable (fertilizer stress).
-Notably, although multicollinearity may not significantly affect the overall predictive performance of the model as evaluated by cross-validation, it can substantially impact the interpretation of individual variable importance.
-In the presence of multicollinearity, estimated coefficients for individual predictors may become unstable and unreliable.
-This instability can lead to biased assessments of variable importance, potentially resulting in overestimation or underestimation of the importance of certain variables in the model.
-In the context of this study, where both the generalization performance and model interpretability through cross-validation are paramount, it is imperative to exclude the ISO speed ratings as a variable because of their demonstrated multicollinearity.
-This exclusion enhances the robustness and interpretability of the statistical model, ensuring more reliable insights into the relationships between trichome density, fertilizer stress, and other relevant variables.
-The resolution value exhibits interactions and a shift in the intercept, necessitating its inclusion as a variable within the statistical model.
-An investigation of the effect of resolution on the relationship between trichome density and nitrate concentration in compound leaves reveals varying coefficients and intercepts in the regression equation.
-These changes in the coefficients with respect to the resolution are demonstrated in Table 5 and Figure 13.
-The resolution inherently varies because of factors such as camera performance and the distance between the camera sensor and the measurement paper.
-Consequently, in the dataset used in this study, the resolution is an indicator of the distance between the camera and the measurement paper.
-Thus, to avoid omitted variable bias, which could lead to misinterpretation of the effects of one variable as those of another, it is crucial to include resolution as a variable in the model.
-Furthermore, the varying coefficients demonstrate the advantages of using higher-resolution images for trichome detection in computer vision applications.
-The results show that increased resolution leads to improved detection accuracy, highlighting the importance of holding the measurement paper close to the camera and using a high-resolution image sensor for such tasks.
-Table 5: Interaction when Stratified Based on Resolution
-Percentile Resolution (pixel) Equation
-Higher than 85th $>$ $1.08\times 10^{6}$ $y=3.06\times 10^{-2}x+170$
-15th to 85th $1.08\times 10^{6}-5.56\times 10^{5}$ $y=1.90\times 10^{-2}x+205$
-Lower than 15th $<$ $5.56\times 10^{5}$ $y=1.28\times 10^{-2}x+254$
-Figure 13: Effect of resolution on the relationship between trichome density and nitrate concentration in leaves.
-Table 6: VIF Value after Dimensionality Reduction
-Feature VIF
-Nearest Neighbor Distance 8.63
-Resolution 6.49
-Exposure Time 3.10
-Table 6 displays the VIF values for all the variables except for the ISO speed ratings.
-After excluding ISO speed ratings, no variables with a VIF exceeding 10 were identified, indicating that multicollinearity was no longer a concern.
-To improve the interpretability of the model, ISO speed ratings were excluded as a variable in subsequent analyses because of their strong correlation with other variables and their potential to obscure the individual effects of each predictor.
-Instead, the modeling process included variables such as NND, resolution, and exposure time, which were deemed to have a more independent impact on the response variable.
-4.4 Performance of the Proposed Method
-Figure 14 shows the learning curve of the LightGBM model over 100 boosting rounds with default settings.
-A gradual increase is observed in the AUC-PR values for the training sets after each boosting round.
-This improvement indicates that the model has been adequately trained.
-Figure 14: Learning curve of the LightGBM model.
-Table 7 shows the predictive performance of the classification model at the different nitrate concentration thresholds.
-With respect to the trade-off between precision and recall, the model achieves highly balanced values.
-Table 7: Predictive Performance in Terms of the Need for Additional Fertilizer using LOOCV at Various Nitrate Ion (ppm) Thresholds when Using the 25 Image Features
-Evaluation Metrics 1600 1750 1900
-Precision 0.68 0.76 0.79
-Recall 0.66 0.76 0.79
-F1 Score 0.67 0.76 0.79
-PR AUC 0.76 0.84 0.87
-ROC AUC 0.64 0.68 0.58
-Table 8: Comparison of Predictive Performance between the Proposed Model and the Baseline Model in a Representative Parameter Setting
-Model Precision Recall F1 Score ROC AUC PR AUC
-Proposed 0.79 0.78 0.79 0.81 0.87
-Baseline 0.62 0.49 0.55 0.50 0.72
-For subsequent analyses, we defined a representative scenario using a nitrate concentration threshold of 1,783 ppm and 25 images captured from a single compound leaf.
-Using this representative scenario, predictions were made, allowing for an accurate assessment of the need for additional fertilization.
-The predictive performance, as evaluated by the mROC, is 0.64, and, as evaluated by the mPR, it is 0.82 despite variations in the measurement data caused by differences in optical conditions.
-Figure 15 shows a confusion matrix depicting true positives, false positives, true negatives, and false negatives.
-The predictions do not show substantial bias toward false positives or false negatives, indicating a balanced performance in identifying both classes correctly.
-This lack of bias is also evident from the balanced results presented in Table 7, which includes the F1 score and AUC-PR and AUC-ROC values, as well as the PR curve in Figure 16 and the ROC curve in Figure 17.
-Table 8 compares the performance of the proposed model and the baseline model.
-The baseline model uses a simple majority class predictor.
-This comparison utilizes the metrics defined in Equations 7-11, using the representative parameter setting.
-The table demonstrates the improved performance of the proposed model across all evaluated metrics.
-Figure 15: Confusion matrix for classification outcomes during LOOCV in a representative parameter setting.
-Figure 16: Precision-recall (PR) curve analysis of the model during LOOCV in a representative parameter setting.
-Figure 17: Receiver operating characteristic (ROC) curve analysis of the model during LOOCV in a representative parameter setting.
-Figure 18 illustrates the relationship between the number of input images and the model’s predictive performance.
-Using a single image yields a mean AUC-PR (mPR) of 0.80, and increasing the number of input images leads to better predictive performance.
-Figure 18: Model prediction performance when various thresholds and numbers of images are set.
-Figure 19 illustrates the global feature contributions based on SHAP values, providing insights into the overall model interpretation.
-NND has a high mean absolute SHAP value, suggesting its significant contribution to the model’s predictions across the dataset.
-Figure 19: Feature contribution during LOOCV with a representative model setting.
-4.5 Effect of Fertilizer on Trichome Density
-To quantitatively evaluate the relationship between trichome density on tomato leaves and nitrogen nutritional status, we conducted modeling via NND.
-Figure 20 presents the results of an NND prediction model employing nitrate ion concentration, image resolution, and exposure time as explanatory variables.
-The coefficient of determination ($R^{2}$) for this model is 0.63, indicating that the combination of explanatory variables (nitrate ion concentration) and moderating variables (image resolution, exposure time) can account for approximately 62.48% of the variation in trichome density (represented by NND) of leaflets obtained from fertilizer intervention experiments.
-Notably, this study was designed primarily to focus on the development of image-processing technology rather than being a rigorous hypothesis-testing study for interpreting natural phenomena in complex system.
-Typically, a hypothesis-based approach starts with a specific hypothesis or theory, and experiments are designed to test the validity of this hypothesis in a deductive manner [glass2008brief].
-In contrast, while a simple hypothesis existed for conducting experimental work as shown in Figure 3, this study primarily adopted a data-driven approach, where hypotheses and models are generated inductively based on patterns and relationships found within the dataset.
-Therefore, it is appropriate to interpret the results of this study as an analysis stemming from a data-driven approach.
-However, the relationship between trichome density on tomato leaves and nitrogen nutritional status has been quantitatively verified in previous studies through hypothesis-based research [hoffland2000nitrogen].
-As this study aims to capture general trends rather than subtle differences, the need for strict hypothesis-based experimentation is relatively low.
-Consequently, by adopting an inductive research method, the likelihood of fatal bias is considered low, and the data-driven approach of this study supports the possibility that fertilizer stress may affect trichome density, particularly in young leaves.
-Nevertheless, the existence of unexplained variation (approximately 37.52%) in the model suggests the need for further investigation of additional factors and model improvements in future research.
-Figure 20: Modeling of NND using nitrate concentration, resolution, and exposure time.
-5 Discussion
-Tomatoes pose a challenge for detecting fertilizer stress through noncontact optical measurements of young leaf chlorophyll.
-This challenge arises from the nonlinear relationship between optical measurements and leaf nitrogen content, which distinguishes tomatoes from other crops [padilla2018proximal, gianquinto2011a].
-To address this problem, our study focused on trichomes, which are densely packed hair-like structures on leaf surfaces that have been less explored in previous research.
-We investigated the feasibility of detecting fertilizer stress by developing a device that utilizes trichome density as a primary predictor.
-Our findings suggest the possibility of overcoming the limitations of conventional noncontact optical methods in assessing tomato fertilizer conditions.
-Our approach opens new avenues for developing diagnostic devices that consider leaf surface structure.
-However, the diagnostic kits developed in this study have some limitations regarding their performance.
-We validated the performance of the kit through experiments using a well-designed setup that simulated different levels of fertilizer stress commonly found in agricultural production conditions.
-In addition, the tomato cultivation method used in the present study involved cost-effective hydroponic cultivation via plastic buckets.
-The yields achieved were comparable to those obtained in commercial cultivation.
-This finding suggests that the cultivation method employed in this study closely simulated agricultural production conditions.
-However, it is uncertain whether all farmers will be able to establish optimal growing conditions for tomatoes to achieve higher yields, considering the financial costs involved in acquiring the necessary equipment.
-Consequently, environmental stresses beyond fertilizer stress could be encountered, which could decrease the predictive performance of our novel method under these circumstances.
-Traditionally, farmers have relied on visual inspection of leaf color, size, texture and overall plant health to assess the nutrient status of plants.
-However, the accuracy of this method highly depends on the farmers’ experience and motivation, resulting in considerable variations among individuals.
-Consequently, standardizing the assessment of nutrient status becomes problematic, making it difficult to compare and evaluate the effectiveness of different approaches.
-The absence of standardization in nutrient status assessment hinders highly advanced evaluation of the usefulness of diagnostic kits for a certain farmer solely on the basis of a comparison of their predictive accuracy.
-Thus, further demonstrations and validations employing more lagging indicators, such as financial metrics, are necessary to assess the practical applicability of these diagnostic kits in real-world agricultural settings.
-Although various nutrient diagnostic kits are commercially available for assessing plant nutrient status, their high cost and potential difficulties in application to delicate plant tissues, such as young tomato leaves, limit their widespread adoption in practical farming settings.
-To overcome these challenges, we propose a novel, simple, and cost-effective method for assessing nutrient status in tomato plants, with a specific focus on young leaves, which are typically relatively sensitive to nutrient deficiencies and play a crucial role in plant growth and development.
-In addition to tomatoes, various plant species, including agriculturally important crops such as watermelon (Citrullus lanatus), egg plants (Solanum melongena), tobacco (Nicotiana tabacum), soybean (Glycine max), and cannabis (Cannabis sativa), are known to have abundant trichomes.
-The proposed trichome-based method for assessing nutrient status has application potential in other plant species with dense trichomes near their growing points, expanding its scope and potentially improving real-time nutrient management practices in a wide range of crops, thereby enhancing the efficiency and precision of fertilizer application.
-By optimizing fertilizer application and reducing waste, this method could lead to increased crop yields, lower production costs, and reduced environmental impact, ultimately contributing to more sustainable agricultural practices.
-Further research is needed to validate the effectiveness of this method in various crop species and investigate the potential integration of this approach with other precision agriculture technologies, such as remote sensing, for more comprehensive and efficient nutrient management strategies.
-6 Conclusions
-This study focused on developing a novel method for detecting fertilizer stress in tomato plants by measuring the density of trichomes on their leaves.
-The study also included setting up and evaluating the experimental environment necessary for conducting these assessments.
-The developed diagnostic kit enables noninvasive and early detection of fertilizer stress in plants, obviating the need for customized sensing device design and permitting the use of a universal diagnostic kit at a lower cost than traditional devices while accurately assessing the nutrient status of tomato plants.
-The findings of this study present a novel approach to designing diagnostic devices for detecting fertilizer stress in plants by considering the structure of the plant surface.
-Appendix
-Appendix A Simulation Demonstrating the Robustness of the Nearest Neighbor Distance
-This appendix presents a simulation demonstrating the robustness of the nearest neighbor distance (NND) as a density metric compared with the conventional point count per unit area.
-The simulation involves generating a set of points within a defined area.
-Damage is simulated by removing a portion of these points and then comparing the changes in the NND and the point count.
-To mimic the random spatial distribution of trichomes on a leaf surface, we employed a Poisson point process for point generation.
-The Poisson process is often used to model random events occurring independently in a given space or time, making it suitable for simulating the distribution of trichomes, which are expected to emerge randomly on the leaf surface.
-The detailed steps are outlined in Algorithm 7.
-Algorithm 7 Trichome distribution simulation
-1:$\lambda$: Expected number of events (trichomes)
-2:$A$: Total area of the region
-3:$S$: Random seed value
-4:$P$: Set of points representing trichome locations
-5:$P^{\prime}$: Set of points after simulated damage
-6:function SetSeed($S$)
-7:Set a random seed to $S$
-8:end function
-9:function PoissonRandom($\lambda$)
-10:Generate random numbers from the Poisson distribution return number of points
-11:end function
-12:function UniformRandom($a,b$)
-13:Generate random numbers from a uniform distribution return random number between $a$ and $b$
-14:end function
-15:function GeneratePoints($N$)
-16:Generate $N$ random points in the area return set of points $P$
-17:end function
-18:function SimulateDamage($P$)
-19:Remove points with $x>500$ return set of remaining points $P^{\prime}$
-20:end function
-21:
-22:Main Execution
-23:SetSeed($S$)
-24:$N\leftarrow\textsc{PoissonRandom}(\lambda)$
-25:$P\leftarrow\textsc{GeneratePoints}(N)$
-26:$P^{\prime}\leftarrow\textsc{SimulateDamage}(P)$
-Figure 21 shows an example of the point distribution before and after simulated damage.
-The points were randomly generated via a Poisson process within a 1000 $\times$ 1000 area.
-The damage was simulated by removing points within the right half of the area, as indicated by the "damage boundary" line.
-Figure 22 shows the histograms of NNDs before and after simulated damage.
-Figure 21: Point distribution before and after simulated damage.
-The simulation results demonstrate that NND is more robust to point removal than the conventional density metric based on the point count per unit area.
-Figure 23 compares the NND and point count before and after simulated damage.
-Although both metrics decreased after damage, the decrease in NND was smaller than that in the point count.
-Figure 22: Histogram of nearest neighbor distances.
-To statistically validate this observation, we performed a Wilcoxon signed-rank test to compare the absolute change rates of the NND and the point count after damage, as shown in Figure 24.
-The test revealed a statistically significant difference between the two metrics ($p<0.001$), supporting the hypothesis that NND is less affected by localized point loss than the point count per unit area is.
-This finding highlights the advantage of using NND as a density metric in scenarios where localized point loss might occur, such as trichome detachment from leaf surfaces.
-Figure 23: Comparison of NND and point count after simulated damage.
-Figure 24: Comparison of NND and point count absolute change rates after simulated damage.
-Appendix B Comprehensive Model Performance across All Nitrate Ion Concentration Thresholds
-Although Table 7 in the main text focused on the nitrate ion concentration range of 1600-1900 ppm, this section provides a more comprehensive evaluation of the model’s performance across the entire range of nitrate ion concentrations observed in our study.
-Table 9 presents the detailed results for the proposed and baseline models.
-For this comprehensive analysis, we divided the full range of observed nitrate ion concentrations into 10 equal intervals: Each interval serves as a threshold for model evaluation.
+Appendix C Comprehensive Model Performance across All Nitrate Ion Concentration Thresholds
+This section extends our model evaluation beyond the 1600-1900 ppm nitrate ion concentration range presented in Table 7 of the main text to encompass the entire range of observed concentrations in our study.
+For a comprehensive assessment, we divided the full range of observed nitrate ion concentrations into 10 equal intervals, with each serving as a threshold for model evaluation.
 This approach allows for a thorough assessment of model performance across all possible thresholds within our dataset.
-Table 9 illustrates the predictive performance of the proposed and baseline models using LOOCV at these various nitrate ion concentration thresholds.
-Both models utilized features extracted from 25 images per compound leaf.
-For clarity, the table presents the results for the proposed model as the main values, with the corresponding results for the baseline model shown in parentheses below.
+Table 9 presents a detailed performance comparison between our proposed model and the baseline model, using LOOCV and features extracted from 25 images per compound leaf.
+The proposed model’s results are displayed as main values, with corresponding baseline model results shown in parentheses below, enabling direct performance comparison across all threshold intervals.
 Table 9: Predictive Performance Comparison between Proposed and Baseline Models using LOOCV at Various Nitrate Ion Concentration Thresholds
@@
 (0.95) (0.49) (0.65) (0.96) (0.50)
+Acknowledgments
+The authors would like to express their sincere gratitude to Prof.
+ZHANG Shuhuai for accepting them into his laboratory; to Prof.
+OHKAWA Hiroshi, Assoc., for his insights regarding the presence of trichomes on the surfaces of tomato leaves; and to Prof.
+Emeritus ARAKAWA Osamu from the Faculty of Agriculture and Life Science at Hirosaki University for his invaluable advice on experimental design.
+Their contributions were critical to the success of this research.
+Furthermore, we illustrated this manuscript via AI tools (OpenAI, ChatGPT).
+The authors reviewed and corrected the AI-generated images in Figure 5 and Figure 9 to ensure accuracy.
 […]